### v4

# Spectral Independence and Local-to-Global Techniques for Optimal Mixing of Markov Chains

Abstract
This monograph is an exposition on an exciting new technique known as spectral independence, which has been instrumental in analyzing the convergence rate of Markov Chain Monte Carlo (MCMC) algorithms. For a high-dimensional distribution defined on labelings of the vertices of an $n$-vertex graph, the spectral independence condition, introduced by Anari, Liu, and Oveis Gharan (2020), is a bound on the maximum eigenvalue of the $n\times n$ influence matrix whose entries capture the influence between pairs of vertices (closely related to the covariance between the variables). In the first part of the monograph, we present results showing that spectral independence (and related techniques) imply fast mixing of simple Markov chains such as the Gibbs sampler. These proofs utilize local-to-global theorems which we will detail in this work.
We focus on two applications: the hard-core model on independent sets of a graph (which is a binary graphical model) and random bases of a matroid. We demonstrate several methods for establishing spectral independence, including the correlation decay approach introduced by Weitz (2006) and the analytic approach introduced by Barvinok (2015) using zero-freeness of the associated complex-valued partition function. As a consequence of the techniques presented in this monograph we obtain fast mixing of the Gibbs sampler on general graphs in the so-called tree-uniqueness region, polynomial-time mixing on general graphs at the critical point for the uniqueness threshold, and polynomial-time mixing on random regular graphs beyond the uniqueness threshold. We also utilize the local-to-global framework of spectral independence and the Trickle-Down theorem of Oppenheim (2018) to prove fast mixing of the bases-exchange walk for generating a random basis of an arbitrary matroid, which was established by Anari, Liu, Oveis Gharan and Vinzant (2019).
Our focus in this monograph is on the analysis of the spectral gap of the associated Markov chains from a functional analysis perspective; we present proofs of the associated local-to-global theorems and the Trickle-Down Theorem from this same Markov chain perspective. The monograph is self-contained and aims to present the proofs in a unified fashion.
1 Introduction
This monograph aims to present the new technique of spectral independence together with its implications on the mixing time of the single-site update Markov chain known as the Glauber dynamics. Spectral independence is a simple, yet seemingly quite powerful technique to obtain optimal upper bounds on the convergence rate of Markov chains to its stationary distribution. We will introduce the spectral independence technique, explain the basic properties of the associated influence matrix, and discuss connections to simplicial complexes. We then prove that spectral independence implies fast mixing of the Glauber dynamics.
We also present the prominent result of Anari, Liu, Oveis Gharan, and Vinzant [ALOV19] which shows fast mixing of a Markov chain, known as the bases-exchange walk, for randomly sampling bases of a matroid. This utilizes the local-to-global theorems presented for spectral independence and the Trickle-Down Theorem of Oppenheim [Opp18] to bound the spectral gap of the associated local walks.
Our aim in this work is to present the analyses from a Markov chain perspective, and to be as self-contained as possible.
1.1 Setting
Our setting is high-dimensional probability distributions defined by a graphical representation; this setting corresponds to the equilibrium distribution for spin systems in statistical physics and undirected graphical models in machine learning. Let $G=(V,E)$ be a graph on $n=|V|$ vertices. Consider a distribution $\mu$ on a subset of $\{0,1\}^{V}$. Let $\Omega\subseteq\{0,1\}^{V}$ be the support of $\mu$, that is,
$\Omega=\left\{\sigma\in\{0,1\}^{V}:\mu(\sigma)>0\right\}.$
We refer to the above setting as a binary system since each vertex has an assignment in $\{0,1\}$. Many of the results in these notes generalize to multispin systems which are distributions defined on $\{1,\dots,q\}^{V}$ for an integer $q\geq 2$; we will comment on the generalization to $q\geq 3$ spins at the end of this monograph in Section 12.2, and outside that section of the text we only consider the binary $q=2$ case.
An example to keep in mind is the so-called hard-core model that we now define.
1.2 Running example: Hard-core Model
The hard-core model is defined by a graph $G=(V,E)$ and an activity $\lambda>0$. The motivation of the hard-core model in statistical physics is as an idealized model of a gas where $\lambda$ corresponds to the fugacity of the gas (the fugactiy is related to the density). Configurations of the model are defined on independent sets of $G$: an independent set is a subset $I\subset V$ of vertices which does not contain an edge, which means that for all $\{u,v\}\in E$ either $u\notin I$ and/or $v\notin I$. Let $\Omega=\Omega_{G}$ denote the collection of independent sets of $G$ (regardless of their sizes).
For an independent set $\sigma\in\Omega$, we can view $\sigma$ as an $n$-dimensional indicator vector in $\{0,1\}^{n}$ where the $i$-th coordinate is assigned 1 if the $i$-th vertex is in $\sigma$ and is assigned 0 otherwise. Thus,
$\Omega=\Omega_{G}=\{\sigma\in\{0,1\}^{V}:\forall\{v,w\}\in E,\sigma(v)+\sigma(w)\leq 1\}.$
For an independent set $\sigma$, when a vertex $v\in\sigma$ (corresponding to $\sigma(v)=1$) then we often refer to $v$ as occupied, whereas for $w\notin\sigma$ (corresponding to $\sigma(w)=0$) we refer to $w$ as unoccupied. Each independent set $\sigma\in\Omega$ is assigned the following weight:
$w(\sigma)=\lambda^{|\sigma|},$
where $|\sigma|$ is the number of vertices in the independent set $\sigma$, in other words, $|\sigma|=|\{v\in V:\sigma(v)=1\}|$. The Gibbs distribution $\mu=\mu_{G,\lambda}$ is the probability distribution on independent sets that is proportional to the weights. Specifically, for $\sigma\in\Omega$,
$\mu(\sigma)=\frac{\lambda^{|\sigma|}}{Z},$
where the normalizing factor $Z=Z_{G,\lambda}=\sum_{\eta\in\Omega}\lambda^{|\eta|}$ is known as the partition function.
1.3 Glauber dynamics/Gibbs sampler
The following Markov chain is a simple single-site update Markov chain designed to sample from the Gibbs distribution. The chain is known as the Glauber dynamics in the statistical physics community and as the Gibbs sampler in the machine learning community. We begin with the definition of the Glauber dynamics for the special case of the hard-core model (the reader new to the field can further consider the case $\lambda=1$ for simplicity).
From a state $X_{t}\in\Omega$, the transitions $X_{t}\rightarrow X_{t+1}$ of the Glauber dynamics for the hard-core model are defined as follows:
  1. 1.
    Choose a vertex $v$ uniformly at random from $V$.
  2. 2.
    Let
    $X^{\prime}=\begin{cases}X_{t}\cup\{v\}&\mbox{with probability }\lambda/(1+\lambda)\\ X_{t}\setminus\{v\}&\mbox{with probability }1/(1+\lambda)\end{cases}$
  3. 3.
    If $X^{\prime}\in\Omega$ (i.e., $X^{\prime}$ is a valid independent set) then set $X_{t+1}=X^{\prime}$, and otherwise set $X_{t+1}=X_{t}$.
Let $P$ denote the transition matrix of the Glauber dynamics. Note that $P$ is an $N\times N$ stochastic matrix where $N=|\Omega|$. The Glauber dynamics is irreducible and aperiodic and hence the Glauber dynamics is ergodic. To see that the dynamics is irreducible, let $\tau=\emptyset$ and notice that every state $\sigma\in\Omega$ can reach the empty set, i.e., $P^{t}(\sigma,\tau)>0$ for some $t\geq 0$, by successively removing vertices from $\sigma$, and similarly for every $\sigma^{\prime}\in\Omega$, $P^{t^{\prime}}(\tau,\sigma^{\prime})>0$ for some $t^{\prime}\geq 0$. Hence, $P^{s}(\sigma,\sigma^{\prime})>0$ for some $s\geq 0$ where this time $s=t+t^{\prime}$ depends on $\sigma,\sigma^{\prime}$ (though one can easily argue that there is a universal such time $s$). To see that the Glauber dynamics is aperiodic, notice that for every state $\sigma\in\Omega$ there is a so-called self-loop, namely $P(\sigma,\sigma)>0$, since the updated vertex $v$ can have $X^{\prime}(v)=X_{t}(v)$ with positive probability. Since the Glauber dynamics is ergodic, there is a unique stationary distribution $\mu$. Moreover, the Gibbs distribution $\mu$ is the unique stationary distribution, since it satisfies the detailed balance conditions, which state that for all pairs $\sigma,\sigma^{\prime}\in\Omega$, $\mu(\sigma)P(\sigma,\sigma^{\prime})=\mu(\sigma^{\prime})P(\sigma^{\prime},\sigma)$; hence, the Glauber dynamics is a reversible chain.
Before proceeding, we present the general definition of the Glauber dynamics for more general distributions. For a distribution $\mu$ on $\Omega\subset\{0,1\}^{V}$, the Glauber dynamics is defined as follows. From a state $X_{t}\in\Omega$,
  1. 1.
    Choose a vertex $v$ uniformly at random from $V$.
  2. 2.
    For all $w\neq v$, let $X_{t+1}(w)=X_{t}(v)$.
  3. 3.
    Choose $X_{t+1}(v)$ from the conditional Gibbs distribution $\mu(\sigma(v)=\cdot\mid\sigma(w)=X_{t}(w)\mbox{ for all }w\neq v)$. In words, fix the spin/label at all vertices except $v$ and resample the spin/label at $v$ conditional on the fixed configuration on the rest of the vertices.
The mixing time of the Glauber dynamics, denoted as $T_{\mathrm{mix}}(\varepsilon)$ for $\varepsilon>0$, is the minimum number of steps $t$ such that, for the worst initial state $X_{0}$, the distribution of $X_{t}$ is within total variation distance $\leq\varepsilon$ of the stationary distribution $\mu$. For a pair of distributions $\mu,\omega$ on $\Omega$, denote their total variation distance as follows:
$\|\mu-\omega\|_{\mathrm{TV}}=\frac{1}{2}\sum_{\sigma\in\Omega}|\mu(\sigma)-\omega(\sigma)|=\max_{S\subset\Omega}\mu(S)-\omega(S).$
Then the mixing time is defined as:
$T_{\mathrm{mix}}(\varepsilon)=\min\{t:\forall X_{0}\in\Omega,\|\mu-P^{t}(X_{0},\cdot)\|_{\mathrm{TV}}\leq\varepsilon\}.$
Setting $\varepsilon=1/4$ we will refer to $T_{\mathrm{mix}}=T_{\mathrm{mix}}(1/4)$ as the mixing time. The choice of constant $1/4$ is somewhat arbitrary as one can “boost” the mixing time to obtain, for any $\varepsilon>0$, $T_{\mathrm{mix}}(\varepsilon)\leq T_{\mathrm{mix}}(1/4)\times\lceil\log_{2}(1/\varepsilon)\rceil$. Oftentimes, the constant $1/(2e)$ is used in place of $1/4$ in the definition of the mixing time and then the corresponding boosting has $\ln(1/\varepsilon)$ in place of $\log_{2}(1/\varepsilon)$, see [LP17, Section 4.5].
We say the mixing time is optimal if $T_{\mathrm{mix}}=O(n\log{n})$ as this matches the lower bound established by Hayes and Sinclair [HS07] for any graph of constant maximum degree $\Delta$. Note an upper bound of $T_{\mathrm{mix}}=O(n\log{n})$ yields a more general bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log{n}\log(1/\varepsilon))$ via the above boosting result. However, the proofs of optimal mixing in these notes will yield a stronger upper bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$, and hence we will state this improved bound when relevant.
In contrast to the mixing time, a weaker notion of convergence is the relaxation time. For $N=|\Omega|$, let $\lambda_{1}=1>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{N}>-1$ denote the eigenvalues of the transition matrix $P$ of the Glauber dynamics (recall, the Glauber dynamics is reversible, i.e., satisfies the detailed balance conditions). Note, $\lambda_{1}$ corresponds to the principal eigenvector which is the stationary distribution $\mu$. Moreover, $\lambda_{2}<1$ since the Glauber dynamics is an ergodic Markov chain and hence the stationary distribution is unique. In addition, $\lambda_{N}>-1$ since the Glauber dynamics is reversible and aperiodic.
The relaxation time is defined as $T_{\mathrm{relax}}:=(1-\lambda_{*})^{-1}$ where $\lambda_{*}=\max\{\lambda_{2},|\lambda_{N}|\}$. The quantity $1-\lambda_{*}$ is known as the absolute spectral gap, see Section 3.4 for further discussion. The relaxation time measures the decay rate of the variance of the chain with respect to the stationary distribution; this will be formally presented in Section 3. The optimal relaxation time for the Glauber dynamics is $O(n)$, whereas the optimal mixing time is $O(n\log{n})$. An optimal mixing time of the form $O(n\log(n/\varepsilon))$ implies an optimal relaxation time (see Remark 9.30 in Section 9.5), but the reverse implication does not necessarily hold (see Section 3.7 for related results).
One of the primary applications of an (approximate) sampling algorithm is for the following problem of estimating the partition function. Consider the algorithmic problem of obtaining a Fully Polynomial-time Randomized Approximation Scheme ($\mathsf{FPRAS}$) for estimating the partition function $Z_{G,\lambda}$ of the hard-core model for a given $G=(V,E)$ and $\lambda\geq 0$; more precisely, our goal is to construct an algorithm which, given as input a graph $G=(V,E)$ and parameters $\lambda^{*}\geq 0$ and $\delta>0$, produces an estimate $\mathsf{EST}$ where:
$\mathrm{Pr}\left((1-\delta)Z_{G,\lambda}\leq\mathsf{EST}\leq(1+\delta)Z_{G,\lambda}\right)\geq 3/4,$
and runs in time $\mathrm{poly}(n,1/\delta)$ where $n=|V|$. (The success probability of $3/4$ can be boosted to $\geq 1-\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0$ with an additional $\log(1/\varepsilon^{\prime})$ factor in the running time.) Given a parameter $\lambda^{*}$, if for all $\lambda\leq\lambda^{*}$ we have an optimal mixing time of $O(n\log(n/\varepsilon))$ for the Glauber dynamics, then there is an $\mathsf{FPRAS}$ with running time $O((n/\delta)^{2}\log(n)\log(n/\delta))$ for estimating the partition function $Z_{G,\lambda^{*}}$ at $\lambda^{*}$, see [ŠVV09, Hub15, Kol18]. In fact, for the purposes of an approximate counting algorithm, it suffices to have a bound on the relaxation time (instead of the mixing time) of an appropriate sampling algorithm. In particular, if we have a relaxation time of $O(n)$ for all $\lambda\leq\lambda^{*}$ then we obtain an $\mathsf{FPRAS}$ with running time of $O^{*}((n/\delta)^{2})$ where the $O^{*}()$ notation hides logarithmic factors, see [ŠVV09, Section 7]. Moreover, we can estimate other quantities of interest, such as the number of independent sets of a particular size, see [HK25, DP23].
Further properties and fundamental definitions regarding Markov chains are presented in Section 3. We lower bound the spectral gap of the Glauber dynamics (and hence upper bound the relaxation time) using a functional analytic approach to analyze the decay rate of the variance of the Markov chain with respect to the stationary distribution; consequently, in Section 3.3 we introduce the variance functional and the Dirichlet form which is a measure of the local variation of a Markov chain. In Section 3.5 we present the notion of approximate tensorization of variance, which is a useful technique for lower bounding the spectral gap of the Glauber dynamics. The relations between the relaxation time and the mixing time of a Markov chain are established in Section 3.7.
1.4 Key Definitions: Spectral Independence and the Influence Matrix
Spectral independence was introduced by Anari, Liu, and Oveis Gharan [ALO20]. It is defined by an $n\times n$ influence matrix which captures the pairwise influence or correlations between pairs of vertices.
Definition 1.1 (Influence Matrix).
Let $G=(V,E)$ be a graph where $V=\{1,\dots,n\}$, and $\mu$ be a distribution on $\Omega$ where $\Omega\subseteq\{0,1\}^{V}$. Let $\Psi$ be the following real-valued $n\times n$ matrix; we refer to $\Psi$ as the influence matrix. For $1\leq i,j\leq n$, let
$\Psi_{\mu}(i,j)=\Psi(i\rightarrow j)$ $:=\mu\left(\sigma(j)=1\mid\sigma(i)=1\right)-\mu\left(\sigma(j)=1\mid\sigma(i)=0\right)$ (1)
$=\mathrm{Pr}_{\sigma\sim\mu}\left[\sigma(j)=1\mid\sigma(i)=1\right]-\mathrm{Pr}_{\sigma\sim\mu}\left[\sigma(j)=1\mid\sigma(i)=0\right].$
Note, when $i=j$ then $\Psi(i,i)=1$ and hence the diagonal entries of $\Psi$ are defined to be $1$ (see Remark 1.3 below for further discussion regarding an alternative definition of the diagonal entries).
A closely related notion is the modified influence matrix which we formally define in Section 2.5. The modified influence matrix differs from Eq. 1 in that the last term changes from $\mu\left(\sigma(j)=1\mid\sigma(i)=0\right)$ to $\mu\left(\sigma(j)=1\right)$. This leads to a weaker notion than spectral independence (we will later refer to the alternative notion as weak spectral independence); the connections between the influence matrix and modified influence matrix are discussed in further detail in Section 2.5.
We require the generalization of the above definitions to arbitrary “pinnings” which are a fixed assignment $\tau$ assigned to a subset $S$ of vertices. For a subset of vertices $S\subset[n]=\{1,\dots,n\}$, a pinning on $S$ is an assignment $\tau:S\rightarrow\{0,1\}$. Let $\Omega_{\tau}$ denote the assignments on all of $V$ which are consistent with $\tau$; more formally, for $S\subset V$ and a pinning $\tau:S\rightarrow\{0,1\}$, let
$\Omega_{\tau}:=\{\sigma\in\Omega:\sigma(S)=\tau(S)\},$
where $\sigma(S)$ is the restriction of $\sigma$ to $S$, and hence $\sigma(S)=\tau(S)$ iff $\sigma(v)=\tau(v)$ for all $v\in S$.
We only consider pinnings $\tau$ that can be extended to a valid configuration $\sigma$ on the entire vertex set $V$. Hence, we say a pinning $\tau$ is valid if $\Omega_{\tau}\neq\emptyset$. Let $\mathcal{P}$ denote the collection of all valid pinnings (see Section 4.1 for further introduction).
For a pinning $\tau\in\mathcal{P}$, let $\mu_{\tau}$ denote the conditional Gibbs distribution, in other words, $\mu_{\tau}$ is the distribution $\mu$ conditional on the fixed assignment $\tau$ on $S$. For $\sigma\in\Omega_{\tau}$, we have
$\mu_{\tau}(\sigma)=\mu(\sigma\mid\sigma(S)=\tau)=\frac{\mu(\sigma(\overline{S})\cup\tau(S))}{\mu(\Omega_{\tau})},$
$\mbox{where}\ \ \ \mu(\Omega_{\tau})=\sum_{\eta\in\Omega_{\tau}}\mu(\eta)=\sum_{\begin{subarray}{c}\eta\in\Omega:\\ \eta(S)=\tau(S)\end{subarray}}\mu(\eta).$
For the case of the hard-core model on a graph $G=(V,E)$, pinnings correspond to induced subgraphs of $G$ in the following manner. In the hard-core model, for $S\subset V$, a valid pinning $\tau$ on $S$ is an independent set on $S$. Since $\tau$ is an independent set, an equivalent realization of the pinning $\tau$ is to consider the induced subgraph on $V^{\prime}\subset V$ where $V^{\prime}$ is formed by starting from $V$ and then proceeding as follows: for every vertex $v\in S$ where $\tau(v)=0$ then remove $v$ (this is because fixing $\tau(v)=0$ means $v$ is fixed to be outside of the independent set), and for every vertex $v\in S$ where $\tau(v)=1$ then remove $v$ and all $w\in N(v)$ (since $\tau(v)=1$ then $v$ is fixed to be in the independent set and hence all neighbors of $v$ are not in the independent set); let $V^{\prime}$ be the remaining vertices. Let $H$ be the induced subgraph on $V^{\prime}$. The distribution $\mu_{H}$ (which is the Gibbs distribution on $H$ without any pinning) is identical to the distribution $\mu_{\tau}$ (which is the conditional Gibbs distribution on $G$ with pinning $\tau$). Hence, for the case of the hard-core model, the collection of valid pinnings for $G$ can be realized by considering all induced subgraphs of $G$ (without any pinnings).
Returning to the general setting, for a pinning $\tau\in\mathcal{P}$ where $\tau:S\rightarrow\{0,1\}$ for $S\subset V$, let $T\subset V\setminus S$ denote the set of vertices which are “free” in the following sense:
[…]
Note, a vertex $i\in V\setminus(S\cup T)$ is “frozen”, i.e., it can only attain one spin and hence we can fix the configuration on that vertex and remove the corresponding two rows/columns from the matrix $\Psi_{\tau}$; henceforth, we only define the influence matrix on the free vertices.
For $i,j\in T$, let
$\Psi_{\mu_{\tau}}(i\rightarrow j)$ $:=\mu_{\tau}\left(\sigma(j)=1\mid\sigma(i)=1\right)-\mu_{\tau}\left(\sigma(j)=1\mid\sigma(i)=0\right)$ (2)
$=\mu\left(\sigma(j)=1\mid\sigma(i)=1,\sigma(S)=\tau\right)-\mu\left(\sigma(j)=1\mid\sigma(i)=0,\sigma(S)=\tau\right).$
Since the empty pinning $\emptyset$ is always a valid pinning, thus $\Psi=\Psi_{\emptyset}$ is identical to the definition in Eq. 1. As in Eq. 1, observe that the diagonal entries of $\Psi_{\tau}$ are defined to be $1$, see Remark 1.3 for further discussion.
The matrix $\Psi$ can be asymmetric, and the entries of $\Psi$ can be positive or negative. Nevertheless, all of the eigenvalues of $\Psi$ are real and non-negative, this is established in Lemma 2.2 in Section 2.1. Intuitively, the fact that $\Psi$ has real non-negative eigenvalues is because it is closely related to the covariance matrix $\mathrm{Cov}$. In particular, the influence matrix $\Psi$ is equal to a positive diagonal matrix times the covariance matrix $\mathrm{Cov}$, and then the real non-negative eigenvalues of $\Psi$ follows from the positive semidefiniteness of the covariance matrix.
Since all eigenvalues of $\Psi$ are real valued we can denote the maximum eigenvalue by $\lambda_{\max}(\Psi)$. Now we can define the spectral independence condition as follows.
Definition 1.2 (Spectral Independence).
For $\eta\geq 0$, we say that $\mu$ is $\eta$-spectrally independent if for all pinnings $\tau\in\mathcal{P}$,
$\lambda_{\max}(\Psi_{\mu_{\tau}})\leq 1+\eta.$ (3)
Note that the spectral independence condition only depends on the distribution $\mu$, and it has no explicit dependence on the Glauber dynamics. The definition was extended to non-binary spin spaces, such as the Potts model and colorings, in [FGYZ21, CGŠV21] (see also [CLV21a, BCC+22] for a general formulation); we present these generalizations and discuss the related results in Section 12.2.
When $\mu$ is a product distribution then $\eta=0$. Our goal is to show that $\eta$ is constant, and hence in some sense it is not far from a product distribution.
Remark 1.3.
Note the diagonals of the influence matrix $\Psi$ are 1 since if $i=j$ then conditioning on $i$ prescribes $j$. We could have defined the influence matrix so that the off-diagonal entries remain the same and the diagonals are 0; this would decrease all of the eigenvalues by 1, and hence with this alternative definition we would change the spectral independence requirement from $1+\eta$ to $\eta$.
For the special case of the hard-core model, the configurations of the model are independent sets. Hence, we denote a configuration by the set of vertices in the independent set $I\subset V$. We refer to those vertices $v\in I$ as occupied and those $w\notin I$ as unoccupied. In this case of the hard-core model by referring to the subset of occupied vertices $I\subset V$ this simplifies the notation. For example, returning to the definition of the influence matrix in Definition 1.1, for the hard-core model on $G=(V,E)$, for $u,v\in V$, the $(u,v)$-entry of the influence matrix $\Psi$ can be compactly expressed as the following:
$\Psi(u\rightarrow v):=\mu(v|u)-\mu(v|\overline{u}),$
where $u,v$ denote the events $u\in I,v\in I$ respectively, and $\overline{u}$ denotes the event $u\notin I$.
As discussed earlier, for the hard-core model, since the configurations correspond to independent sets then a pinning $\tau$ on $S$ can be realized by considering the appropriate induced subgraph $H$ of $G$. Consequently, for a graph $G=(V,E)$ if we show for all induced subgraphs $H$ of $G$ that $\lambda_{\max}(\Psi_{H})\leq 1+\eta$ (without any pinnings) then that implies the spectral independence condition Definition 1.2 holds for $G$ (since every pinning $\tau$ can be realized by an appropriate induced subgraph $H$).
In Section 2 we prove several fundamental properties of the influence matrix $\Psi$, including the fact that the eigenvalues of $\Psi$ are real and non-negative; we also present several basic techniques for bounding the maximum eigenvalue of the influence matrix $\Psi$ in Sections 2.2, 2.3 and 2.4, and introduce the notion of weak spectral independence in Section 2.5.
1.5 Main Results: Fast Mixing via Spectral Independence
Anari, Liu, and Oveis Gharan [ALO20] proved the following result that spectral independence implies polynomial mixing time of the Glauber dynamics.
Theorem 1.4 ([ALO20]).
Suppose $\mu$ is a distribution fully supported on $\Omega\subset\{0,1\}^{n}$ such that the Glauber dynamics for $\mu$ is ergodic. If $\mu$ is $\eta$-spectrally independent, then the relaxation time of the Glauber dynamics for $\mu$ satisfies:
$T_{\mathrm{relax}}\leq O(n^{1+\eta}).$
Therefore, the mixing time satisfies $T_{\mathrm{mix}}=O(n^{1+\eta}\log(1/\mu^{*}))$ where $\mu^{*}:=\min_{\sigma\in\Omega}\mu(\sigma)$.
For the hard-core model with fugacity $\lambda$, note that $\log(1/\mu^{*})=O_{\lambda}(n)$ and hence Theorem 1.4 gives a mixing time upper bound of $O(n^{2+\eta})$.
The above result was improved by Chen, Liu, and Vigoda [CLV21a] to show optimal bounds on the relaxation time when the maximum degree is constant.
Theorem 1.5 ([CLV21a]).
For all constant $\Delta\geq 2$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$ and fugacity $\lambda>0$, if the hard-core model on $G$ is $\eta$-spectrally independent for a constant $\eta>0$, then the relaxation time of the Glauber dynamics satisfies:
$T_{\mathrm{relax}}\leq O(n).$
In particular, there exists a constant $C(\eta,\Delta)$ where the relaxation time satisfies:
$T_{\mathrm{relax}}\leq C(\eta,\Delta)n.$
Therefore, the mixing time satisfies $T_{\mathrm{mix}}=O(n^{2})$.
The above result shows that spectral independence implies optimal relaxation time. In fact, spectral independence is a necessary condition for optimal relaxation time: Anari, Jain, Koehler, Pham, and Vuong [AJK+24] proved that optimal relaxation time of the Glauber dynamics implies spectral independence.
Theorem 1.6.
Let $\mu$ be a distribution on $\{0,1\}^{n}$. If the Glauber dynamics for $\mu$ has relaxation time bounded by $\leq Cn$ then $\lambda_{\max}(\Psi_{\mu})\leq C$. Furthermore, if for any pinning $\tau$ on a subset $S$ of vertices, the Glauber dynamics for $\mu_{\tau}$ has relaxation time bounded by $\leq C\cdot(n-|S|)$, then $\mu$ is (C1)(C-1)-spectrally independent.
If one further assumes a lower bound on the marginal probability for every vertex and every valid pinning, then Chen, Liu and Vigoda [CLV21a] proved an optimal upper bound on the mixing time, rather than the relaxation time.
Definition 1.7 (Marginally bounded).
For $b>0$, we say that $\mu$ is $b$-marginally bounded if for all pinnings $\tau\in\mathcal{P}$, for all $v\in V$, all $s\in\{0,1\}$, then the following holds:
$\mu_{\tau}(\sigma(v)=s)>b\mbox{ or }\mu_{\tau}(\sigma(v)=s)=0.$
Theorem 1.8 ([CLV21a]).
For all constant $\Delta\geq 2$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$ and fugacity $\lambda>0$, if the hard-core model on $G$ is $\eta$-spectrally independent and $b$-marginally bounded for constants $\eta,b>0$, then the mixing time of the Glauber dynamics satisfies:
$T_{\mathrm{mix}}\leq O(n\log{n}).$
In particular, there exists a constant $C(\eta,\Delta,b)$ where the mixing time satisfies for all $\varepsilon>0$,
$T_{\mathrm{mix}}(\varepsilon)\leq C(\eta,\Delta,b)n\log(n/\varepsilon).$
As noted in the theorem statement, Theorem 1.4 holds for any distribution on $\Omega\subseteq\{0,1\}^{n}$ where the Glauber dynamics is ergodic. In contrast, Theorems 1.5 and 1.8 hold for any 2-spin system (i.e., binary graphical model with pairwise interactions), the generalizations to $q>2$ spins are presented in Section 12.2. Examples of 2-spin systems are the hardcore model, the Ising model, and matchings (when viewed as independent sets of the line graph).
As pointed out earlier, Theorem 1.8 establishes an optimal bound on the mixing time of $O(n\log{n})$ as Hayes and Sinclair [HS07] proved that for any graph with constant maximum degree $\Delta$, the mixing time is $\geq C^{\prime}(\Delta)n\log{n}$ for some constant $C^{\prime}(\Delta)$.
Closely connected to the influence matrix is a notion of a local walk, which is introduced in Section 4. The local walk is a Markov chain which is defined on spin assignments for individual vertices in such a way that the spectral gap of the local walk is closely related to the maximum eigenvalue of the influence matrix, and hence the spectral independence property yields a lower bound on the spectral gap of the local walk.
The convergence results stated above (Theorems 1.4, 1.5 and 1.8) are established using a so-called local-to-global theorem of Alev and Lau [AL20], which is known as the Random Walk Theorem, presented in Section 4. The Random Walk Theorem relates the spectral gap of the local walk to the spectral gap of the Glauber dynamics, which corresponds to the global walk in the framework presented in Section 4. Applying the Random Walk Theorem, in Section 4.5, we prove Theorem 1.4 of Anari, Liu, and Oveis Gharan [ALO20]which shows that spectral independence implies polynomial mixing of the Glauber dynamics where the exponent in the polynomial depends on the spectral independence constant.
The Random Walk Theorem is proved in Section 5, where we introduce the associated up and down processes and explore their connection to the local walk. Subsequently, in Section 6, we present the proof of Theorem 1.5 of Chen, Liu, and Vigoda [CLV21a], which shows that spectral independence implies an optimal bound on the spectral gap of the Glauber dynamics; this proof utilizes an improved version of the Random Walk Theorem which is presented in Section 6.2 and proved in Section 6.6. Finally, in Section 6.8 we present the high-level ideas used to further extend the proof from variance contraction to entropy contraction and obtain optimal mixing time. Theorem 1.6 is proved in Section 9.5.
1.6 Application: Hard-core Model
An exciting consequence of the optimal mixing result stated in Theorem 1.8 is a beautiful computational phase transition for the hard-core model. In particular, for every $\Delta\geq 3$ there is a critical point $\lambda_{c}(\Delta)$ where a computational phase transition occurs on graphs of maximum degree $\Delta$. Namely, for all $\lambda<\lambda_{c}(\Delta)$, the Glauber dynamics has optimal mixing time on any graph of maximum degree $\Delta$; this is established using Theorem 1.8. On the other side, when $\lambda>\lambda_{c}(\Delta)$ then it is computationally hard to generate samples from a distribution which is close to the Gibbs distribution on all graphs of maximum degree $\Delta$.
The critical point $\lambda_{c}(\Delta)$ is a consequence of phase transition for the Gibbs distribution on $\Delta$-regular trees. The critical point $\lambda_{c}(\Delta)$ is defined as follows:
$\lambda_{c}(\Delta):=\frac{(\Delta-1)^{\Delta-1}}{(\Delta-2)^{\Delta}}.$
Note, $\lambda_{c}(\Delta)$ is a decreasing function, and for all $\Delta\leq 5$ then $\lambda_{c}(\Delta)>1$, while for $\Delta\geq 6$ then $\lambda_{c}(\Delta)<1$ ($\lambda_{c}(\Delta)=1$ at $\Delta\approx 5.14104$).
The function $\lambda_{c}(\Delta)$ corresponds to the critical point for the following decay of correlations property on trees. Fix an integer $\Delta\geq 3$. For an integer $\ell\geq 1$, let $T_{\ell}$ denote the complete $\Delta$-regular tree of height $\ell$. More precisely, $T_{\ell}$ is the tree where the root is at level 0, all leaves are at level $\ell$, and every non-leaf vertex has $\Delta-1$ children. (Whether the root vertex has degree $\Delta$ or $\Delta-1$ makes negligible difference to the following discussion.)
Let $\mu_{\ell}$ denote the Gibbs distribution on the tree $T_{\ell}$, and let $p_{\ell}$ denote the probability that the root vertex is in an independent set drawn from the distribution $\mu_{\ell}$. Note, the maximum sized independent set in $T_{\ell}$ contains all of the leaves and then every other level (as we move up the tree). Thus, for even $\ell$ the root is in the maximum sized independent set, and for odd $\ell$ the root is not in the maximum sized independent set. Consequently, for large $\lambda$, we expect that $p_{2\ell}>p_{2\ell+1}$ as the root has a bias to be in the independent set for even heights and a bias to be out of the independent set for odd heights.
There is a phase transition on complete $\Delta$-regular trees at $\lambda_{c}(\Delta)$ in the following sense. More formally, for all $\Delta\geq 3$ the following holds:
$\mbox{For all }\lambda\leq\lambda_{c}(\Delta),$ $\lim_{\ell\rightarrow\infty}p_{2\ell+1}=\lim_{\ell\rightarrow\infty}p_{2\ell}$
$\mbox{For all }\lambda>\lambda_{c}(\Delta),$ $\lim_{\ell\rightarrow\infty}p_{2\ell+1}<\lim_{\ell\rightarrow\infty}p_{2\ell}$
This phase transition is derived and explained in more detail in Section 9.1.
In statistical physics terminology, $\lambda_{c}(\Delta)$ is the critical point for the uniqueness/non-uniqueness phase transition on the infinite $\Delta$-regular tree. A Gibbs measure for the hard-core model on the infinite $\Delta$-regular tree is any measure such that for every finite subset of vertices and any pinning outside the subset, the conditional distribution gives a hard-core model with the corresponding pinning. In particular, we can define two extremal Gibbs measures, $\mu_{\mathrm{even}}$ and $\mu_{\mathrm{odd}}$, which can be obtained by taking the limit of finite-volume Gibbs distributions on $T_{\ell}$ with even and odd boundary conditions, respectively, which are fixed assignments (i.e., pinnings) to all of the leaves.
When $\lambda\leq\lambda_{c}(\Delta)$ then the extremal measures are the same $\mu_{\mathrm{even}}=\mu_{\mathrm{odd}}$ and hence there is a unique infinite-volume Gibbs measure; this is called the tree uniqueness region. On the other side, when $\lambda>\lambda_{c}(\Delta)$ then $\mu_{\mathrm{even}}\neq\mu_{\mathrm{odd}}$ and hence there are multiple infinite-volume Gibbs measures; this is called the tree non-uniqueness region. We refer the interested reader to [Geo11, FV17] for a more comprehensive introduction to infinite-volume Gibbs measures and related phase transitions.
The amazing aspect is that the decay of correlations property on $\Delta$-regular trees directly relates to the computational phase transition on all graphs of maximum degree $\Delta$. We can now formally state the optimal mixing result in the tree uniqueness region.
Theorem 1.9.
For all $\Delta\geq 3$, all $\delta>0$, there exists $C(\Delta,\delta)$, for all $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the mixing time of the Glauber dynamics for the hard-core model satisfies for all $\varepsilon>0$,
$T_{\mathrm{mix}}(\varepsilon)\leq C(\Delta,\delta)n\log(n/\epsilon).$
Remark 1.10.
The independent works of Chen, Feng, Yin, and Zhang [CFYZ22] and Chen and Eldan [CE22] improved Theorem 1.9 to obtain $T_{\mathrm{mix}}\leq C(\delta)n\log(n/\varepsilon)$ and thus extend the theorem to classes of graphs where the maximum degree $\Delta$ grows as a function of $n$. We will show in Section 10 (see Remark 10.2) how the tools presented for the proof of the following theorem (Theorem 1.11) yield $T_{\mathrm{relax}}\leq C(\delta)n$, thus extending Theorem 1.9 to unbounded degrees for the relaxation time.
A remarkable result of Chen, Chen, Yin, and Zhang [CCYZ25] proved polynomial mixing time for the Glauber dynamics all the way up to and including the critical point.
Theorem 1.11 ([T]heorem 1.1).
CCYZ-critical-hard-core] There exist constants $C,C^{\prime}>0$, such that for all $\Delta\geq 3$, for any graph $G=(V,E)$ of maximum degree $\Delta$, the relaxation time of the Glauber dynamics for the hard-core model at fugacity $\lambda\leq\lambda_{c}(\Delta)$ satisfies
$T_{\mathrm{relax}}\leq C^{\prime}n^{C}.$
Theorem 1.11 is proved in Section 10.
On the other side, it is computationally hard to approximately sample from the Gibbs distribution and approximate the partition function in the tree non-uniqueness region; in fact, not only is it hard to obtain an $\mathsf{FPRAS}$ (i.e., an arbitrarily close approximation) for the partition function, it is hard to approximate the partition function within some exponential factor even when restricted to triangle-free graphs. The below result is not proved in this monograph, we refer the reader to [GŠV15, CGG+16] for more results in this context (including so-called #BIS-hardness on bipartite graphs in the tree non-uniqueness region).
Theorem 1.12 ([Sly10, SS14, GŠV16, GŠV15]).
For all $\Delta\geq 3$, all $\delta>0$, there exists $\varepsilon=\varepsilon(\Delta,\delta)$, for all $\lambda\geq(1+\delta)\lambda_{c}(\Delta)$, assuming $\mathsf{RP}\neq\mathsf{NP}$, there is no algorithm which for any triangle-free graph $G=(V,E)$ of maximum degree $\Delta$ produces an estimate $\mathsf{EST}$ where:
$\mathrm{Pr}\left(Z_{G,\lambda}2^{-\varepsilon n}\leq\mathsf{EST}\leq Z_{G,\lambda}2^{\varepsilon n}\right)\geq 3/4,$
and runs in time polynomial in $n=|V|$.
The above hardness result is for general graphs, or triangle-free graphs. If we restrict attention to a restricted class of graphs then the threshold for rapid mixing of the Glauber dynamics might occur at a different point. A natural class of graphs to consider are random $\Delta$-regular graphs. Since these graphs are locally tree-like (in particular, they have few short cycles with high probability over the choice of graphs) then we might expect that the tree uniqueness threshold $\lambda_{c}(\Delta)$ is still the threshold for rapid mixing. Surprisingly, Chen, Chen, Chen, Yin and Zhang [CCC+25] showed that when restricted to random $\Delta$-regular graphs then we can obtain fast mixing beyond the uniqueness threshold.
Theorem 1.13 ([T]heorem 1.1).
CCCYZ-random-regular] For all $\Delta\geq 3$, there exists $C(\Delta)$ where the following holds. For $G$ chosen uniformly at random from all $\Delta$-regular graphs, then with probability 1o(1)1-o(1) over the choice of $G$, for all $\lambda<\frac{1}{3\sqrt{\Delta}}$ the Glauber dynamics for the hard-core model on $G$ at fugacity $\lambda$ satisfies:
$T_{\mathrm{mix}}(\varepsilon)\leq C(\Delta)n\log(n/\varepsilon).$
Theorem 1.13 is proved in Section 11. Inspired by Theorem 1.13 it is natural to conjecture that the rapid mixing threshold on random $\Delta$-regular graphs is closely related to the reconstruction threshold, which considers the influence of typical boundary conditions on complete $\Delta$-regular trees, as opposed to worst-case boundaries as considered in the uniqueness threshold; we refer the interested reader to [BST16, RŠV+14] for further discussion regarding the reconstruction threshold in the hard-core model.
We outline results for other spin systems in Section 12.2, including the Ising model results in Section 12.1 and colorings results in Section 12.2.
1.7 Methods for Establishing Spectral Independence
In Section 9 we present general techniques for establishing spectral independence, focusing on three approaches as outlined below. As a consequence, for the hard-core model, we prove that spectral independence holds up to the critical point $\lambda_{c}(\Delta)$ on any graph of maximum degree $\Delta$.
Theorem 1.14.
For all $\Delta\geq 3$, all $\delta>0$, for all $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the Gibbs distribution $\mu_{G,\lambda}$ is $\eta(\delta)$-spectrally independent where $\eta(\delta)\leq 16\sqrt{2}/\delta$.
Combining Theorem 1.14 with the results from Section 1.5, which show that spectral independence implies fast mixing, yields the majority of the fast mixing results for the hard-core model stated in Section 1.6. Towards proving Theorem 1.14, we present several techniques for establishing spectral independence.
In Section 9.2 we present the correlation-decay approach of Weitz [Wei06] and how it yields spectral independence. At a high-level, we show that to upper bound the row-sum of the influence matrix, which is the sum of the influences of a fixed vertex, one can consider influences in a tree, where the tree corresponds to self-avoiding walks in the graph. Then one can use the potential function method to prove that the sum of the influences in the tree decay exponentially fast (in the distance) when the fugacity is in the tree-uniqueness region. In Section 9.3 we show that stability of the partition function, so-called zero-freeness, also implies spectral independence; such conditions were used in the approximate counting algorithm introduced by Barvinok [Bar16a]. Finally, in Section 9.5, we show that a contractive coupling, and more generally optimal relaxation time for a local chain, implies spectral independence.
1.8 Matroids and Trickle-Down Theorem
We utilize the spectral independence approach to establish fast mixing of a Markov chain, known as the bases-exchange walk, for generating a random basis of a matroid, see Section 7 for a more comprehensive introduction and the associated proofs. Anari, Liu, Oveis Gharan, and Vinzant [ALOV19] presented the first proof of fast mixing of the bases-exchange walk for arbitrary matroids. In fact, the work of [ALOV19] for matroids inspired the idea of spectral independence, however we present these results in reverse chronological order. We do not directly utilize spectral independence to prove fast mixing of the bases-exchange walk, instead we utilize the framework of down-up chains and the associated Random Walk Theorem (Theorem 4.1) which are presented in Section 5. One additional tool utilized for the analysis of the bases-exchange walk is the Trickle-Down Theorem of Oppenheim [Opp18] which is used to bound the spectral gap of the local walks considered in the Random Walk Theorem, see Section 8.
We refer the reader to Section 7 for an introduction to matroids and associated terminology, including bases of a matroid. A matroid $\mathcal{M}=(E,\mathcal{I})$ consists of a ground set $E$ and a collection of independent sets $\mathcal{I}$. For the example of a graphic matroid the ground set is the set of edges $E$ of a graph $G=(V,E)$. The independent sets are subsets of the ground set with the requirement that every subset of an independent set is also an independent set (hence the collection $\mathcal{I}$ is downward closed), and the collection of independent sets also need to satisfy the so-called exchange property, see Section 7. The exchange property implies that all maximal independent sets are of the same size; these maximal independents are referred to as the bases of the matroid and their size is the rank of the matroid. For the example of a graphic matroid, the independent sets are acyclic subgraphs (viewed as subsets of edges) of a graph $G=(V,E)$ and the bases are spanning forests of $G$.
The bases-exchange walk is a Markov chain on the bases of a matroid where the transitions utilize the exchange property in the following manner: from a basis $B_{t}$ at time $t$, we choose an element $e\in B_{t}$ uniformly at random, and an element $f$ uniformly from those elements such that $B^{\prime}=B_{t}\cup\{f\}\setminus\{e\}$ is a basis, and then we set $B_{t+1}=B^{\prime}$ as the state at time $t+1$. The main result presented in Section 7 is a bound on the relaxation time of the bases-exchange walk.
Theorem 1.15 ([ALOV19]).
For a matroid $\mathcal{M}=(E,\mathcal{I})$, the relaxation time of the bases-exchange walk satisfies $T_{\mathrm{relax}}\leq r(\mathcal{M})$, where $r(\mathcal{M})$ is the rank of the matroid. Consequently, the mixing time of the bases-exchange walk satisfies $T_{\mathrm{mix}}=O(r(\mathcal{M})\log(|\mathcal{I}|))=O(r(\mathcal{M})^{2}\log n)$ where $n=|E|$.
The mixing time bound was improved by Cryan, Guo, and Mousa [CGM21] to $O(r(\mathcal{M})\log{\log(|\mathcal{I}|)})=O(r(\mathcal{M})\log r(\mathcal{M})+r(\mathcal{M})\log{\log{n}})$; moreover, the $\log\log n$ term was subsequently removed in [ALO+21], achieving $O(r(\mathcal{M})\log r(\mathcal{M}))$ mixing time.
The results for generating a random basis of a matroid are presented beginning in Section 7 where we review the relevant definitions for matroids and present the main result for the mixing time of the bases-exchange walk. The analysis of the bases-exchange walk is presented in Section 8; the proof utilizes the Random Walk Theorem, presented earlier in Section 5, and the Trickle-down Theorem, which is presented in Section 8.2. The Trickle-down Theorem is used to obtain an optimal bound on the spectral gap of the associated local walks. The proof of fast mixing of the bases-exchange walk is presented in Section 8.3, and the proof of the Trickle-down Theorem is presented in Section 8.5.
2 Properties of the Influence Matrix
Here we establish several basic properties of the influence matrix $\Psi$, defined in Section 1.4. In Section 2.1 we prove that the eigenvalues of the influence matrix $\Psi$ are real-valued and non-negative. In Sections 2.2, 2.3 and 2.4 we present several useful techniques for bounding the maximum eigenvalue of the influence matrix and hence establishing spectral independence. We first present in Section 2.2 a simple approach for bounding the maximum eigenvalue by the maximum absolute rowsum. Since the influence matrix is not symmetric it is often convenient to consider a symmetrized version of the influence matrix, which we present in Section 2.3. In Section 2.4 we introduce a semidefinite inequality which is equivalent to spectral independence.
In Section 2.5 we present the modified influence matrix $\widetilde{\Psi}$ which is closely related to the definition of the influence matrix. The modified influence matrix leads to a weaker notion, hence referred to as weak spectral independence; the connections are discussed in Section 2.5.
2.1 Nonnegative eigenvalues
Here we will prove that the influence matrix has non-negative real eigenvalues by relating it to the covariance matrix.
For a matrix $A$, its signature is the number of positive, negative, and zero eigenvalues. We will utilize the following basic property regarding the signature of a symmetric matrix.
Lemma 2.1.
For a symmetric matrix $A$ and a positive diagonal matrix $D$, the signature of $A$ and $DA$ are the same.
Proof.
Note that all eigenvalues of $A$ are real since it is symmetric. Observe that $DA$ is similar to the matrix $D^{-1/2}(DA)D^{1/2}=D^{1/2}AD^{1/2}$ which is a symmetric matrix and hence also has real eigenvalues. Thus, $DA$ and $A$ have the same signature by Sylvester’s law of inertia, which says that for a symmetric matrix $A$ and a non-singular matrix $B$, the signature of $A$ and $BAB^{T}$ are the same, where the signature of a matrix refers to the number of positive, negative, and zero eigenvalues. ∎
Lemma 2.2.
All eigenvalues of $\Psi$ are non-negative real numbers.
Proof.
For every $i,j\in[n]$, the covariance of $\sigma(i),\sigma(j)$ is given by
$\mathrm{Cov}_{\mu}(i,j)=\mathrm{Cov}_{\sigma\sim\mu}(\sigma(i),\sigma(j))$ $=\mathbb{E}_{\mu}[\sigma(i)\sigma(j)]-\mathbb{E}_{\mu}[\sigma(i)]\cdot\mathbb{E}_{\mu}[\sigma(j)]$ (4)
$=\mu\left(\sigma(i)=\sigma(j)=1\right)-\mu(\sigma(i)=1)\cdot\mu(\sigma(j)=1)$
$=\mu(\sigma(i)=1)\cdot\Big{(}\mu(\sigma(j)=1\mid\sigma(i)=1)-\mu(\sigma(j)=1)\Big{)}.$
Plugging in
$\mu(\sigma(j)=1)=\mu(\sigma(j)=1\mid\sigma(i)=1)\cdot\mu(\sigma(i)=1)+\mu(\sigma(j)=1\mid\sigma(i)=0)\cdot\mu(\sigma(i)=0),$
we obtain that
$\mathrm{Cov}_{\mu}(i,j)$
$=\mu(\sigma(i)=1)\cdot\Big{(}\mu(\sigma(j)=1\mid\sigma(i)=1)-\mu(\sigma(j)=1\mid\sigma(i)=1)\cdot\mu(\sigma(i)=1)$
$\hskip 90.3375pt-\mu(\sigma(j)=1\mid\sigma(i)=0)\cdot\mu(\sigma(i)=0)\Big{)}$
$=\mu(\sigma(i)=1)\cdot\Big{(}\mu(\sigma(j)=1\mid\sigma(i)=1)\cdot(1-\mu(\sigma(i)=1))-\mu(\sigma(j)=1\mid\sigma(i)=0)\cdot\mu(\sigma(i)=0)\Big{)}$
$=\mu(\sigma(i)=1)\cdot\mu(\sigma(i)=0)\cdot\Big{(}\mu(\sigma(j)=1\mid\sigma(i)=1)-\mu(\sigma(j)=1\mid\sigma(i)=0)\Big{)}.$ (5)
Let $\mathrm{Var}_{\mu}$ denote the vector of variances, with
$\mathrm{Var}_{\mu}(i)=\mathrm{Var}_{\sigma\sim\mu}(\sigma(i))=\mu(\sigma(i)=1)\cdot\mu(\sigma(i)=0)$
for all $i\in[n]$. Plugging in the definition of the influence matrix $\Psi$ into Eq. 5 then we have for all $i,j\in[n]$,
$\mathrm{Cov}_{\mu}(i,j)=\mathrm{Var}_{\mu}(i)\cdot\Psi_{\mu}(i,j).$ (6)
Let $D=\mathrm{diag}(\mathrm{Var}_{\mu})$ denote the diagonal matrix of variances, and hence $D^{-1}$ is the diagonal matrix with $D^{-1}(i,i)=1/D(i,i)$. Then we have the matrix identity:
$\Psi_{\mu}=D^{-1}\mathrm{Cov}_{\mu}.$
Since $\mathrm{Cov}_{\mu}$ is symmetric positive semidefinite (see [HJ12, Chapter 7]), and $D$ (and hence $D^{-1}$ as well) is a diagonal matrix with positive diagonal entries (this follows from the definition of the set $T$ of free vertices in Section 1.4), then, by Lemma 2.1, we have that $\Psi_{\mu}$ has non-negative real eigenvalues. ∎
2.2 Bounding Spectral Radius of Influence Matrix
To upper bound the maximum eigenvalue of the influence matrix, most approaches use the maximum absolute row sum. Recall, the influence matrix $\Psi$ is defined to be a signed matrix, meaning the entries can be positive, negative, or zero. By considering the unsigned matrix formed by taking the absolute value of each entry of $\Psi$, and then considering the maximum row sum this yields an upper bound on the maximum eigenvalue of $\Psi$.
Lemma 2.3.
For any $n\times n$ influence matrix $\Psi$,
$\lambda_{\max}(\Psi)\leq\max_{i}\sum_{j}|\Psi(i,j)|.$
The below proof and statement of Lemma 2.3 hold for arbitrary matrices $M$ in place of the influence matrix $\Psi$ with the largest eigenvalue replaced by the spectral radius of $M$.
Proof.
Consider an $n\times n$ matrix $M$ with all real eigenvalues. Let $v$ be the eigenvector corresponding to the largest (in absolute value) eigenvalue $\lambda$ of $M$. Thus,
$Mv=\lambda v.$ (7)
Let $1\leq i\leq n$ be the position in $v$ where $|v_{i}|$ is maximized (note that $|v_{i}|>0$). We have
$|\lambda|\,|v_{i}|$ $=|(Mv)_{i}|$ by Eq. 7
$=\big{|}\sum_{j}M_{ij}v_{j}\big{|}$
$\leq\sum_{j}|M_{ij}|\,|v_{j}|$ by the triangle inequality
$\leq|v_{i}|\sum_{j}|M_{ij}|.$ since $v_{i}$ is the maximum.
Dividing both sides by $|v_{i}|$ yields the lemma. ∎
More generally, one can consider any matrix $p$-norm $\|\Psi\|_{p}$ for any $p\geq 1$. The absolute row sum in Lemma 2.3 corresponds to $p=\infty$, and the absolute column sum corresponds to $p=1$. Recall, the matrix $p$-norm is defined as follows:
$\|\Psi\|_{p}:=\max_{\mathbf{x}:\|\mathbf{x}\|_{p}=1}\|\Psi\mathbf{x}\|_{p},$
where $\mathbf{x}$ is an $n$-dimensional real-valued unit column vector, and for $\mathbf{y}$, which is an $n$-dimensional real-valued column vector, the vector $p$-norm of $\mathbf{y}$ is $\|\mathbf{y}\|_{p}=\left(\sum_{i}|\mathbf{y}_{i}|^{p}\right)^{1/p}$.
Lemma 2.4.
For any $n\times n$ influence matrix $\Psi$, for any $p\geq 1$,
$\lambda_{\max}(\Psi)\leq\|\Psi\|_{p}.$
Proof.
Let $v$ be the eigenvector corresponding to the largest (in absolute value) eigenvalue $\lambda$ of $M$. Thus,
$Mv=\lambda v.$
We have
$|\lambda|\,\|v\|_{p}$ $=\|\lambda v\|_{p}$
$=\big{\|}Mv\big{\|}_{p}$
$\leq\|M\|_{p}\|v\|_{p},$
and dividing both sides by $\|v\|_{p}$ completes the proof of the lemma. ∎
2.3 Correlation Matrix
As pointed out earlier, the influence matrix $\Psi$ is not necessarily symmetric. It is often convenient to work with a symmetric matrix, and in this section we introduce a symmetric version of the influence matrix.
Recall from Definition 1.1, the influence matrix $\Psi$ has $(u,v)$-entry:
$\Psi(u\rightarrow v):=\mu(\sigma(v)=1|\sigma(u)=1)-\mu(\sigma(v)=1|\sigma(u)=0).$
Also recall from Eq. 6 that the following holds for all $u,v\in V$:
$\Psi(v\rightarrow u)\cdot\mathrm{Var}_{\mu}(v)=\mathrm{Cov}_{\mu}(u,v)=\Psi(u\rightarrow v)\cdot\mathrm{Var}_{\mu}(u).$ (8)
Let $D=\mathrm{diag}(\mathrm{Var}_{\mu})$ be diagonal matrix with entries $\mathrm{Var}_{\mu}(v)=\mu(\sigma(v)=1)\cdot\mu(\sigma(v)=0)$. Equation (8) means that $D\Psi=\mathrm{Cov}_{\mu}$ is symmetric. Since $D\Psi$ is symmetric then $D^{1/2}\Psi D^{-1/2}$ is also symmetric, as it is obtained from $D\Psi$ by multiplying by the same diagonal matrix on the left and on the right. Now we can define the symmetric version of the influence matrix.
Definition 2.5.
Let the correlation matrix be defined as $D^{1/2}\Psi D^{-1/2}$ where $D=\mathrm{diag}(\mathrm{Var}_{\mu})$ is the diagonal matrix of variances, with entries $D(i,i)=\mathrm{Var}_{\mu}(i)=\mu(\sigma(i)=1)\cdot\mu(\sigma(i)=0)$.
An important feature of the correlation matrix is that the largest eigenvalue is the same as for the influence matrix.
Lemma 2.6.
For an $n\times n$ influence matrix $\Psi$, the correlation matrix $D^{1/2}\Psi D^{-1/2}$ has $(u,v)$-entries $\mathrm{Cov}_{\mu}(u,v)/\sqrt{\mathrm{Var}_{\mu}(u)\cdot\mathrm{Var}_{\mu}(v)}$, whose absolute value equals $\sqrt{\Psi(u\rightarrow v)\Psi(v\rightarrow u)}$. Moreover, the correlation matrix $D^{1/2}\Psi D^{-1/2}$ is similar to the influence matrix $\Psi$, and thus we have:
$\lambda_{\max}(\Psi)=\lambda_{\max}(D^{1/2}\Psi D^{-1/2}).$ (9)
Proof.
The $(u,v)$-entry in $D^{1/2}\Psi D^{-1/2}$ is
$\sqrt{\mathrm{Var}_{\mu}(u)}\cdot\Psi(u\rightarrow v)\cdot\frac{1}{\sqrt{\mathrm{Var}_{\mu}(v)}}=\sqrt{\mathrm{Var}_{\mu}(u)}\cdot\frac{\mathrm{Cov}_{\mu}(u,v)}{\mathrm{Var}_{\mu}(u)}\cdot\frac{1}{\sqrt{\mathrm{Var}_{\mu}(v)}}=\frac{\mathrm{Cov}_{\mu}(u,v)}{\sqrt{\mathrm{Var}_{\mu}(u)\cdot\mathrm{Var}_{\mu}(v)}}.$
Notice that the square of the $(u,v)$-entry is
$\frac{\left(\mathrm{Cov}_{\mu}(u,v)\right)^{2}}{\mathrm{Var}_{\mu}(u)\cdot\mathrm{Var}_{\mu}(v)}=\Psi(u\rightarrow v)\cdot\Psi(v\rightarrow u),$
establishing the lemma. ∎
2.4 Semidefinite Ordering
Let $D={\mathrm{diag}}(\mathrm{Var}_{\mu})$ be the diagonal matrix of variances, with entries $D(i,i)=\mathrm{Var}_{\mu}(i)=\mu(\sigma(i)=1)\cdot\mu(\sigma(i)=0)$. Bounding the largest eigenvalue of the influence matrix $\Psi$ by $\lambda$ is equivalent to the following statement:
$D^{1/2}\Psi D^{-1/2}\preceq\lambda I,$
which is equivalent to the following:
$D\Psi\preceq\lambda D.$
Recall the $n\times n$ covariance matrix $\mathrm{Cov}_{\mu}$ from Eq. 4, and $\mathrm{Cov}_{\mu}=D\Psi$. Therefore, we have that $\lambda_{\max}(\Psi)\leq 1+\eta$ is equivalent to the following semidefinite inequality:
$\mathrm{Cov}_{\mu}\preceq(1+\eta)D.$ (10)
Recall, condition Eq. 3 in $\eta$-spectral independence, we have shown that this condition (namely, $\lambda_{\max}(\Psi)\leq 1+\eta$) is equivalent to Eq. 10.
2.5 Modified Influence Matrix
Definition 2.7 (Modified Influence Matrix).
Let $G=(V,E)$ be a graph where $V=\{1,\dots,n\}$, and $\mu$ be a distribution on $\Omega$ where $\Omega\subset\{0,1\}^{V}$. Let $\widetilde{\Psi}$ be the following real-valued $n\times n$ matrix; we refer to $\widetilde{\Psi}$ as the modified influence matrix. For $1\leq i,j\leq n$, let
$\widetilde{\Psi}(i\rightarrow j)=\widetilde{\Psi}(i,j):=\mu\left(\sigma(j)=1\mid\sigma(i)=1\right)-\mu\left(\sigma(j)=1\right).$ (11)
Remark 2.8.
We remark that the modified influence matrix was first studied in [AASV21] and was called the correlation matrix. However, we define the correlation matrix in Definition 2.5 which matches the standard definition in statistics, as can be seen from Lemma 2.6.
We consider an analog of spectral independence for the modified influence matrix, which we refer to as weak spectral independence.
Definition 2.9 (Weak Spectral Independence).
For $\eta>0$, we say that $\mu$ is $\eta$-weak spectrally independent if for all pinnings $\tau\in\mathcal{P}$,
$\lambda_{\max}(\widetilde{\Psi}_{\tau})\leq 1+\eta.$ (12)
Recall, for $\eta$-spectral independence, $\lambda_{\max}(\Psi)\leq 1+\eta$ is equivalent to $\mathrm{Cov}_{\mu}\preceq(1+\eta)D$, see Eq. 10. Let $\widetilde{D}={\mathrm{diag}}(m_{\mu})$ be the diagonal matrix of means, with entries
$\widetilde{D}(i,i)=m_{\mu}(i)=\mathbb{E}_{\sigma\sim\mu}[\sigma(i)]=\mu(\sigma(i)=1).$
Analogous to Eq. 10, we have that $\lambda_{\max}(\widetilde{\Psi})\leq 1+\eta$ is equivalent to:
$\mathrm{Cov}_{\mu}\preceq(1+\eta)\widetilde{D}$. (13)
Remark 2.10.
Note that (10) implies (13), since for a random variable $X\in[0,1]$ we have $\mathrm{Var}(X)\leq\mathbb{E}(X)$. Consequently, $\eta$-spectral independence implies $\eta$-weak spectral independence. Moreover, under the assumption of $b$-marginal boundedness the notions of $\eta$-spectral independence and $\eta^{\prime}$-weak spectral independence are equivalent for some constants $\eta,\eta^{\prime}>0$; namely, $\eta^{\prime}$-weak spectral independence implies $((1+\eta^{\prime})/b-1)$-spectral independence.
Note that for the rapid mixing results presented in Section 1 we need $\eta$-spectral independence for a constant $\eta>0$. In particular, to prove Theorem 1.5 which establishes optimal relaxation time we need $\eta$-spectral independence for constant $\eta>0$. For Theorem 1.8 which establishes optimal mixing time, as noted above in Remark 2.10, under the assumption of $b$-marginal boundedness the notions of $\eta$-spectral independence and $\eta^{\prime}$-weak spectral independence are equivalent for some constants $\eta,\eta^{\prime}>0$.
In Section 7, for the application to the bases-exchange walk for matroids, we implicitly show that 0-weak spectral independence implies optimal relaxation time, see Remark 8.5. Moreover, in Section 8.4 we show that the log-concavity of the partition function is closely related to 0-weak spectral independence.
3 Markov Chain Fundamentals
To prove Theorem 1.5 we will bound the spectral gap of the Glauber dynamics. To do that, we consider a functional analysis formulation of the spectral gap. In particular, we will bound the decay rate of the variance of the chain with respect to the stationary distribution for an arbitrary real-valued function on the state space. This yields tight bounds on the spectral gap and relaxation time of the chain. Subsequently in Section 6.8, we will extend the notions presented in this section from variance to entropy, and present the main ideas for bounding the (modified) log-Sobolev constant which captures the decay rate of relative entropy; this yields optimal upper bounds on the mixing time, rather than the relaxation time.
3.1 Mixing Time and Total Variation Distance
For a pair of distributions $\mu$ and $\nu$ on a finite state space $\Omega$, their total variation distance is defined as:
$\|\mu-\nu\|_{\mathrm{TV}}=\frac{1}{2}\sum_{\sigma\in\Omega}|\mu(\sigma)-\nu(\sigma)|=\max_{S\subset\Omega}\mu(S)-\nu(S).$
For a Markov chain $(X_{t})$ with transition matrix $P$, state space $\Omega$, and unique stationary distribution $\pi$, its mixing time is the number of steps $T_{\mathrm{mix}}$ from the worst-case initial state $X_{0}\in\Omega$ to reach a distribution which is within total variation distance $\leq 1/4$ of $\pi$:
$T_{\mathrm{mix}}=\min\{t:\forall X_{0}\in\Omega,\|\mu-P^{t}(X_{0},\cdot)\|_{\mathrm{TV}}\leq 1/4.$
More generally, for any $\varepsilon>0$, let
$T_{\mathrm{mix}}(\varepsilon)=\min\{t:\forall X_{0}\in\Omega,\|\mu-P^{t}(X_{0},\cdot)\|_{\mathrm{TV}}\leq\varepsilon.$
We say the mixing time is optimal if $T_{\mathrm{mix}}=O(n\log{n})$ as this matches the lower bound established by Hayes and Sinclair [HS07] for any graph of constant maximum degree $\Delta$. Note an upper bound of $T_{\mathrm{mix}}=O(n\log{n})$ yields a more general bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log{n}\log(1/\varepsilon))$ via the above boosting result. However, the proofs of optimal mixing in these notes will yield a stronger upper bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$, and hence we will state this improved bound when relevant.
We will often establish a mixing time bound of $T_{\mathrm{mix}}=O(n\log{n})$, which as noted earlier is referred to as optimal mixing time. An optimal mixing time bound of $T_{\mathrm{mix}}=O(n\log{n})$ imples the following mixing time bound for any $\varepsilon>0$ (see [LP17, Section 4.5]):
$T_{\mathrm{mix}}(\varepsilon)\leq T_{\mathrm{mix}}(1/4)\times\lceil\log_{2}(1/\varepsilon)\rceil.$
In fact, many proof techniques, as is the case in this monograph, we obtain the stronger bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon)$.
Closely connected to variation distance is the coupling technique. For a pair of distributions $\mu,\nu$, both of which are on the same state space $\Omega$, a coupling $\omega$ of $\mu$ and $\nu$ is a distribution on $\Omega\times\Omega$ such that the projections on each coordinate follow the respective distributions. More formally, $\omega$ is a valid coupling of $\mu$ and $\nu$ if the following hold:
  • For all $\tau\in\Omega$,
    $\sum_{\sigma\in\Omega}\omega(\sigma,\tau)=\nu(\tau),$
  • For all $\sigma\in\Omega$,
    $\sum_{\tau\in\Omega}\omega(\sigma,\tau)=\mu(\sigma).$
The goal is to design a coupling $\omega$ which minimizes the disagreement probability which is the probability of $\sigma\neq\tau$ where $(\sigma,\tau)$ is a sample from the coupled distribution $\omega$. This disagreement probability provides an upper bound on the total variation distance of the coupled distributions $\mu$ and $\nu$, and there always exists a coupling where the disagreement probability equals the total variation distance; this is summarized in the following lemma, often known as the coupling lemma (see, e.g., [Ald83, Lemma 3.6]).
Lemma 3.1.
For a pair of distributions $\mu$ and $\nu$ on a finite state space $\Omega$, let $\mathcal{P}(\mu,\nu)$ denote the family of all valid couplings of $\mu$ and $\nu$.
  • For any $\omega\in\mathcal{P}(\mu,\nu)$, for a sample $(\sigma,\tau)\sim\omega$,
    $\|\mu-\nu\|_{\mathrm{TV}}\leq\mathrm{Pr}\left(\sigma\neq\tau\right).$
  • There exists a coupling $\omega\in\mathcal{P}(\mu,\nu)$ where:
    $\|\mu-\nu\|_{\mathrm{TV}}=\mathrm{Pr}\left(\sigma\neq\tau\right).$
For a more comprehensive introduction to coupling methods in the MCMC setting see [LP17, Jer03].
3.2 Reversibility
We say an ergodic Markov chain is defined by $(\Omega,P,\mu)$ to refer to $\Omega$ as the (finite) state space, $P$ is the $N\times N$ transition matrix where $N=|\Omega|$, and $\mu$ is the stationary distribution (which is unique since we will only consider ergodic Markov chains in this text). For some of the upcoming algebraic statements in terms of eigenvalues we will need to restrict attention to reversible chains; these are Markov chains whose transition matrix $P$ and stationary distribution $\mu$ satisfy the following condition known as the reversibility condition, namely, for all pairs $\sigma,\sigma^{\prime}\in\Omega$,
$\mu(\sigma)P(\sigma,\sigma^{\prime})=\mu(\sigma^{\prime})P(\sigma^{\prime},\sigma).$ (14)
It is straightforward to verify the reversibility condition Eq. 14 for the Glauber dynamics since $P(\sigma,\sigma^{\prime})>0$ iff $\sigma=\sigma^{\prime}$ or $\sigma\oplus\sigma^{\prime}=\{v\}$ for some $v\in V$ (i.e., the Glauber dynamics changes at most one vertex in each transition). For a discussion of non-reversible chains, we refer the interested reader to Montenegro and Tetali’s monograph [MT06].
3.3 Dirichlet Form and Variance
Consider a function $f:\Omega\rightarrow\mathbb{R}$. The expectation of $f$ with respect to the stationary distribution $\mu$ is denoted as follows:
$\mu(f):=\mathbb{E}_{\mu}[f]:=\sum_{\eta\in\Omega}\mu(\eta)f(\eta).$
We will also study the variance of $f$ with respect to $\mu$ which has the following convenient, equivalent formulations that can be easily verified by simple algebra:
$\mathrm{Var}_{\mu}(f)$ $:=\mu\left(\left(f-\mu(f)\right)^{2}\right)$
$=\sum_{\sigma\in\Omega}\mu(\sigma)\left(f(\sigma)-\mu(f)\right)^{2}$
$=\frac{1}{2}\sum_{\sigma,\eta\in\Omega}\mu(\sigma)\mu(\eta)(f(\sigma)-f(\eta))^{2},$
see [Jer03, (5.4-5.5)] for a proof of the last identity. Notice that the last formulation of variance is a sum over all pairs of states of the difference in their functional value (squared to measure the absolute value).
It will be convenient later to use inner product and $2$-norm to represent the variance functional. For two functions $f,g:\Omega\rightarrow\mathbb{R}$, the inner product of $f$ and $g$ with respect to $\mu$ is defined as
$\left\langle f,g\right\rangle_{\mu}:=\sum_{\eta\in\Omega}\mu(\eta)f(\eta)g(\eta)=\mu(fg),$
and the $2$-norm of $f$ with respect to $\mu$ is defined as
$\left\lVert f\right\rVert_{2,\mu}=\sqrt{\left\langle f,f\right\rangle_{\mu}}.$
Thus, we can represent the variance of $f$ as
$\mathrm{Var}_{\mu}(f)=\left\lVert f-\mu(f)\right\rVert_{2,\mu}^{2}.$
The variance functional is closely related to the $\chi^{2}$-divergence. Specifically, for any distribution $\nu$ over $\Omega$, the $\chi^{2}$-divergence of $\nu$ from $\mu$ is defined as
$D_{\chi^{2}}(\nu\,\|\,\mu):=\sum_{\eta\in\Omega}\mu(\eta)\left(\frac{\nu(\eta)}{\mu(\eta)}-1\right)^{2}=\sum_{\eta\in\Omega}\frac{\left(\nu(\eta)-\mu(\eta)\right)^{2}}{\mu(\eta)}.$
Observe that $D_{\chi^{2}}(\nu\,\|\,\mu)\geq 0$, with an equality iff $\nu=\mu$. If we let $f=\nu/\mu$ be the relative density of $\nu$ with respect to $\mu$, i.e., $f(\eta)=\nu(\eta)/\mu(\eta)$ for each $\eta\in\Omega$, then it holds $\mu(f)=1$ and $\mathrm{Var}_{\mu}(f)=D_{\chi^{2}}(\nu\,\|\,\mu)$.
In contrast to the variance which measures the “global variation” of the function $f$, the Dirichlet form is a measure of the “local variation” for a specified Markov chain. The terms global vs. local are with respect to the graph $(\Omega,P)$. The Dirichlet form measures the difference in the functional value over pairs of adjacent states with respect to the graph $(\Omega,P)$ as formulated in the following definition.
Definition 3.2 (Dirichlet Form).
For a Markov chain on state space $\Omega$ with transition matrix $P$, and for a function $f:\Omega\rightarrow\mathbb{R}$, define the Dirichlet form as follows:
$\mathcal{E}_{P}(f):=\frac{1}{2}\sum_{\sigma,\eta\in\Omega}\mu(\sigma)P(\sigma,\eta)(f(\sigma)-f(\eta))^{2}.$
Equivalently, we can write the Dirichlet form as
$\mathcal{E}_{P}(f)=\left\langle f,(I-P)f\right\rangle_{\mu},$
where $I$ is the identity matrix, and (IP)f(I-P)f is a function from $\Omega$ to $\mathbb{R}$ obtained as a matrix-vector product of IPI-P and $f$ (viewed as a column vector).
3.4 Spectral Gap
For an ergodic, reversible Markov chain $(\Omega,P,\mu)$, recall that $P$ denotes the transition matrix and $\mu$ is the stationary distribution which means that $\mu$ is a row eigenvector of $P$ with eigenvalue $1$. Since the Markov chain is reversible then the eigenvalues of $P$ are real-valued, and hence we denote the eigenvalues of $P$ as follows:
$1>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{N}>-1.$
The fact that $\lambda_{N}>-1$ follows from the aperiodicity of an ergodic Markov chain. The spectral gap is defined as $\gamma=1-\lambda_{2}$, and the absolute spectral gap is $1-\lambda_{*}$ where $\lambda_{*}=\max\{\lambda_{2},|\lambda_{N}|\}$ is the second largest eigenvalue in absolute value.
Dyer, Greenhill, and Ullrich [DGU14] proved that the heat-bath block dynamics is positive semidefinite (PSD), that is, all of the eigenvalues of $P$ are non-negative. The heat-bath block dynamics is defined by a collection of blocks $B_{1},\dots,B_{L}$ where each $B_{i}\subset V$ and $\cup_{i}B_{i}=V$; the steps $X_{t}\rightarrow X_{t+1}$ of the block dynamics operate by first choosing a random $B_{i}$ and then choosing the configuration $X_{t+1}(B_{i})$ on $B_{i}$ from the Gibbs distribution conditional on $X_{t}(V\setminus B_{i})$ (the configuration on $V\setminus B_{i}$ stays the same). The transition matrix of heat-bath update for one block is positive semidefinite (since it is a block diagonal matrix with positive semidefinite blocks of rank 1). The positive semidefinitness now follows since the heat bath chain is a mixture of block updates. The Glauber dynamics is the special case of the heat-bath block dynamics where the blocks are single vertices. Hence, for the Glauber dynamics, the absolute spectral gap is the same as the spectral gap.
The absolute spectral gap measures the decay rate of variance whereas the spectral gap is potentially off by a factor of $1/2$ (though this distinction is irrelevant for the case of the Glauber dynamics). The spectral gap is convenient to work with due to the following formulation known as a Poincaré inequality (see [LP17, Lemma 13.7] for a proof of the equivalence).
Lemma 3.3 (Spectral gap).
For a reversible Markov chain defined by $(\Omega,P,\mu)$, the spectral gap $\gamma$ is the largest constant such that the following inequality holds for any function $f:\Omega\rightarrow\mathbb{R}$,
$\gamma\cdot\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P}(f)\,.$ (15)
For non-reversible Markov chains, we can define $\gamma$ as in Eq. 15 which is referred to as the Poincaré constant (see [Jer95, Remark 5.3]); however it is not necessarily equivalent to the spectral gap $1-\lambda_{2}$ in the case of non-reversible chains.
Recall, the relaxation time is defined as the inverse of the absolute spectral gap:
$T_{\mathrm{relax}}:=(1-\lambda_{*})^{-1}.$
Let $P_{\textsc{gd}}$ denote the transition matrix of the Glauber dynamics. As noted above, for the Glauber dynamics the transition matrix $P_{\textsc{gd}}$ is PSD and hence the absolute spectral gap is the same as the spectral gap, and thus we have:
$T_{\mathrm{relax}}(P_{\textsc{gd}})\leq\gamma^{-1}.$ (16)
3.5 Approximate Tensorization of Variance
A useful tool for bounding the spectral gap of the Glauber dynamics is the following property of the Gibbs distribution known as approximate tensorization of variance.
Definition 3.4.
For a graph $G=(V,E)$, let $\mu$ be a probability distribution on $\Omega\subseteq\{0,1\}^{V}$. For a constant $C\geq 1$, we say that $\mu$ satisfies $C$-approximate tensorization of variance if the following holds for all $f:\Omega\rightarrow\mathbb{R}$,
$\mathrm{Var}_{\mu}(f)\leq C\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)].$ (17)
It is important to stress that approximate tensorization of variance is a property of the Gibbs distribution and does not explicitly involve the Glauber dynamics, but as we will see shortly it is equivalent to bounding the spectral gap. When $\mu$ is a product distribution, i.e., $\mu=\otimes_{v\in V}\mu(v)$, then Eq. 17 holds with $C=1$, which is the optimal constant (it is necessary that $C\geq 1$, as can be seen by taking functions depending only on a single vertex).
Let us decipher the right-hand side of Eq. 17 before proceeding: for the term $\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]$, we first sample a configuration $\tau$ on $V\setminus\{v\}$ where $\tau$ is drawn from the projection of $\mu$ on $V\setminus\{v\}$, then we consider the variance of the function $f$ with respect to the marginal distribution of $\mu$ on $v$ conditional on $\tau$. More formally, for a configuration $\tau\in\{0,1\}^{V\setminus\{v\}}$ on all vertices except $v$, let $\mu(\tau)$ denote the marginal probability of $\tau$ in the Gibbs distribution $\mu$, i.e., $\mu(\tau)=\sum_{\sigma\in\Omega:\sigma(V\setminus\{v\})=\tau}\mu(\sigma)$. Conversely, let $\mu_{\tau}$ denote the Gibbs distribution $\mu$ conditional on $\tau$, and thus for $i\in\{0,1\}$,
$\mu_{\tau}(\eta(v)=i)=\mu(\eta(v)=i\mid\eta(V\setminus\{v\})=\tau).$
Then we have the following:
$\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]=\sum_{\tau\in\{0,1\}^{V\setminus\{v\}}}\mu(\tau)\mu_{\tau}(\eta(v)=0)\mu_{\tau}(\eta(v)=1)\left(f(\tau_{v,0})-f(\tau_{v,1})\right)^{2},$ (18)
where, for $i\in\{0,1\}$, $\tau_{v,i}$ is the extension of $\tau$ to an assignment on all vertices where for $w\in V$:
$\tau_{v,i}(w):=\begin{cases}i,&\mbox{if }w=v;\\ \tau(w),&\mbox{if }w\neq v.\end{cases}$
For a configuration $\sigma\in\Omega$ on all vertices $V$, for $v\in V$, let $\sigma_{v}=\sigma_{v,1-\sigma(v)}$ denote the configuration which flips the spin at $v$ and keeps all other spins the same. Let $P_{\textsc{gd}}$ denote the transition matrix for the Glauber dynamics. Using the observation that $\mu^{\sigma(V\setminus\{v\})}(\eta(v)=1-\sigma(v))=nP_{\textsc{gd}}(\sigma,\sigma_{v})$ and recalling the definition of the Dirichlet form (see Definition 3.2), we can rewrite Eq. 18 as follows:
$\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]$ $=\frac{n}{2}\sum_{v\in V}\sum_{\sigma\in\Omega}\mu(\sigma)P_{\textsc{gd}}(\sigma,\sigma_{v})\left(f(\sigma)-f(\sigma_{v})\right)^{2}$ (19)
$=n\mathcal{E}_{P_{\textsc{gd}}}(f).$ (20)
Therefore,
$\mathrm{Var}_{\mu}(f)\leq C\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]\quad\iff\quad\frac{1}{Cn}\cdot\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P}(f),$
and hence $C$-approximate tensorization of variance is equivalent to the spectral gap satisfying $\gamma\geq 1/(Cn)$.
Corollary 3.5.
A distribution $\mu\in\{0,1\}^{V}$ satisfies $C$-approximate tensorization of variance iff the Glauber dynamics for $\mu$ has spectral gap $\gamma\geq 1/(Cn)$ (or equivalently, the Glauber dynamics has relaxation time $T_{\mathrm{relax}}\leq Cn$).
The analog of approximate tensorization of variance with entropy in place of variance is presented in Section 6.8, see Definition 6.8. Whereas approximate tensorization of variance implies optimal bounds of $O(n)$ time on the relaxation time, the stronger condition of approximate tensorization of entropy implies optimal bounds of $O(n\log{n})$ time on the mixing time. We refer the interested reader to Section 6.8 for further details.
3.6 Preliminaries
Here we present a functional characterization of spectral independence, utilizing the variance functional.
Definition 3.6.
We say that a measure $\mu$ on $\{0,1\}^{n}$ satisfies approximate subadditivity of variance with constant $C$ if for every $f:\Omega\to\mathbb{R}$ we have
$\sum_{i=1}^{n}\mathrm{Var}_{\mu}[\mathbb{E}[f|X_{i}]]\leq C\cdot\mathrm{Var}_{\mu}[f].$ (21)
Lemma 3.7.
Measure $\mu$ satisfies $\lambda_{\max}(\Psi_{\mu})\leq C$ if and only if $\mu$ satisfies approximate subadditivity of variance with constant $C$.
Proof.
Without loss of generality, we can assume $\mathbb{E}_{\mu}[f]=0$. Let $p_{i}=\mathrm{Pr}(X_{i}=1)\in(0,1)$ for $i\in[n]$. Let $a_{i}=\mathbb{E}[f|X_{i}=1]$ and note that $\mathbb{E}[f|X_{i}=0]=a_{i}p_{i}/(1-p_{i})$, since $\mathbb{E}_{\mu}[f]=0$. In terms of $p_{i}$ and $a_{i}$’s the LHS of (21) is
$\sum_{i=1}^{n}a_{i}^{2}\frac{p_{i}}{1-p_{i}}.$ (22)
We only need to verify (21) for $f$ that minimizes $\mathrm{Var}_{\mu}[f]$ subject to $\mathbb{E}[f|X_{i}=1]=a_{i}$ for $i\in[n]$ and $\mathbb{E}_{\mu}[f]=0$. From the method of Lagrange multipliers it follows that such $f$ is linear, that is, for some $c_{0},c_{1},\dots,c_{n}$ we have
$f=c_{0}+\sum_{i=1}^{n}c_{i}X_{i}.$
Let $c=(c_{1},\dots,c_{n})$. In terms of $c_{i}$’s the main term on the RHS of (21) is
$\mathrm{Var}_{\mu}[f]=c^{T}\mathrm{Cov}_{\mu}c.$
The $c_{i}$’s, $a_{i}$’s, and $p_{i}$’s satisfy the following equation (for $i\in[n]$):
$(\mathrm{Cov}_{\mu}c)_{i}=\mathbb{E}[fX_{i}]-\mathbb{E}[f]\mathbb{E}[X_{i}]=\mathrm{Pr}[X_{i}=1]\mathbb{E}[f|X_{i}=1]-0=p_{i}a_{i}.$ (23)
Let $D=\mathrm{diag}(\mathrm{Var}(X_{i}))$ be the variance matrix. We have $D_{ii}=p_{i}(1-p_{i})$ for $i\in[n]$ and recall that $C$-spectral independence means $\mathrm{Cov}_{\mu}\preceq CD$, see Eq. 10. Thus we can express the LHS of (21) as $c^{T}\mathrm{Cov}_{\mu}D^{-1}\mathrm{Cov}_{\mu}c$ and (21) holding for every $f$ is equivalent to the following
$\mathrm{Cov}_{\mu}D^{-1}\mathrm{Cov}_{\mu}\preceq C\,\mathrm{Cov}_{\mu}.$ (24)
If $\mu$ is $C$ spectrally independent then we have $D^{-1}\preceq C\mathrm{Cov}_{\mu}^{-1}$ and by multiplying by $\mathrm{Cov}_{\mu}$ on both left and right we obtain (24) and hence $C$-approximate subadditivity of variance (technically one should use the pseudo-inverse for $\mathrm{Cov}_{\mu}$ since it is not guaranteed to be invertible; this does not change the proof).
If $\mu$ satisfies $C$-approximate subadditivity of entropy then we have (24); multiplying both sides by $\mathrm{Cov}_{\mu}^{-1}$ (technically we are using the pseudo-inverse for $\mathrm{Cov}_{\mu}$) we obtain $D^{-1}\preceq C\,\mathrm{Cov}_{\mu}^{-1}$, which implies $\mathrm{Cov}_{\mu}\preceq C\cdot D$ and hence, by Eq. 10, $\lambda_{\max}(\Psi_{\mu})\leq C$. ∎
3.7 Mixing Time Bounds via Spectral Gap
We now show how the spectral gap implies an upper bound on the mixing time. For the Glauber dynamics $(X_{t})$, let $P_{\textsc{gd}}$ denote the transition matrix, let $\nu_{t}$ denote the distribution of the random variable $X_{t}$ at time $t$, and let $\mu$ denote the stationary distribution. Let
$\mu^{*}=\min_{x\in\Omega}\mu(x).$
We will show the following bound:
$T_{\mathrm{mix}}(P_{\textsc{gd}})\leq\frac{1}{2\gamma}\ln(4/\mu^{*}).$ (25)
We will also establish a more general bound which accounts for a so-called warm start when the initial distribution $\nu_{0}$ is somewhat close to the stationary distribution $\mu$.
Let $f_{t}=\nu_{t}/\mu$ denote the relative density at time $t$, i.e., $f_{t}(x)=\nu_{t}(x)/\mu(x)$ for each $x\in\Omega$. Notice that $\mu(f_{t})=1$. Recall the $\chi^{2}$-divergence from stationarity defined as:
$D_{\chi^{2}}(\nu_{t}\,\|\,\mu)=\mathrm{Var}_{\mu}(f_{t})=\left\|f_{t}-1\right\|_{2,\mu}^{2}=\sum_{x\in\Omega}\mu(x)\left(\frac{\nu_{t}(x)}{\mu(x)}-1\right)^{2}.$
Meanwhile, the $L^{1}$-distance from stationarity is related to the total variation distance:
$\left\|f_{t}-1\right\|_{1,\mu}=\sum_{x\in\Omega}|\nu_{t}(x)-\mu(x)|=2\left\|\nu_{t}-\mu\right\|_{\mathrm{TV}}.$
We can bound the convergence time to the stationary distribution $\mu$ as a function of the initial distance. Let
$M:=D_{\chi^{2}}(\nu_{0}\,\|\,\mu)$ (26)
denote the initial $\chi^{2}$-divergence from the stationary distribution. Note that $M$ is typically exponentially large in $n$:
$M=\left(\sum_{x}\frac{\nu_{0}(x)^{2}}{\mu(x)}\right)-1\leq\frac{1}{\mu^{*}}\sum_{x}\nu_{t}(x)^{2}\leq\frac{1}{\mu^{*}}.$ (27)
In contrast, when $M=O(1)$ (e.g., if $\nu_{0}(x)\leq 2\mu(x)$ for all $x\in\Omega$) then we refer to $\nu_{0}$ as a warm start. We will prove the following bound on the distance to stationarity of the Glauber dynamics after $t$ steps as a function of the initial distance $M$:
$4\left\|\nu_{t}-\mu\right\|_{\mathrm{TV}}^{2}\leq D_{\chi^{2}}(\nu_{t}\,\|\,\mu)\leq M\exp(-2\gamma t).$ (28)
The above bound holds for the Glauber dynamics, and more generally holds for any reversible chain whose transition matrix is positive semidefinite (consequently, the absolute spectral gap is the same as the spectral gap); we will prove it below. Plugging in the worst case bound for the initial distance $M$ from Eq. 27 into Eq. 28 yields the upper bound on the mixing time stated earlier in Eq. 25.
For general Markov chains (not necessarily reversible), defined by a transition matrix $P$, recall $\gamma$ is defined in Eq. 15. Note that for non-reversible chains the eigenvalues of the transition matrix are not necessarily real. For general chains (not necessarily reversible) a similar result as in Eq. 28 holds with $\gamma/2$ in place of $2\gamma$ for the lazy version $P_{zz}:=\frac{1}{2}(P+I)$ (which transitions as in the original chain $P$ with probability $1/2$ and stays in the same state with probability $1/2$), see [Mih89] and [Jer03, Theorem 5.6].
Notice that if we have a warm start for the initial distance then the mixing time is on the order of the inverse spectral gap. Consequently, in many approximate counting algorithms, such as the simulated annealing algorithm of [ŠVV09, Hub15, Kol18, HK25], the running time avoids the extra $O(\log(1/\mu^{*}))$ factor.
The main task is to prove Eq. 28; then Eq. 25 follows as a corollary as noted above by plugging in the worst-case bound on $M$ from Eq. 27. We refer the interested reader to [Jer03, Chapter 5] for an alternative (and somewhat simpler) proof of Eq. 28 for the lazy chain $P_{zz}$. The below proof is stated for the Glauber dynamics but holds for any PSD transition matrix.
Proof of Eq. 28.
[…]
$\mathrm{Var}(P_{\textsc{gd}}f)\leq(1-\gamma)^{2}\mathrm{Var}(f),\quad\forall f:\Omega\to\mathbb{R};$
$\quad\iff\quad$ $D_{\chi^{2}}(\nu P_{\textsc{gd}}\,\|\,\mu)\leq(1-\gamma)^{2}D_{\chi^{2}}(\nu\,\|\,\mu),\quad\text{$\forall$ distribution $\nu$ on $\Omega$}.$ (29)
We will establish Eq. 29 later in the proof. Note, for the analog of Eq. 29 for the lazy version of general chains (including non-reversible chains), see [Mih89] and, e.g., [Jer03, Corollary 5.7].
To complete the proof of Eq. 28, assuming Eq. 29 holds, we will begin by showing the first inequality in Eq. 28. Namely, by Cauchy–Schwarz inequality the $L^{1}$-distance is upper bounded by the $L^{2}$-distance (see also [MT06, (7.1)]):
$2\left\|\nu_{t}-\mu\right\|_{\mathrm{TV}}=\left\|f_{t}-1\right\|_{1,\mu}=\mu\left(|f_{t}-1|\right)\leq\sqrt{\mu\left((f_{t}-1)^{2}\right)}=\left\|f_{t}-1\right\|_{2,\mu}=\sqrt{D_{\chi^{2}}(\nu_{t}\,\|\,\mu)}.$
For the second inequality in Eq. 28, we deduce from Eq. 29 that
$D_{\chi^{2}}(\nu_{t}\,\|\,\mu)=D_{\chi^{2}}(\nu_{0}P^{t}\,\|\,\mu)\leq(1-\gamma)^{2t}D_{\chi^{2}}(\nu_{0}\,\|\,\mu)\leq M\exp(-2\gamma t),$
where $M$ is the $\chi^{2}$-divergence of the initial distribution $\nu_{0}$ to the stationary distribution $\mu$, see Eq. 26.
It remains to prove Eq. 29. We claim the following holds:
$\mathrm{Var}_{\mu}(P_{\textsc{gd}}f)=\mathrm{Var}_{\mu}(f)-\mathcal{E}_{P_{\textsc{gd}}^{2}}(f).$ (30)
Without loss of generality, we assume $\mu(f)=0$. We observe that
$\mathrm{Var}_{\mu}(f)=\left\lVert f\right\rVert_{2,\mu}^{2}=\langle f,f\rangle_{\mu}$
and
$\mathrm{Var}_{\mu}(P_{\textsc{gd}}f)=\left\lVert P_{\textsc{gd}}f\right\rVert_{2,\mu}^{2}=\langle P_{\textsc{gd}}f,P_{\textsc{gd}}f\rangle_{\mu}=\langle f,P_{\textsc{gd}}^{2}f\rangle_{\mu},$
where the last equality follows from the reversibility of $P_{\textsc{gd}}$. Therefore, we deduce that
$\mathrm{Var}_{\mu}(f)-\mathrm{Var}_{\mu}(P_{\textsc{gd}}f)=\langle f,(I-P_{\textsc{gd}}^{2})f\rangle_{\mu}=\mathcal{E}_{P_{\textsc{gd}}^{2}}(f).$
By the Poincaré inequality Eq. 15 $\gamma\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P_{\textsc{gd}}}(f)$, we have that $(1-(1-\gamma)^{2})\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P_{\textsc{gd}}^{2}}(f)$ since every eigenvalue of $P_{\textsc{gd}}^{2}$ is the square of an eigenvalue of $P_{\textsc{gd}}$, and the Glauber dynamics $P_{\textsc{gd}}$ is positive semidefinite. From Eq. 30, we then obtain Eq. 29 that
$\mathrm{Var}(P_{\textsc{gd}}f)\leq(1-\gamma)^{2}\mathrm{Var}(f),$
completing the proof. ∎
We showed in Eq. 25 how a lower bound on the spectral gap of the Glauber dynamics (which is equivalent to an upper bound on the relaxation time) implies an upper bound on the mixing time with an extra $O(\log(1/\mu^{*}))$ factor. In Remark 9.30 in Section 9.5 we discuss the reverse connection: namely, an optimal upper bound on the mixing time of the Glauber dynamics of the form $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$ for all $\varepsilon>0$ implies an optimal upper bound on the relaxation time of $T_{\mathrm{relax}}=O(n)$.
4 Rapid Mixing from Spectral Independence
The main result of this section is Theorem 1.4, which states that if $\mu$ is $\eta$-spectrally independent for a constant $\eta$, then the Glauber dynamics has polynomial mixing time.
4.1 Pinnings
Recall that a pinning is a fixed assignment of spins to a subset of vertices. For $S\subset V$ and a pinning $\tau:S\rightarrow\{0,1\}$, denote the assignments consistent with $\tau$ as $\Omega_{\tau}:=\{\sigma\in\Omega:\sigma(S)=\tau(S)\}$.
We will partition the pinnings based on the number of pinned vertices. For $0\leq k\leq n$, denote the set of valid pinnings of size $k$ by
$\mathcal{P}_{k}:=\{\tau\in\{0,1\}^{S}:\mbox{ for some }S\subseteq V,\mbox{ such that }|S|=k,\Omega_{\tau}\neq\emptyset\}.$
Recall, the set of valid configurations is denoted by
$\Omega=\{\sigma\in\{0,1\}^{V}:\mu(\sigma)>0\}.$
Observe that $\Omega=\mathcal{P}_{n}$, that is the set of valid configurations is the same as the set of pinnings on all $n$ vertices (this follows from the observation that for $\tau\in\Omega$, $\Omega_{\tau}=\{\tau\}\neq\emptyset$ and hence $\tau\in\mathcal{P}_{n}$).
Finally, denote the collection of all (valid) pinnings by:
$\mathcal{P}=\bigcup_{k=0}^{n}\mathcal{P}_{k}.$
For a pinning $\tau\in\mathcal{P}$, let $\mu_{\tau}$ denote the conditional Gibbs distribution (that is, the distribution $\mu$ conditional on the fixed assignment $\tau$ on $S$).
4.2 Rapid Mixing Theorem
We will bound the mixing time of the Glauber dynamics by considering the spectral gap. We will prove the following result of Anari, Liu, and Oveis Gharan [ALO20]. Namely, that if the Gibbs distribution $\mu$ is $\eta$-spectrally independent (note, Definition 1.2 requires it for all pinnings $\tau\in\mathcal{P}$) then the spectral gap satisfies:
$\gamma(\mathrm{P_{\textsc{gd}}})\geq\frac{C(\eta)}{n^{1+\eta}},$ (31)
for a constant $C(\eta)>0$. By Eq. 16 this implies relaxation time $O(n^{1+\eta})$ and mixing time $O(n^{2+\eta})$, and hence this proves Theorem 1.4.
4.3 Local to Global: Random Walk Theorem of Alev-Lau
The key technical tool in the proof of Eq. 31 is the following local-to-global theorem of Alev and Lau [AL20], which improved upon Kaufman and Oppenheim [KO20]. Our input graph $G=(V,E)$ has $n=|V|$ vertices and each vertex has two possible spin assignments 0 or $1$. The state space of the local walk are the $2n$ (vertex,spin) pairs $(v_{i},s_{i})\in V\times\{0,1\}$. The transition probabilities are defined based on their pairwise joint probabilities. The stationary distribution of the local walk corresponds to the marginal probability $\mu(\sigma(v_{i})=s_{i})$ of that particular vertex $v_{i}$ having the specified spin $s_{i}$.
The matrix $Q$ is a real-valued $2n\times 2n$ matrix. For $1\leq i\neq j\leq n$, $s_{i},s_{j}\in\{0,1\}$, let
$Q((i,s_{i}),(j,s_{j}))=\frac{1}{n-1}\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}\mid\sigma(i)=s_{i}],$
and when $i=j$, then for $s,s^{\prime}\in\{0,1\}$, let $Q((i,s),(i,s^{\prime}))=0$. Moreover, for $0\leq k\leq n-2$, for a subset $S\subset V$ where $|S|=k$, for a pinning $\tau:S\rightarrow\{0,1\}$ where $\tau\in\mathcal{P}$, for $i,j\notin S$ where $i\neq j$, and $s_{i},s_{j}\in\{0,1\}$, let
$Q_{\tau}((i,s_{i}),(j,s_{j}))$ $=$ $\frac{1}{n-k-1}\mathrm{Pr}_{\sigma\sim\mu}(\sigma(j)=s_{j}\mid\sigma(i)=s_{i},\sigma(S)=\tau)$ (32)
$=$ $\frac{1}{n-k-1}\mathrm{Pr}_{\sigma\sim\mu_{\tau}}(\sigma(j)=s_{j}\mid\sigma(i)=s_{i}),$
and when $i=j$ we set the entries to 0 as before. Notice that $Q=Q_{\emptyset}$, that is the local walk $Q$ corresponds to the empty pinning. In the definition of the local walk $Q$ and the generalization in Eq. 32, we restrict attention to pairs $(i,s_{i})$ where $\mu_{\tau}(\sigma(i)=s_{i})>0$.
The Markov chain $Q$ is called the “local walk” because it captures local information of the Gibbs distribution, namely the marginal probabilities of (vertex,spin) pairs. In contrast, the Glauber dynamics is referred to as a “global walk” because it captures the probability of a configuration on the entire graph.
The upcoming Random Walk Theorem of Alev and Lau [AL20] is referred to as a local-to-global theorem because it relates the behavior of local chains $Q$ to the Glauber dynamics which is a global chain. To do this we need to consider the local walk on each level. Namely, for $k$ where $0\leq k\leq n-2$, for $\tau\in\mathcal{P}_{k}$, consider the local walk $Q_{\tau}$. Let $\gamma_{k}$ be the minimum spectral gap for a local walk $Q_{\tau}$ where $\tau\in\mathcal{P}_{k}$.
We can now formally state the Random Walk Theorem of Alev and Lau [AL20] (which improves upon earlier work of [KO20]).
Theorem 4.1 (Random Walk Theorem [AL20]).
$\gamma(P_{\mathrm{Glauber}})\geq\frac{1}{n}\prod_{k=0}^{n-2}\gamma_{k},$ (33)
where
$\gamma_{k}=\min_{\tau\in\mathcal{P}_{k}}\gamma(Q_{\tau})$
is the spectral gap for the local walk with a worst-case pinning of $k$ vertices.
We will present the proof of the Random Walk Theorem in Section 5.
4.4 Local Walk Connection to Influence Matrix
We begin by bounding $\lambda_{2}$ of the local walk $Q_{\tau}$ in terms of the influence matrix $\Psi_{\tau}$ for the spectral independence technique. In fact, we can relate the entire spectrum of $Q_{\tau}$ with the spectrum of $\Psi_{\tau}$ in the following manner.
Let us first extend the definition of the influence matrix in Definition 1.1 to a fixed pinning. For a subset $S\subset V$ where $|S|=k$, for a pinning $\tau:S\rightarrow\{0,1\}$ where $\tau\in\mathcal{P}$, recall (see equation (2)) the influence matrix $\Psi_{\tau}$ is defined as:
$\Psi_{\tau}(i\rightarrow j):=\mu\left[\sigma(j)=1\mid\sigma(i)=1,\sigma(S)=\tau\right]-\mu\left[\sigma(j)=1\mid\sigma(i)=0,\sigma(S)=\tau\right].$
Lemma 4.2.
For a pinning $\tau\in\mathcal{P}_{k}$,
$\lambda_{2}(Q_{\tau})=\frac{1}{n-k-1}\left(\lambda_{\mathrm{max}}(\Psi_{\tau})-1\right).$ (34)
Moreover,
$\mbox{spectrum}(Q_{\tau})=\mbox{spectrum}\left(\frac{1}{n-k-1}\left(\Psi_{\tau}-I\right)\right)\cup\{1\}\cup\left\{n-k-1\mbox{ copies of }\frac{-1}{n-k-1}\right\}.$ (35)
As a consequence of the above lemma we have that $\gamma_{k}\geq 1-\eta/(n-k-1)$, see Section 4.5 for more details.
Proof of Lemma 4.2.
We will prove the lemma for the case without a pinning. Recall, $\mu$ denotes the Gibbs distribution.
Consider the local random walk $Q$. Notice that this walk has an $n$-partite structure, namely, for every $v_{i}\in V$, the states $(v_{i},0)$ and $(v_{i},1)$ are an independent set, and each of these $n$ independent sets are fully connected by a (weighted) bipartite clique. This $n$-partite structure yields eigenvalues 1/(n1)-1/(n-1) for $Q$. The stationary distribution for $Q$ is an eigenvector with eigenvalue $1$ (note that when $n\geq 3$ then the chain is clearly aperiodic and hence ergodic).
Let $\pi$ denote the stationary distribution of the local walk $Q$. For a vertex $j\in V$ and a spin/label $s_{j}\in\{0,1\}$, we have $\pi((j,s_{j}))=(1/n)\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}]$, and hence the distribution $\pi$ corresponds to the normalized marginal distribution of the Gibbs distribution $\mu$. Moreover, let $\pi_{j}$ denote the marginal distribution at vertex $j$; that is, $\pi_{j}((j,s_{j}))=\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}]$ for $s_{j}\in\{0,1\}$, and the remaining entries set to 0.
We will “zero-out” these trivial eigenvalues $1$ and 1/(n1)-1/(n-1) in the following manner:
$M_{\pi}=Q-\frac{n}{n-1}\mathbf{1}\pi^{T}+\frac{1}{n-1}\sum_{i=1}^{n}\mathbf{1}_{i}\pi_{i}^{T}.$
Then we notice that $M_{\pi}$ has the following block structure:
$M_{\pi}=\begin{pmatrix}A_{\pi}&-A_{\pi}\\ B_{\pi}&-B_{\pi}\end{pmatrix}$, (36)
where
$A_{\pi}-B_{\pi}=\frac{1}{n-1}(\Psi-I).$ (37)
To see (36) and (37) note that for $i\neq j$ we have
$M_{\pi}((i,s_{i}),(j,s_{j}))=\frac{1}{n-1}\Big{(}\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}\mid\sigma(i)=s_{i}]-\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=s_{j}]\Big{)},$
and
$M_{\pi}((i,s_{i}),(i,s_{i}))=M_{\pi}((i,s_{i}),(i,1-s_{i}))=0.$
Note that if $w$ is a left-eigenvector of $A_{\pi}-B_{\pi}$ then $\bigl{(}\begin{smallmatrix}w\\ -w\end{smallmatrix}\bigr{)}$ is a left-eigenvector of $M_{\pi}$ with the same eigenvalue. Moreover, vectors of the form $(v^{T}v^{T})$ (a space of dimension $n$) are both: right-eigenvectors of $M_{\pi}$ with eigenvalue 0, and are perpendicular to the left-eigenvectors of the form $\bigl{(}\begin{smallmatrix}w\\ -w\end{smallmatrix}\bigr{)}$. These vectors of the form $(v^{T}v^{T})$ yield right-eigenvectors of $\frac{n}{n-1}\mathbf{1}\pi^{T}-\frac{1}{n-1}\sum_{i=1}^{n}\mathbf{1}_{i}\pi_{i}^{T}$, where $\mathbf{1}$ has eigenvalue $1$ and the subspace perpendicular to $\pi$ (a space of dimension n1n-1) has eigenvectors with eigenvalue $\frac{1}{n-1}$. This implies (35).
This completes the proof of the lemma for the case without pinning. The proof easily generalizes to an arbitrary pinning $\tau\in\mathcal{P}_{k}$ with (n1)(n-1) replaced by (nk1)(n-k-1). ∎
4.5 Proof of Rapid Mixing (Theorem 1.4)
We can now utilize Lemma 4.2 and Theorem 4.1 to conclude $\mathrm{poly}(n)$ relaxation time for the Glauber dynamics and thereby prove Theorem 1.4.
Proof of Theorem 1.4.
Recall, the definition of spectral independence in Definition 1.2; it states that the maximum eigenvalue of the influence matrix $\Psi$ is at most $1+\eta$. Hence, from Lemma 4.2, we get that $\lambda_{2}(Q)\leq\eta/(n-1)$. Moreover, Definition 1.2 is for the worst-case pinning and hence we can apply Lemma 4.2 to any pinning. Therefore, we have that
$\gamma_{k}\geq 1-\eta/(n-k-1),$ (38)
where $\gamma_{k}$ is defined below (33). However, the above bound is not useful when $k\geq n-\eta-1$, but in this case we have the following bound.
When $k\geq n-\eta-1$ then the local walk $Q_{\tau}$ for $\tau\in\mathcal{P}_{k}$ is a walk on $2(n-k)\leq 2(\eta+1)$ states. Since $\eta$ is a constant, then this is a walk on a constant number of states and the marginal probabilities (and hence the transition probabilities) are independent of $n$, and hence the spectral gap $\gamma(Q_{\tau})$ is lower bounded by a constant $C=C(\eta)$, independent of $n$. Therefore, the following holds:
 $\mbox{ if }k\geq n-\eta-1,\mbox{ then }\gamma_{k}\geq C\mbox{ for some constant $C>0$.}$ (39)
Now we can apply (33) in the Random Walk Theorem (Theorem 4.1) and we obtain the following:
$\gamma(P_{\mathrm{Glauber}})$ $\geq\frac{1}{n}\prod_{k=0}^{n-2}\gamma_{k}$
$\geq C^{\lceil\eta\rceil}\times\frac{1}{n}\prod_{k=0}^{\lfloor n-\eta-2\rfloor}\gamma_{k}$ by Eq. 39
$\geq C^{\lceil\eta\rceil}\times\frac{1}{n}\prod_{k=0}^{\lfloor n-\eta-2\rfloor}\left(1-\frac{\eta}{n-k-1}\right)$ by Eq. 38
$\geq\frac{C^{\prime}}{n}\prod_{k=0}^{\lfloor n-\eta-2\rfloor}\left(1-\frac{\eta}{n-k-1}\right)$
$\geq\frac{C^{\prime}}{n}\exp\left(-\sum_{k=0}^{\lfloor n-\eta-2\rfloor}\frac{\eta}{n-k-1-\eta}\right)$ since $1-x\geq\exp\left(-\frac{x}{1-x}\right)$
$\geq\frac{C^{\prime}}{n}\exp\left(-\eta\sum_{i=1}^{\lfloor n-\eta-1\rfloor}\frac{1}{i}\right)$
$\geq\frac{C^{\prime\prime}}{n}\exp(-\eta(1+\ln(n-\eta)))$
$\geq\frac{C^{\prime\prime\prime}}{n}(n-\eta)^{-\eta}$
$\geq\frac{C^{\prime\prime\prime}}{n^{1+\eta}},$
where $C,C^{\prime},C^{\prime\prime},C^{\prime\prime\prime}$ are positive constants (which may depend on $\eta$).
We have shown that the spectral gap of the Glauber dynamics is $\Omega(1/n^{1+\eta})$. Applying Eq. 16 we obtain that the relaxation time satisfies $T_{\mathrm{relax}}=O(n^{1+\eta})$. Moreover, since $\log(\mu^{*})\leq O(n\log{q})$ and we are considering the case $q=2$, then, from Eq. 16, we have the mixing time bound of $T_{\mathrm{mix}}=O(n^{2+\eta})$. This establishes the upper bounds on the relaxation time and mixing time stated in Theorem 1.4. ∎
5 Random Walk Theorem: Proof of Local-to-Global
In this section we will prove the Random Walk Theorem (Theorem 4.1). We will use random walks between adjacent levels of the associated simplicial complex which corresponds to walks between sets $\mathcal{P}_{k}$ and $\mathcal{P}_{k+1}$. These will be denoted as up and down walks, and their composition as up-down and down-up walks.
5.1 Up and Down Walks
We begin with the definitions of the down-up and up-down chains which are at the heart of the proof. We do not explicitly utilize simplicial complexes, though readers familiar with simplicial complexes will understand the natural connection.
We start by defining a probability distribution on the collection of partial assignments to a set of $k$ vertices where $0\leq k\leq n$. Recall, $\tau\in\mathcal{P}_{k}$ is an assignment of spins $\{0,1\}$ to a subset $S\subseteq V$ where $|S|=k$. We define the distribution $\pi_{k}$ on $\mathcal{P}_{k}$ as follows:
$\mbox{for }\tau\in\mathcal{P}_{k},\ \ \ \ \pi_{k}(\tau)=\frac{1}{\binom{n}{k}}\mu(\tau),$
where
$\mu(\tau):=\sum_{\begin{subarray}{c}\eta\in\{0,1\}^{n}:\\ \tau(S)=\eta(S)\end{subarray}}\mu(\eta).$ (40)
For any $S\subset V$, observe that $\sum_{\tau:S\rightarrow\{0,1\}}\mu(\tau)=1$, and hence $\sum_{\tau\in\mathcal{P}_{k}}\pi_{k}(\tau)=1$. Moreover, for $k=n$, $\mathcal{P}_{n}=\Omega$ and $\pi_{n}=\mu$. (Note that $\pi_{1}$ is the same as the distribution $\pi$ defined in the proof of Lemma 4.2 in Section 4.4.)
It is useful to observe the following basic fact. Consider $0\leq k\leq r\leq n$ and $\sigma\in\mathcal{P}_{k}$ where $\sigma:S\rightarrow\{0,1\}$ for $S\subset V$ where $|S|=k$. First, observe that
$\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\mu(\eta)=\binom{n-k}{r-k}\mu(\sigma).$
Hence, we have the following:
$\pi_{k}(\sigma)=\frac{\mu(\sigma)}{{\binom{n}{k}}}=\frac{1}{\binom{n-k}{r-k}}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\frac{\mu(\eta)}{{\binom{n}{k}}}=\frac{1}{\binom{r}{k}}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\frac{\mu(\eta)}{\binom{n}{r}}={\binom{r}{k}}^{-1}\sum_{\eta\in\mathcal{P}_{r}:\sigma\subset\eta}\pi_{r}(\eta).$ (41)
We generalize the above definition to a fixed pinning $\eta\in\mathcal{P}$. For $0\leq\ell\leq n$, consider $\eta\in\mathcal{P}_{\ell}$ where $\eta$ is on $S$ (i.e., for $S\subset V$, $\eta:S\rightarrow\{0,1\}$). For $1\leq j\leq n$-$\ell$, let $\mathcal{P}_{\eta,j}$ be the set of assignments $\tau$ of spins $\{0,1\}$ to a subset $S^{\prime}\subseteq V\setminus S$ where $|S^{\prime}|=j$ and $\tau\cup\eta\in\mathcal{P}_{j+\ell}$.
We define the probability distribution for $j$ levels above $\eta$ with respect to the conditional Gibbs distribution $\mu_{\eta}$. Hence, for $j$ where $1\leq j\leq n$-$\ell$, define $\pi_{\eta,j}$ as follows:
$\mbox{for }\tau\in\mathcal{P}_{\eta,j}\mbox{ where }\tau:S^{\prime}\rightarrow\{0,1\}\mbox{ with }S^{\prime}\subset V\setminus S,\ \ \ \ \pi_{\eta,j}(\tau)=\frac{1}{\binom{n-\ell}{j}}\mu(\sigma(S^{\prime})=\tau\mid\sigma(S)=\eta).$
We will only use the above definitions of $\mathcal{P}_{\eta,j}$ and $\pi_{\eta,j}$ for the case $j=1$, in which case it is closely related to the local walk $Q_{\eta}$, see Remarks 5.2 and 5.3.
For every $1\leq k\leq n$, we define the following random walks which are known as Up and Down Operators.
Down Walk:
The down-walk is denoted by $\mathrm{P}_{k}^{\downarrow}$. From $\tau\in\mathcal{P}_{k}$ we remove an element of $\tau$ chosen uniformly at random. Note the elements of $\tau$ are (vertex,spin) pair. Hence, for $(j,s_{j})\in\tau$,
$\mathrm{P}_{k}^{\downarrow}(\tau,\tau\setminus(j,s_{j}))=1/k.$
Up Walk:
The up-walk is denoted by $\mathrm{P}_{k}^{\uparrow}$ and corresponds to the following stochastic process from $\mathcal{P}_{k}$ to $\mathcal{P}_{k+1}$. Starting from $\tau\in\mathcal{P}_{k}$ where $\tau$ is an assignment on $S\subset V,|S|=k$, then choose a random $j\not\in S$ and spin $s_{j}\in\{0,1\}$ where the probability of picking $(j,s_{j})$ is proportional to $\pi_{k+1}(\tau\cup(j,s_{j}))$. Hence,
$\mathrm{P}_{k}^{\uparrow}(\tau,\tau\cup\{(j,s_{j}\})=\frac{\pi_{k+1}(\tau\cup(j,s_{j}))}{(k+1)\pi_{k}(\tau)}=\frac{\mu(\tau\cup(j,s_{j}))}{(n-k)\mu(\tau)}=\frac{1}{(n-k)}\mu(\sigma(j)=s_{j}\mid\sigma(S)=\tau),$
where the denominator in the first identity comes from the observation that for $S\subset V,|S|=k$:
$\sum_{\begin{subarray}{c}(j^{\prime},s_{j^{\prime}}):\\ j^{\prime}\notin S,s^{\prime}_{j}\in\{0,1\}\end{subarray}}\pi_{k+1}(\tau\cup(j^{\prime},s_{j^{\prime}}))=\frac{1}{\binom{n}{k+1}}\sum_{j^{\prime}\notin S}\sum_{s_{j^{\prime}}\in\{0,1\}}\mu(\tau\cup(j^{\prime},s_{j^{\prime}}))\\ =\frac{1}{\binom{n}{k+1}}\sum_{j^{\prime}\notin S}\mu(\tau)=\frac{n-k}{\binom{n}{k+1}}\mu(\tau)=(k+1)\pi_{k}(\tau).$
From the definition of the up and down chains, we have that the reversibility condition is satisfied:
$\pi_{k}(\tau)\mathrm{P}_{k}^{\uparrow}(\tau,\tau\cup(j,s_{j}))=\pi_{k+1}(\tau\cup(j,s_{j}))\mathrm{P}_{k+1}^{\downarrow}(\tau\cup(j,s_{j}),\tau).$ (42)
Up-Down Walk:
The up-down walk is denoted by $\mathrm{P}_{k}^{\wedge}$ and corresponds to the following Markov chain from $\mathcal{P}_{k}$ to $\mathcal{P}_{k}$:
$\mathrm{P}_{k}^{\wedge}=\mathrm{P}_{k}^{\uparrow}\mathrm{P}_{k+1}^{\downarrow}.$
Consider $\sigma\in\mathcal{P}_{k-1}$ where $\sigma:S\rightarrow\{0,1\}$ for $S\subset V$. For $i,j\notin S$ and $s_{i},s_{j}\in\{0,1\}$, the transition $\mathrm{P}_{k}^{\wedge}(\sigma\cup(i,s_{i}),\sigma\cup(j,s_{j}))$ operates as follows. Starting from $\tau=\sigma\cup(i,s_{i})\in\mathcal{P}_{k}$, in the up-step, we choose a random $(j,s_{j})$ where the probability of picking $(j,s_{j})$ is proportional to $\pi_{k+1}(\tau\cup(j,s_{j}))$. Then in the down step, we remove an element $(i,s_{i})$ chosen uniformly at random from $\tau\cup(j,s_{j})$. More formally, we have the following:
$\mathrm{P}_{k}^{\wedge}(\sigma\cup(i,s_{i}),\sigma\cup(j,s_{j}))=\frac{\pi_{k+1}(\sigma\cup(i,s_{i})\cup(j,s_{j}))}{(k+1)^{2}\pi_{k}(\sigma)}.$ (43)
Down-Up Walk:
The down-up walk is denoted by $\mathrm{P}_{k}^{\vee}$. From $\tau\in\mathcal{P}_{k}$, we first remove a uniformly random element $(i,s_{i})$ from $\tau$ and then add an element $(j,s_{j})$ with probability proportional to the weight of the resulting set $\pi_{k}(\tau\cup(j,s_{j})\setminus(i,s_{i}))$,
$\mathrm{P}_{k}^{\vee}=\mathrm{P}_{k}^{\downarrow}\mathrm{P}_{k-1}^{\uparrow}.$
Observe that the stationary distribution of $\mathrm{P}_{k}^{\wedge}$ and $\mathrm{P}_{k}^{\vee}$ is $\pi_{k}$. The reversibility condition Eq. 42 shows that the two operators $\mathrm{P}_{k}^{\downarrow},\mathrm{P}_{k-1}^{\uparrow}$ are adjoint to each other, which implies that both up-down walks and down-up walks are reversible and positive semidefinite (PSD); see [CCGP25, Section 2] for more discussion.
Remark 5.1 (Glauber dynamics).
For a spin system, the down-up chain $\mathrm{P}_{n}^{\vee}$ is equivalent to the Glauber dynamics.
The above definitions can be considered as walks on levels of the simplicial complex defined by the Gibbs distribution $\mu$ on $\Omega\subset\{0,1\}^{V}$ where the ground set of the simplicial complex is the $2n$ pairs of (vertex,spin) assignments $(v,\sigma(v))\in V\times\{0,1\}$.
Let us illustrate the above definitions for the special case of independent sets (this is the hard-core model with $\lambda=1$); here the Gibbs distribution is uniformly distributed over all independent sets (of any size) of $G$. The corresponding simplicial complex has dimension $n=|V|$ since for every independent set $I$ we have $\{(v,1)\,|\,v\in I\}\cup\{(v,0)\,|\,v\not\in I\}$ in $\mathcal{P}_{n}$. The Glauber dynamics which updates the spin at a randomly chosen vertex is equivalent to $\mathrm{P}_{n}^{\vee}$.
Remark 5.2 (Local Walk is Non-Backtracking $\mathrm{P}_{1}^{\wedge}$).
Notice that the local walk $Q$ defined in Section 4.3 is similar to the up-down chain. For simplicity let us first consider the case without a pinning. Observe that for the up-down chain starting from $(i,s_{i})\in\mathcal{P}_{1}$ the following holds when $i\neq j$:
$\mathrm{P}_{1}^{\wedge}((i,s_{i}),(j,s_{j}))$ $=\mathrm{P}_{1}^{\uparrow}((i,s_{i}),(i,s_{i})\cup(j,s_{j}))\mathrm{P}_{2}^{\downarrow}((i,s_{i})\cup(j,s_{j}),(j,s_{j}))$
$=\frac{\pi_{2}((i,s_{i})\cup(j,s_{j}))}{2\pi_{1}((i,s_{i}))}\times\frac{1}{2}$
$=\frac{1}{2}\frac{n}{\binom{n}{2}}\times\frac{\mu((i,s_{i})\cup(j,s_{j}))}{2\mu(i,s_{i}))}$
$=\frac{1}{2}\frac{1}{n-1}\times\frac{\mu((i,s_{i})\cup(j,s_{j}))}{\mu(i,s_{i}))}$
$=\frac{1}{2}Q((i,s_{i}),(j,s_{j})).$
Observe that $\pi_{1}$ is the stationary distribution for both $\mathrm{P}_{1}^{\wedge}$ and $Q$, and that Remark 5.2 implies that
$\mathrm{P}_{1}^{\wedge}=(Q+I)/2.$
The extra factor $1/2$ comes from the step of the down-walk which drops an element chosen uniformly at random from $(i,s_{i}),(j,s_{j})$, whereas the local walk $Q$ corresponds to the non-backtracking walk which chooses $(j,s_{j})$ to avoid the self-loop.
Remark 5.3.
More generally, for any $1\leq k<n$, consider $\eta\in\mathcal{P}_{k-1}$ where $\eta$ is an assignment on $S\subset V$. Observe that $\pi_{\eta,1}$ is the stationary distribution for the local walk $Q_{\eta}$. For $(i,s_{i}),(j,s_{j})\in V\setminus S\times\{0,1\}$, we have the following for $i\neq j$:
$\mathrm{P}_{k}^{\wedge}(\eta\cup(i,s_{i}),\eta\cup(j,s_{j}))=\frac{1}{k+1}Q_{\eta}((i,s_{i}),(j,s_{j})).$ (44)
Before proceeding, let us verify the above identity Eq. 44. For $i\neq j$,
$\mathrm{P}_{k}^{\wedge}(\eta\cup(i,s_{i}),\eta\cup(j,s_{j}))$ $=\mathrm{P}_{k}^{\uparrow}(\eta\cup(i,s_{i}),\eta\cup(i,s_{i})\cup(j,s_{j}))\mathrm{P}_{k+1}^{\downarrow}(\eta\cup(i,s_{i})\cup(j,s_{j}),\eta\cup(j,s_{j}))$
$=\frac{\pi_{k+1}(\eta\cup(i,s_{i})\cup(j,s_{j}))}{(k+1)\pi_{k}(\eta\cup(i,s_{i}))}\times\frac{1}{(k+1)}$
$=\frac{\mu(\eta\cup(i,s_{i})\cup(j,s_{j}))}{(k+1)(n-k)\mu(\eta\cup(i,s_{i}))}$
$=\frac{1}{(k+1)}Q_{\eta}((i,s_{i}),(j,s_{j})).$
5.2 Spectrum of Up-Down Chains
We begin by pointing out that the spectral gaps for the up-down chain and down-up chain are the same.
Lemma 5.4.
$\gamma(\mathrm{P}_{k-1}^{\wedge})=\gamma(\mathrm{P}_{k}^{\vee}).$
Moreover, they have the same non-zero spectrum of eigenvalues:
$\mathrm{spectrum}_{\neq 0}(\mathrm{P}_{k-1}^{\wedge})=\mathrm{spectrum}_{\neq 0}(\mathrm{P}_{k}^{\vee}).$
Proof.
Note that $\mathrm{P}_{k-1}^{\wedge}=P^{\uparrow}_{k-1}P^{\downarrow}_{k}$ and $\mathrm{P}_{k}^{\vee}=P^{\downarrow}_{k}P^{\uparrow}_{k+1}$. Hence, the lemma follows from the general linear algebra fact that for an $n\times m$ matrix $A$ and $m\times n$ matrix $B$ then the non-zero spectrum of $AB$ is the same as the non-zero spectrum of $BA$. For a more detailed proof see [Mou22, Corollary 3.3.2]. ∎
5.3 Proof Setup
We can now state a more general version of the Random Walk Theorem (Theorem 4.1) which bounds the spectral gap of the down-up chain at any level. We will prove, for all $2\leq k\leq n$:
$\gamma(\mathrm{P}_{k}^{\vee})\geq\frac{1}{k}\prod_{i=0}^{k-2}\gamma_{i},$ (45)
where
$\gamma_{i}=\min_{\tau\in\mathcal{P}_{i}}\gamma(Q_{\tau})$
is the spectral gap for the local walk with a worst-case pinning of $i$ vertices. Note, the claim trivially holds for $k=1$ as $\gamma(\mathrm{P}_{1}^{\vee})=1$. Since the Glauber dynamics is equivalent to $\mathrm{P}_{n}^{\vee}$, then Theorem 4.1 is the special case of Eq. 33 where $k=n$.
5.4 Key Technical Lemma
The following lemma will be the key tool in the inductive proof of Theorem 4.1. It relates the Dirichlet form for the up-down walk $\mathrm{P}_{k}^{\wedge}$ at level $k$ with the down-up walk $\mathrm{P}_{k}^{\vee}$ at level $k$. Recall that the spectral gap of the $\mathrm{P}_{k}^{\wedge}$ is equal to the spectral gap of $\mathrm{P}_{k+1}^{\wedge}$ (see Lemma 5.4). Hence, the lemma is in essence an inductive statement as it relates the chain $\mathrm{P}_{k}^{\wedge}$ to $\mathrm{P}_{k}^{\vee}$, rather than $\mathrm{P}_{k+1}^{\vee}$. In this manner the lemma will immediately yield an inductive proof of the Random Walk Theorem.
Lemma 5.5.
For all $0\leq k<n$, and any $f:\mathcal{P}_{k}\to\mathbb{R}$, the following holds:
$\mathcal{E}_{\mathrm{P}_{k}^{\wedge}}(f)\geq\frac{k}{k+1}\gamma_{k-1}\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f).$
At an intuitive level, the reader can understand the factor $k/(k+1)$ because in the LHS the down operator is from level $k+1$ to $k$ and hence a transition occurs with probability $1/(k+1)$, whereas in the RHS the down operator is from level $k$ to k1k-1 and hence occurs with probability $1/k$.
We will utilize the following two claims to prove Lemma 5.5. The first claim captures the local-to-global nature of the process. In particular, for the chain $\mathrm{P}_{k}^{\wedge}$ we can express its Dirichlet form in terms of local walks for pinnings at level k1k-1. Recall that the Dirichlet form for a chain $P$ is the local variation of a functional $f$ over pairs of states connected by a transition of $P$.
For the chain $\mathrm{P}_{k}^{\wedge}$, consider a pair of distinct states $\sigma,\tau$ where $\mathrm{P}_{k}^{\wedge}(\sigma,\tau)>0$. Notice that it must be the case that $|\sigma\oplus\tau|=2$ where $\oplus$ is the symmetric difference, i.e., $\sigma\oplus\tau=(\sigma\setminus\tau)\cup(\tau\setminus\sigma)$. This is because in the up-step we can add one element $(v_{i},s_{i})$ to $\sigma$ and then in the down-step we can remove one element $(v_{j},s_{j})$ from $\sigma\cup(v_{i},s_{i})$. Hence, $\sigma\cap\tau=\eta$ where $\eta\in\mathcal{P}_{k-1}$. Then the transition $\mathrm{P}_{k}^{\wedge}(\sigma,\tau)$ is closely related to the transition $Q_{\eta}((v_{i},s_{i}),(v_{j},s_{j}))$, which yields the following identity.
Claim 5.6.
For all $0<k\leq n$, and all $f:\mathcal{P}_{k}\to\mathbb{R}$,
$\mathcal{E}_{\mathrm{P}_{k}^{\wedge}}(f)=\frac{k}{k+1}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathcal{E}_{Q_{\eta}}(f_{\eta}),$
where for $\eta:S\rightarrow\{0,1\}$ and $(j,s_{j})\in(V\setminus S)\times\{0,1\}$, then $f_{\eta}((j,s_{j}))=f(\eta\cup(j,s_{j}))$.
The second claim works in the reverse manner and relates the variance on level $1$ conditional on $\eta\in\mathcal{P}_{k-1}$ to the global chain $\mathrm{P}_{k}^{\vee}$.
Claim 5.7.
For all $0<k\leq n$, and all $f:\mathcal{P}_{k}\to\mathbb{R}$,
$\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f)=\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathrm{Var}_{\pi_{\eta,1}}(f_{\eta}).$
Using these two claims it is straightforward to prove the key technical lemma.
Proof of Lemma 5.5.
For $1\leq k\leq n$, and $\eta\in\mathcal{P}_{k-1}$ we know that $\gamma(Q_{\eta})\geq\gamma_{k-1}$, and hence since $\pi_{\eta,1}$ is the stationary distribution for $Q_{\eta}$ then for any $g:(V\setminus S)\times\{0,1\}\rightarrow\mathbb{R}$ we have that
$\mathcal{E}_{Q_{\eta}}(g)\geq\gamma_{k-1}\mathrm{Var}_{\pi_{\eta,1}}(g).$ (46)
We can now prove the lemma:
$\mathcal{E}_{\mathrm{P}_{k}^{\wedge}}(f)$ $=\frac{k}{k+1}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathcal{E}_{Q_{\eta}}(f_{\eta})$ by Claim 5.6
$\geq\frac{k}{k+1}\gamma_{k-1}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathrm{Var}_{\pi_{\eta,1}}(f_{\eta})$ by Eq. 46 with $g=f_{\eta}$
$=\frac{k}{k+1}\gamma_{k-1}\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f),$ by Claim 5.7
as claimed. ∎
Proof of Claim 5.6.
As a warm-up let us consider the simpler case where $k=1$. Recall that $\pi_{1}$ is the stationary distribution for $\mathrm{P}_{1}^{\wedge}$ and $Q$. In Remark 5.2 we showed that for $(i,s_{i}),(j,s_{j})\in V\times\{0,1\}$ where $i\neq j$,
$\mathrm{P}_{1}^{\wedge}((i,s_{i}),(j,s_{j}))=\frac{1}{2}Q((i,s_{i}),(j,s_{j})).$
For any $f:\mathcal{P}_{1}\to\mathbb{R}$ (which is $f:V\times\{0,1\}\rightarrow\mathbb{R}$), we have the following:
$\mathcal{E}_{\mathrm{P}_{1}^{\wedge}}(f)$ $=\frac{1}{2}\sum_{(i,s_{i}),(j,s_{j})}\pi_{1}((i,s_{i}))\mathrm{P}_{1}^{\wedge}((i,s_{i}),(j,s_{j}))(f((i,s_{i}))-f((j,s_{j})))^{2}$
$=\frac{1}{4}\sum_{(i,s_{i}),(j,s_{j})}\pi_{1}((i,s_{i}))Q((i,s_{i}),(j,s_{j}))(f((i,s_{i}))-f((j,s_{j})))^{2}$
$=\frac{1}{2}\mathcal{E}_{Q}(f),$
where the summation is over $(i,s_{i}),(j,s_{j})\in V\times\{0,1\}$.
Since $\mathcal{P}_{0}=\{\emptyset\}$, this proves the claim for the case $k=1$.
Before proving the general claim, consider the following observation. For $\eta\cup a\in\mathcal{P}_{k}$, note that:
$\pi_{k}(\eta\cup a)=\frac{1}{\binom{n}{k}}\mu(\eta\cup a)=\frac{k}{(n-k+1)}\pi_{k-1}(\eta)\mu_{\eta}(a)=k\pi_{k-1}(\eta)\pi_{\eta,1}(a).$ (47)
In general, for $k\geq 1$ we have the following for any $f:\mathcal{P}_{k}\to\mathbb{R}$:
$\mathcal{E}_{\mathrm{P}_{k}^{\wedge}}(f)$ $=\frac{1}{2}\sum_{\eta\in\mathcal{P}_{k-1}}\sum_{\begin{subarray}{c}a,b\in V\setminus S\times\{0,1\}:\\ \eta:S\rightarrow\{0,1\}\end{subarray}}\pi_{k}(\eta\cup a)\mathrm{P}_{k}^{\wedge}(\eta\cup a,\eta\cup b)(f(\eta\cup a)-f(\eta\cup b))^{2}$
$=\frac{1}{2(k+1)}\sum_{\eta\in\mathcal{P}_{k-1}}\sum_{a,b}\pi_{k}(\eta\cup a)Q_{\eta}(a,b)(f_{\eta}(a)-f_{\eta}(b))^{2}$ by Eq. 44
$=\frac{k}{2(k+1)}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{a,b}\pi_{\eta,1}(a)Q_{\eta}(a,b)(f_{\eta}(a)-f_{\eta}(b))^{2}$ by Eq. 47
$=\frac{k}{k+1}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathcal{E}_{Q_{\eta}}(f_{\eta}).\qed$
Proof of Claim 5.7.
$\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathrm{Var}_{\pi_{\eta,1}}(f_{\eta})$
$=\frac{1}{2}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{\begin{subarray}{c}a,b\in V\setminus S\times\{0,1\}:\\ \eta:S\rightarrow\{0,1\}\end{subarray}}\pi_{\eta,1}(\eta\cup a)\pi_{\eta,1}(\eta\cup b)(f_{\eta}(a)-f_{\eta}(b))^{2}$
$=\frac{1}{2}\sum_{\eta,a,b}\pi_{k-1}(\eta)\frac{\mu_{\eta}(a)}{(n-k+1)}\times\frac{\mu_{\eta}(b)}{(n-k+1)}(f(\eta\cup\{a\})-f(\eta\cup\{b\}))^{2}$
$=\frac{1}{2}\sum_{\eta,a,b}\pi_{k}(\eta\cup\{a\})\times\frac{1}{k}\times\frac{\mu_{\eta}(b)}{(n-k+1)}(f(\eta\cup\{a\})-f(\eta\cup\{b\}))^{2}$
$=\frac{1}{2}\sum_{\eta,a,b}\pi_{k}(\eta\cup\{a\})P^{\downarrow}_{k}(\eta\cup\{a\},\eta)P^{\uparrow}_{k-1}(\eta,\eta\cup\{b\})(f(\eta\cup\{a\})-f(\eta\cup\{b\}))^{2}$
$=\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f).\qed$
5.5 Inductive Proof of Random Walk Theorem: Proof of Theorem 4.1
We can now prove Eq. 45, which implies the Random Walk Theorem (Theorem 4.1).
Proof of Theorem 4.1.
Our goal is to prove by induction that for all $2\leq\ell\leq n$:
$\gamma(\mathrm{P}_{\ell}^{\vee})\geq\frac{1}{\ell}\prod_{i=0}^{\ell-2}\gamma_{i}.$ (48)
This is equivalent to the following statement in terms of the Dirichlet form and variance:
$\forall f:\Omega\rightarrow\mathbb{R},\ \ \mathcal{E}_{\mathrm{P}_{\ell}^{\vee}}(f)\geq\frac{1}{\ell}\prod_{i=0}^{\ell-2}\gamma_{i}\mathrm{Var}_{\pi_{\ell}}(f).$ (49)
Recall $\gamma(\mathrm{P}_{\ell-1}^{\wedge})=\gamma(\mathrm{P}_{\ell}^{\vee})$ was established in Lemma 5.4. Hence, Eq. 48 is equivalent to the following statement:
$\forall f:\Omega\rightarrow\mathbb{R},\ \ \mathcal{E}_{\mathrm{P}_{\ell-1}^{\wedge}}(f)\geq\frac{1}{\ell}\prod_{i=0}^{\ell-2}\gamma_{i}\mathrm{Var}_{\pi_{\ell-1}}(f)$ (50)
We will prove Eq. 50 by induction and use Eq. 49 in the inductive step.
Now let us assume the inductive hypothesis Eq. 48 for all $\ell<k$ and we will establish it for $\ell=k$.
$\mathcal{E}_{\mathrm{P}_{k-1}^{\wedge}}(f)$ $\geq\frac{(k-1)\gamma_{k-2}}{k}\mathcal{E}_{\mathrm{P}_{k-1}^{\vee}}(f)$ by Lemma 5.5
$\geq\frac{(k-1)\gamma_{k-2}}{k}\times\frac{1}{k-1}\prod_{i=0}^{k-3}\gamma_{i}\mathrm{Var}_{\pi_{k-1}}(f)$ by Eq. 49 for $\ell=k-1$
$=\frac{1}{k}\prod_{i=0}^{k-2}\gamma_{i}\mathrm{Var}_{\pi_{k-1}}(f),$
which establishes Eq. 50 (and hence Eq. 48) by induction. ∎
6 Optimal Mixing Time for Glauber from Spectral Independence
The goal of this section is to prove that when the system is $\eta$-spectrally independent, then the relaxation time of the Glauber dynamics is $O(n)$, thereby proving Theorem 1.5 of [CLV21a]. Our proof follows the same approach as in [CLV21a]. In Section 6.8 we comment on how the proof changes to obtain Theorem 1.8 establishing an optimal $O(n\log{n})$ mixing time, instead of relaxation time.
6.1 Uniform Block Dynamics
Consider the following dynamics, parameterized by $0<\alpha<1$, which we will refer to as the $\alpha n$-uniform block dynamics. The dynamics updates $\alpha n$ random vertices $S$ in each step. Assume $\alpha n$ is an integer. From $Y_{t}\in\Omega$, the transitions $Y_{t}\rightarrow Y_{t+1}$ of the $\alpha n$-uniform block dynamics are as follows:
  1. 1.
    Select $\alpha n$ random vertices, chosen uniformly at random from all $\binom{n}{\alpha n}$ vertices. Let $S$ denote the chosen set.
  2. 2.
    For all $w\notin S$, let $Y_{t+1}(w)=Y_{t}(w)$.
  3. 3.
    Sample $Y_{t+1}(S)$ from the conditional Gibbs distribution $\mu(\sigma(S)\mid\sigma(w)=Y_{t+1}(w)\mbox{ for all }w\notin S).$
Note the $\alpha n$-uniform block dynamics is identical to $\mathrm{P}_{n,(1-\alpha)n}^{\vee}$. Using the Random Walk Theorem framework we will first prove fast mixing of the $\alpha$-block dynamics. In particular we will show that the relaxation time is $O(1)$ where the constant in the big-$O()$ notation is a function of $1/\alpha$. We will then show that this implies fast mixing of the Glauber dynamics when $\alpha=O(1/\Delta)$ where $\Delta$ is the maximum degree of the graph.
6.2 Improved Random Walk Theorem
The Random Walk Theorem is not strong enough to establish fast mixing of the $\alpha n$-uniform block dynamics, we will need the following improved local-to-global theorem which was presented by [CLV21a, Theorem 5.4] and independently [GM20, Theorem 4]. For simplicity we only present the spectral gap version, the stronger version for entropy contraction is presented in [CLV21a, GM20].
For $0<i\leq n$, let
$\Gamma_{i}:=\prod_{j=0}^{i-1}(2\gamma_{j}-1),$
and let $\Gamma_{0}=1$. Note that since, for all $j$, $\gamma_{j}\leq 1$ as the spectral gap is at most $1$, then $(2\gamma_{j}-1)\leq 1$ and hence $\Gamma_{i}\leq 1$ for all $i$.
We will prove, for $0<k\leq n$,
$\gamma(\mathrm{P}_{k}^{\vee})\geq\frac{\Gamma_{k-1}}{\sum_{i=0}^{k-1}\Gamma_{i}}$ (51)
The above inequality will be a special case of the following improvement of the Random Walk Theorem.
Theorem 6.1 (Improved RW Theorem [CLV21a, Theorem A.9],[GM20, Theorem 4]).
For $0\leq\ell<n$,
$\gamma(\mathrm{P}_{n,\ell}^{\vee})\geq\frac{\sum_{i=\ell}^{n-1}\Gamma_{i}}{\sum_{i=0}^{n-1}\Gamma_{i}}$ (52)
Moreover, one could further improve the above result by replacing $(2\gamma_{j}-1)$ by $\gamma_{j}/(2-\gamma_{j})$ in the definition of $\Gamma_{i}$; see the proof of Lemma 6.5 for details.
Before proving the above theorem let us first compare it to the Random Walk Theorem (Theorem 4.1) stated earlier. Looking at Eq. 51 for $k=n$ (or (52) for $\ell=n-1$) we have:
$\gamma(\mathrm{P}_{n}^{\vee})\geq\frac{\Gamma_{n-1}}{\sum_{i=0}^{n-1}\Gamma_{i}}=\frac{\prod_{j=0}^{n-1}(2\gamma_{j}-1)}{\sum_{i=0}^{n-1}\Gamma_{i}}\geq\frac{1}{n}\prod_{j=0}^{n-1}(2\gamma_{j}-1)\ \ \ \mbox{ since $\Gamma_{i}\leq 1$.}$
In contrast, Theorem 4.1 has $2\gamma_{j}-1$ replaced by $\gamma_{j}$ in the lower bound on the spectral gap of $\mathrm{P}_{n}^{\vee}$.
6.3 Fast Mixing of Uniform Block Dynamics
Applying Theorem 6.1 we establish fast mixing of the $\alpha n$-uniform block dynamics.
Lemma 6.2.
For all $\eta>0$ and $0<\alpha<1$ there exists a constant $C=C(\eta,\alpha)$ such that if the system is $\eta$-spectral independence then the spectral gap of the $\alpha n$-uniform block dynamics is $\gamma(\mathrm{P}_{n,(1-\alpha)n}^{\vee})\geq C$.
Proof.
We begin with a useful lower bound on $\Gamma_{(1-\alpha/2)n}$:
$\Gamma_{(1-\alpha/2)n}=\prod_{i=0}^{(1-\alpha/2)n-2}(2\gamma_{i}-1)\\ \geq\prod_{i=0}^{(1-\alpha/2)n-1}\left(1-\frac{2\eta}{n-i-1}\right)\geq\left(1-\frac{4\eta}{\alpha n}\right)^{n}\geq\exp(-8\eta/\alpha),$ (53)
for $n\geq 8\eta/\alpha$, where the last inequality uses that $1-x\geq\exp(-2x)$ for $x\leq 1/2$.
We can now proceed to lower bound the spectral gap of the $\alpha n$-uniform block dynamics:
$\gamma(\mathrm{P}_{n,(1-\alpha)n}^{\vee})$ $\geq\frac{\sum_{i=(1-\alpha)n}^{n-1}\Gamma_{i}}{\sum_{i=0}^{n-1}\Gamma_{i}}$ by Eq. 52
$\geq\frac{\sum_{i=(1-\alpha)n}^{(1-\alpha/2)n}\Gamma_{i}}{n}$ since $\Gamma_{i}\leq 1$
$\geq\frac{(\alpha/2)n\Gamma_{(1-\alpha/2)n}}{n}$ since $\Gamma_{i}\geq\Gamma_{i+1}$
$\geq(\alpha/2)\exp(-8\eta/\alpha)$ by Eq. 53
$=C(\alpha,\eta).\qed$
6.4 Shattering
We established fast mixing of the $\alpha n$-uniform block dynamics, see Lemma 6.2, which updates $\alpha n$ random vertices in each step. We will use that to establish fast mixing of the Glauber dynamics.
The first step is to look at the properties of the updated vertices in a step of the block dynamics. In particular, [CLV21a] showed that a random subset of $\alpha n$ vertices is “shattered” in the sense that the expected size of each component is $O(1)$ when $\alpha<1/(6\Delta)$ where $\Delta$ is the maximum degree of the graph $G$. The shattering occurs because $\alpha\Delta<1$ and hence this corresponds to a sub-critical branching process.
For a subset $S\subset V$, let $\mathcal{C}_{S}$ denote the collection of connected components in the induced subgraph on $S$, and for $v\in V$, let $C_{v}\in\mathcal{C}_{S}$ denote the component containing $v$.
Lemma 6.3.
[CLV21a, Lemma 4.3] For a graph $G=(V,E)$ of maximum degree $\Delta$, for $\alpha>0$, choose a random subset $S\subset V$ where $|S|=\alpha n$. Then for every integer $k\geq 1$, for every $v\in V$,
$\mathrm{Pr}\left(|T_{v}|=k\right)\leq\alpha(6\Delta\alpha)^{k-1},$
where $T_{v}$ is the component containing $v$ in the induced subgraph on $S$.
Proof.
The number of distinct subgraphs of size $k$ that contain $v$ is bounded by $\binom{\Delta k}{k-1}$ (we can enccode them by a binary string of length $\Delta k$ with exactly k1k-1 ones as follows: running the DFS algorithm starting from $v$, for each encountered edge output $1$ if it is traversed and 0 otherwise, and at the end pad the string with zeros). The probability that a particular set of size $k$ is contained in $S$ is $\binom{n-k}{\alpha n-k}/\binom{n}{\alpha n}\leq\alpha^{k}$. Combining the two estimates with $\binom{n}{k}\leq(n{\mathrm{e}}/k)^{k}$ we obtain
$\mathrm{Pr}\left(|T_{v}|=k\right)\leq\left(\frac{\Delta{\mathrm{e}}k}{k-1}\right)^{k}\alpha^{k}\leq\alpha(6\Delta\alpha)^{k-1}.$
6.5 Optimal Relaxation Time of Glauber: Proof of Theorem 1.5
Using the above shattering result (Lemma 6.3) with the fast mixing result for the $\alpha n$-uniform block dynamics (Lemma 6.2), we can prove optimal upper bound on the relaxation time of the Glauber dynamics and hence prove Theorem 1.5.
The high-level proof idea for Theorem 1.8 is the following. Let $S\subset V$ be a random subset where $|S|=\alpha n$ as in Lemma 6.3. Let $\mathcal{C}_{S}$ denote the connected components in the induced subgraph on $S$. When applying the heat-bath dynamics on $S$, then each $T,T^{\prime}\in\mathcal{C}_{S}$ are independent of each other. Hence, given a configuration $\tau$ on $\overline{S}$, then since we have a product distribution over each component $T\in\mathcal{C}_{S}$ we have:
$\mathrm{Var}_{S}[F|\tau]=\sum_{T\in\mathcal{C}_{S}}\mathrm{Var}_{T}[F|\tau].$
Within a component $T\in\mathcal{C}_{S}$ we can apply Theorem 1.4 to conclude a polynomial bound on the spectral gap of the Glauber dynamics within $T$ (with fixed configuration $\tau$ on the boundary of $T$), and hence
$\mathrm{Var}_{T}[F|\tau]\leq C^{\prime}|T|^{\eta+1}\sum_{v\in T}\mathcal{E}_{\mathrm{P^{\tau}_{Gl}}(T)}(f).$ (54)
Finally, we apply Lemma 6.3 to utilize the upper bound on the size of $T$ and conclude the proof.
The bound in Eq. 54 is the only step in the proof where the analogous bound does not hold with respect to entropy in place of variance. This is an issue when we aim to upper bound the mixing time instead of the relaxation time. To obtain a bound on the entropy analog of Eq. 54 we need the additional assumption of marginal boundedness; see Section 6.8 for further details.
Proof of Theorem 1.5.
In the following, for a subset $S\subset V$, we use $\mathrm{P_{HB}^{\tau}}(S)$ to denote the heat-bath block dynamics on $S$; this dynamics updates the configuration on all of $S$ in one step from the Gibbs distribution conditional on the fixed configuration $\tau$ on $\overline{S}$. Similarly, we use $\mathrm{P^{\tau}_{Gl}}$ to denote the Glauber dynamics on a set $S$ with a fixed configuration $\tau$ on $\overline{S}$; in particular, for $v\in S$, with probability $1/|S|$ we update the configuration at $v$ from the Gibbs distribution conditional on the fixed configuration for the other vertices.
For a function $f:\Omega\rightarrow\mathbb{R}$, we will consider the conditional variance. For $\sigma\sim\mu$, let
$F=f(\sigma).$
Note that when we write $\mathrm{Var}(f)$ we are implicitly writing $\mathrm{Var}[F]$ since we first draw a sample $\sigma$ from the Gibbs distribution $\mu$ and then look at the variance of $F=f(\sigma)$.
Consider a configuration $\tau\in\Omega$, and a subset $S\subset V$. In the upcoming analysis we will consider the variance of $f$ within $S$ where we fix $\tau$ on $\overline{S}=V\setminus S$. We denote this as:
$\mathrm{Var}_{S}[F\mid\tau]:=\mathrm{Var}[F\mid\sigma(\overline{S})=\tau(\overline{S})].$
Once again, in words $\mathrm{Var}_{S}[F\mid\tau]$ is the variance of the function $f$ over the choice of $\sigma$ from the Gibbs distribution $\mu$ conditional on $\sigma(\overline{S})=\tau(\overline{S})$.
Again fix $\tau\in\Omega$ and $S\subset V$. Consider the heat-bath dynamics on $S$ with a fixed $\tau$ on $\overline{S}$. The dynamics chooses the configuration on $S$ from $\mu_{\tau(\overline{S})}$, which is the Gibbs distribution $\mu$ conditional on the configuration on $\overline{S}$ being $\tau(\overline{S})$. Hence,
$\mathcal{E}_{\mathrm{P_{HB}^{\tau}}(S)}(f)=\mathrm{Var}[F\mid\sigma(\overline{S})=\tau(\overline{S})]$ (55)
Moreover, if $S=\{v\}$ for $v\in V$ then this is heat-bath on the single vertex $v$. The Glauber dynamics operates by choosing a random vertex $v$ and then applying the heat-bath dynamics on $v$. Hence, we have the following:
$\mathcal{E}_{\mathrm{P_{\textsc{gd}}}}(f)$ $=\frac{1}{2}\sum_{\tau\in\Omega}\sum_{\sigma\in\Omega}\mu(\tau)\mathrm{P_{\textsc{gd}}}(\tau,\sigma)(f(\tau)-f(\sigma))^{2}$
$=\sum_{\tau\in\Omega}\mu(\tau)\frac{1}{n}\sum_{v\in V}\left[\frac{1}{2}\sum_{\begin{subarray}{c}\sigma\in\Omega:\\ \sigma(V\setminus\{v\})=\tau(V\setminus\{v\})\end{subarray}}\mu(\sigma)(f(\tau)-f(\sigma))^{2}\right]$
$=\frac{1}{n}\sum_{\tau\in\Omega}\mu(\tau)\mathrm{Var}[F\mid\sigma(V\setminus\{v\})=\tau(V\setminus\{v\})].$ (56)
Recall, for a subset $S\subset V$, $\mathcal{C}_{S}$ denotes the collection of connected components in the induced subgraph on $S$. For any function $f:\Omega\rightarrow\mathbb{R}$,
$\mathrm{Var}(f)$ $\leq\frac{1}{C}\mathcal{E}_{\mathrm{P}_{n,(1-\alpha)n}^{\vee}}(f)$ by Lemma 6.2
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\mathcal{E}_{\mathrm{P_{HB}^{\tau}}(S)}(f)$ (see below)
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\mathrm{Var}_{S}[F\mid\tau]$ by Eq. 55
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}\mathrm{Var}_{T}[F\mid\tau]$ by independence of $T,T^{\prime}\in\mathcal{C}_{S}$
$\leq\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}C^{\prime}|T|^{\eta+1}\sum_{v\in T}\mathcal{E}_{\mathrm{P^{\tau}_{Gl}}(T)}(f)$ by Theorem 1.4, (see below)
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}C^{\prime}|T|^{\eta+1}\frac{1}{|T|}\sum_{v\in T}\mathrm{Var}_{v}[F\mid\tau]$ by Eq. 56
$=\frac{1}{C}\sum_{\tau\in\Omega}\mu(\tau)\sum_{v\in V}\mathrm{Var}_{v}[F\mid\tau]\sum_{k=1}^{\alpha n}\mathrm{Pr}\left(|T_{v}|=k\right)\times C^{\prime}k^{\eta-1}$ rearranging
$\leq\frac{1}{C}\sum_{\tau\in\Omega}\mu(\tau)\sum_{v\in V}\mathrm{Var}_{v}[F\mid\tau]\sum_{k=1}^{\alpha n}C^{\prime}(6\Delta\alpha)^{k-1}k^{\eta-1}$ by Lemma 6.3
$\leq\frac{1}{C}\sum_{\tau\in\Omega}\mu(\tau)\sum_{v\in V}\mathrm{Var}_{v}[F\mid\tau]\sum_{k=1}^{\alpha n}C^{\prime}2^{-k}$ for $\alpha<\exp(-\eta)/(100\Delta)$
$\leq\frac{1}{C^{\prime\prime}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{v\in C}\mathrm{Var}_{v}[F\mid\tau]$
$=\frac{n}{C^{\prime\prime}}\mathcal{E}_{\mathrm{P_{\textsc{gd}}}}(f)$ by Eq. 56.
In the second line we are implementing the $(n,(1-\alpha)n)$ down-up dynamics by first choosing the subset $S$ of size $\alpha n$ for update and then applying the heat-bath dynamics on this set $S$ with those vertices outside $S$ fixed to $\tau$ chosen from the Gibbs distribution $\mu$. And in the fifth line where we are applying Theorem 1.4, note we are considering the Glauber dynamics which only operates on the vertices in this component $T$, where the vertices outside $T$ have the fixed configuration $\tau$.
This proves that the spectral gap of the Glauber dynamics is $\geq C^{\prime\prime}/n$ for a constant $C^{\prime\prime}=C^{\prime\prime}(\alpha,\Delta)$. Hence, by applying Eq. 16, the relaxation time of the Glauber dynamics is $\leq Cn$ for some constant $C=C(\alpha,\Delta)$. For the mixing time bound, since $\log(1/\mu^{*})=O(n\log{q})$ and we are considering the case $q=2$, then by Eq. 16, we have $T_{\mathrm{mix}}\leq Cn^{2}$ for some constant $C=C(\alpha,\Delta)$. This establishes the bounds on the relaxation time and mixing time stated in Theorem 1.5. ∎
6.6 Improved Random Walk Theorem: Proof of Theorem 6.1
In this section we will prove the improved Random Walk Theorem presented in Theorem 6.1 in Section 6.2. The proof will involve statements comparing Dirichlet forms and variance on different levels and hence we will need to define the projection of a function $f$ on lower levels.
Consider an arbitrary function $f:\Omega\rightarrow\mathbb{R}$. Recall $\mathcal{P}_{n}=\Omega$, and let $f^{(n)}=f$. We will define $f^{(k)}$ for $0\leq k<n$ in the following inductive manner.
For $f^{(k+1)}:\mathcal{P}_{k+1}\rightarrow\mathbb{R}$, let
$f^{(k)}=\mathrm{P}_{k}^{\uparrow}f^{(k+1)}.$
Hence, for $\sigma\in\mathcal{P}_{k}$ we have that $f^{(k)}(\sigma)=\sum_{\tau\in\mathcal{P}_{k+1}}\mathrm{P}_{k}^{\uparrow}(\sigma,\tau)f^{(k+1)}(\tau)$.
We will use the two identities involving the (global) variance and the local variance, measured by the Dirichlet form. The following is an interesting and useful identity for relating the Dirichlet form for the down-up walk from level $i$ to $j$ in terms of the difference of the variance at levels $i$ and $j$.
Lemma 6.4.
For all $n\geq i>j\geq 0$, for all $f^{(i)}:\mathcal{P}_{i}\rightarrow\mathbb{R}$, the following holds:
$\mathcal{E}_{\mathrm{P}_{i,j}^{\vee}}(f^{(i)})=\mathrm{Var}_{\pi_{i}}(f^{(i)})-\mathrm{Var}_{\pi_{j}}(f^{(j)}).$ (57)
In the proof of the Random Walk Theorem (Eq. 45), the key technical inequality stated in Lemma 5.5 was the following:
$\mathcal{E}_{\mathrm{P}_{k}^{\wedge}}(f)\geq\frac{k}{k+1}\gamma_{k-1}\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f).$
For the improved result we will use the following variant. Recall, $f^{(k)}=\mathrm{P}_{k}^{\uparrow}f^{(k+1)}$.
Lemma 6.5.
For any $f^{(k+1)}:\mathcal{P}_{k+1}\rightarrow\mathbb{R}$, the following holds:
$\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})\geq(2\gamma_{k-1}-1)\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)}).$ (58)
We defer the proofs of Lemmas 6.4 and 6.5 to Section 6.7, and we proceed with the proof of Theorem 6.1.
Proof of Theorem 6.1.
We begin by proving Eq. 51 which states that the spectral gap of $\mathrm{P}_{k}^{\vee}$ is at least $\Gamma_{k-1}/(\sum_{i=0}^{k-1}\Gamma_{i}$ where $\Gamma_{i}=\prod_{j=0}^{i-1}(2\gamma_{j}-1)$. We will then use this result for the one-level down-up walk to derive the more general statement in Theorem 6.1 for the down-up walk $\mathrm{P}_{n,\ell}^{\vee}$.
Our inductive hypothesis is the following statement about the spectral gap of $\mathrm{P}_{k}^{\vee}$:
$\left(\sum_{i=0}^{k-1}\Gamma_{i}\right)\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)})\geq\Gamma_{k-1}\mathrm{Var}_{\pi_{k}}(f^{(k)}).$ (59)
We can now complete the proof by induction as follows:
$\left(\sum_{i=0}^{k}\Gamma_{i}\right)\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})$
$=\Gamma_{k}\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\left(\sum_{i=0}^{k-1}\Gamma_{i}\right)\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})$
$\geq\Gamma_{k}\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\left(\sum_{i=0}^{k-1}\Gamma_{i}\right)(2\gamma_{k-1}-1)\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)})$ by Eq. 58
$\geq\Gamma_{k}\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\Gamma_{k-1}(2\gamma_{k-1}-1)\mathrm{Var}_{\pi_{k}}(f^{(k)})$ by inductive hypothesis
$=\Gamma_{k}\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\Gamma_{k}\mathrm{Var}_{\pi_{k}}(f^{(k)})$ $\mbox{since }\Gamma_{k}=\gamma_{k-1}\Gamma_{k-1}$
$=\Gamma_{k}\mathrm{Var}_{\pi_{k+1}}(f^{(k+1)})$ by Lemma 6.4 with $i=k+1,j=k$.
This completes the proof of Eq. 51.
Let us now prove the more general statement in Eq. 52. Establishing the bound stated in Eq. 52 on the spectral gap of $\mathrm{P}_{n,\ell}^{\vee}$ is equivalent to proving the following for all $0<\ell<n$ and all $f:\Omega\rightarrow\mathbb{R}$:
$\left(\sum_{i=0}^{n-1}\Gamma_{i}\right)\mathcal{E}_{\mathrm{P}_{n,\ell}^{\vee}}(f)\geq\left(\sum_{i=\ell}^{n-1}\Gamma_{i}\right)\mathrm{Var}_{\pi_{n}}(f).$ (60)
Now we can prove Eq. 60 using Eq. 59. From Lemma 6.4 we have that $\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)})=\mathrm{Var}_{\pi_{k}}(f^{(k)})-\mathrm{Var}_{\pi_{k-1}}(f^{(k-1)})$, and hence Eq. 59 is equivalent to the following:
$\frac{\mathrm{Var}_{\pi_{k}}(f^{(k)})}{\left(\sum_{i=0}^{k-1}\Gamma_{i}\right)}\geq\frac{\mathrm{Var}_{\pi_{k-1}}(f^{(k-1)})}{\left(\sum_{i=0}^{k-2}\Gamma_{i}\right)}.$ (61)
Then by chaining these inequalities (with $f=f^{(n)}$ we obtain:
$\frac{\mathrm{Var}_{\pi_{n}}(f)}{\left(\sum_{i=0}^{n-1}\Gamma_{i}\right)}\geq\frac{\mathrm{Var}_{\pi_{\ell}}(f^{(\ell)})}{\left(\sum_{i=0}^{\ell-1}\Gamma_{i}\right)}.$ (62)
Finally, applying Lemma 6.4 with $i=n,j=\ell$, we have that Eq. 62 is equivalent to Eq. 60, which completes the proof of Theorem 6.1. ∎
6.7 Proofs of Basic Facts for Dirichlet Form and Variance
Here we provide the missing proofs of Lemmas 6.4 and 6.5. We begin with the proof of Lemma 6.4.
Proof of Lemma 6.4.
Without loss of generality we can assume that $\mathbb{E}_{\pi_{i}}[f^{(i)}]=\mathbb{E}_{\sigma\sim\pi_{i}}[f^{(i)}(\sigma)]=0$ and hence:
$\mathrm{Var}_{\pi_{i}}(f^{(i)})=\sum_{\sigma\in\mathcal{P}_{i}}\pi_{i}(\sigma)f^{(k)}(\sigma)^{2}.$ (63)
Since we have $\mathbb{E}_{\pi_{i}}[f^{(i)}]=0$, it follows that $\mathbb{E}_{\eta\sim\pi_{i-1}}[f^{(i-1)}(\eta)]=\mathbb{E}_{\sigma\sim\pi_{i}}[f^{(i)}(\sigma)]=0$ and hence by induction we have that $\mathbb{E}_{\eta\sim\pi_{j}}[f^{(j)}(\eta)]=0$.
Now we can consider the variance at level $j$:
$\mathrm{Var}_{\pi_{j}}(f^{(j)})$ $=\sum_{\eta\in\mathcal{P}_{j}}\pi_{j}(\eta)f^{(j)}(\eta)^{2}$
$=\sum_{\eta\in\mathcal{P}_{j}}\pi_{j}(\eta)\left[\sum_{\sigma\in\mathcal{P}_{i}}\mathrm{P}_{j,i}^{\uparrow}(\eta,\sigma)f^{(i)}(\sigma)\right]^{2}$
$=\sum_{\eta\in\mathcal{P}_{j}}\pi_{j}(\eta)\sum_{\tau_{1},\tau_{2}}\mathrm{P}_{j,i}^{\uparrow}(\eta,\eta\cup\tau_{1})f^{(i)}(\eta\cup\tau_{1})\mathrm{P}_{j,i}^{\uparrow}(\eta,\eta\cup\tau_{2})f^{(i)}(\eta\cup\tau_{2})$
$=\sum_{\eta\in\mathcal{P}_{j}}\pi_{j}(\eta)\sum_{\tau_{1},\tau_{2}}\frac{\pi_{i}(\eta\cup\tau_{1})\pi_{i}(\eta\cup\tau_{2})}{(\pi_{j}(\eta)\times(j+1)\times\dots\times n)^{2}}f^{(i)}(\eta\cup\tau_{1})f^{(i)}(\eta\cup\tau_{2})$
$=\sum_{\begin{subarray}{c}\sigma_{1},\sigma_{2}\in\mathcal{P}_{i}:\\ \sigma_{1}\cap\sigma_{2}=\eta\in\mathcal{P}_{j}\end{subarray}}\frac{\pi_{i}(\sigma_{1})\pi_{i}(\sigma_{2})}{\pi_{j}(\eta)\times(j+1)\times\dots\times n)^{2}}f^{(i)}(\sigma_{1})f^{(i)}(\sigma_{2})$
$=\sum_{\sigma_{1},\sigma_{2}\in\mathcal{P}_{i}}\pi_{i}(\sigma_{1})\mathrm{P}_{i,j}^{\vee}(\sigma_{1},\sigma_{2})f^{(i)}(\sigma_{1})f^{(i)}(\sigma_{2}),$ (64)
in the third line, let $S$ denote the set of vertices where $\eta$ is an assignment $\eta:S\rightarrow\{0,1\}$, then $\tau_{1},\tau_{2}$ are assignments $\tau_{1}:T_{1}\rightarrow\{0,1\}$ and $\tau_{2}:T_{2}\rightarrow\{0,1\}$ where $T_{1},T_{2}\subset V\setminus S,|T_{1}|=|T_{2}|=i-j$ and $T_{1}\cap T_{2}=\emptyset$.
We can complete the proof by decomposing the Dirichlet form of the down-up walk into two terms as follows:
$\mathcal{E}_{\mathrm{P}_{i,j}^{\vee}}(f^{(i)})$ $=\frac{1}{2}\sum_{\sigma_{1},\sigma_{2}\in\mathcal{P}_{i}}\pi_{i}(\sigma_{1})\mathrm{P}_{i,j}^{\vee}(\sigma_{1},\sigma_{2})(f^{(i)}(\sigma_{1})-f^{(i)}(\sigma_{2}))^{2}.$
$=\sum_{\sigma\in\mathcal{P}_{i}}\pi_{i}(\sigma)f^{(i)}(\sigma)^{2}-\sum_{\sigma_{1},\sigma_{2}\in\mathcal{P}_{i}}\pi_{i}(\sigma_{1})\mathrm{P}_{i,j}^{\vee}(\sigma_{1},\sigma_{2})f^{(i)}(\sigma_{1})f^{(i)}(\sigma_{2})$
$=:\mathrm{Var}_{\pi_{i}}f^{(k)}-\mathrm{Var}_{\pi_{j}}f^{(j)}$ by Eqs. 63 and 64
where the second line follows from the reversibility condition $\pi_{i}(\sigma)\mathrm{P}_{i}^{\vee}(\sigma,\tau)=\pi_{i}(\tau)\mathrm{P}_{i}^{\vee}(\tau,\sigma)$ and the fact that $\sum_{\tau\in\mathcal{P}_{i}}\mathrm{P}_{i}^{\vee}(\sigma,\tau)=1$. This completes the proof of the lemma. ∎
We now prove Lemma 6.5. We will use the following identity in the proof of Lemma 6.5.
Claim 6.6.
$\mathrm{Var}_{\pi_{k+1}}(f^{(k+1)})-\mathrm{Var}_{\pi_{k-1}}(f^{(k-1)})=\sum_{\tau\in\mathcal{P}_{k-1}}\pi_{k-1}(\tau)\mathrm{Var}_{\pi_{\tau,2}}(f_{\tau}^{(2)}).$ (65)
Proof.
Let us begin by examining each of the terms in the LHS of Eq. 65. Once again we can assume $\mathbb{E}_{\pi_{k+1}}[f^{(k+1)}]=\mathbb{E}_{\pi_{k-1}}[f^{(k-1)}]=0$.
$\mathrm{Var}_{\pi_{k+1}}(f^{(k+1)})=\sum_{\sigma\in\mathcal{P}_{k+1}}\pi_{k+1}(\sigma)f^{(k+1)}(\sigma)^{2}$ (66)
And for the variance at level k1k-1 we have:
$\mathrm{Var}_{\pi_{k-1}}(f^{(k-1)})=\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)f^{(k-1)}(\eta)^{2}$ (67)
Now let us analyze the variance of the local walk on the RHS of Eq. 65. First note that for $\eta\in\mathcal{P}_{k-1}$ where $\eta:S\rightarrow\{0,1\}$ for $S\subset V$, for $\tau:T\rightarrow\{0,1\}$ where $T\subset V\setminus S,|T|=2$, then $f_{\eta}^{(2)}(\tau)=f^{(k+1)}(\eta\cup\tau)$. In the following equation array, the sum over $\tau$ is over all $T\subset V\setminus S,|T|=S$, where $\eta$ is on $S$, and all $\tau:T\rightarrow\{0,1\}$.
$\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathrm{Var}_{\pi_{\eta,2}}(f_{\eta}^{(2)})$
$=\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{\tau}\pi_{\eta,2}(\tau)(f^{(k+1)}(\eta\cup\tau)-f^{(k-1)}(\eta))^{2}$
$=\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{\tau}\pi_{\eta,2}(\tau)f^{(k+1)}(\eta\cup\tau)^{2}-2\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)f^{(k-1)}(\eta)\sum_{\tau}\pi_{\eta,2}(\tau)f^{(k+1)}(\eta\cup\tau)$
$\hskip 36.135pt+\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)f^{(k-1)}(\eta)^{2}\sum_{\tau}\pi_{\eta,2}(\tau)$
$=\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{\tau}\pi_{\eta,2}(\tau)f^{(k+1)}(\eta\cup\tau)^{2}-\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)f^{(k-1)}(\eta)^{2}$
$=\sum_{\sigma\in\mathcal{P}_{k+1}}\pi_{k+1}(\sigma)f^{(k+1)}(\sigma)^{2}-\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)f^{(k-1)}(\eta)^{2}$
$=\mathrm{Var}_{\pi_{k+1}}(f^{(k+1)})-\mathrm{Var}_{\pi_{k-1}}(f^{(k-1)}),$
where we used the following two identities:
$\mbox{for every $\eta\in\mathcal{P}_{k-1}$, }f^{(k-1)}(\eta)=\sum_{\tau}\pi_{\eta,2}(\tau)f^{(k+1)}(\eta\cup\tau)$ (68)
$\mbox{for every $\sigma\in\mathcal{P}_{k+1}$, }\sum_{\begin{subarray}{c}\eta\in\mathcal{P}_{k-1}:\\ \eta\subset\sigma\end{subarray}}\pi_{k-1}(\eta)\pi_{\eta,2}(\sigma\setminus\eta)=\pi_{k+1}(\sigma)$ (69)
This completes the proof of the claim. ∎
Now we can complete the proof of Lemma 6.5.
Proof of Lemma 6.5.
Recall, our aim is to prove, that for any $f^{(k+1)}:\mathcal{P}_{k+1}\rightarrow\mathbb{R}$, the following holds:
$\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{k+1})\geq(2\gamma_{k-1}-1)\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)}).$ (58)
Fix $\tau\in\mathcal{P}_{k-1}$. Note that by definition of $\gamma_{j}$ we have that the spectral gap of the local walk $Q$ satisfies $\gamma(Q_{\tau})\geq\gamma_{k-1}$ and hence $\gamma(\mathrm{P}_{\tau,1}^{\wedge})\geq\gamma_{k-1}/2$; therefore, by Lemma 5.4, we also have that $\gamma(\mathrm{P}_{\tau,2}^{\vee})\geq\gamma_{k-1}/2$ which means that for any $f$:
$\mathcal{E}_{\mathrm{P}_{\tau,2}^{\vee}}(f_{\tau}^{(2)})\geq\frac{1}{2}\gamma_{k-1}\mathrm{Var}_{\pi_{\tau,2}}(f_{\tau}^{(2)}).$ (70)
Lemma 6.4 yields that $\mathcal{E}_{\mathrm{P}_{\tau,2}^{\vee}}(f^{(2)}_{\tau})=\mathrm{Var}_{\tau,2}(f^{(2)}_{\tau})-\mathrm{Var}_{\tau,1}(f^{(j)}_{\tau})$, and hence combining with Eq. 70 we have the following:
$\mathrm{Var}_{\pi_{\tau,2}}(f_{\tau}^{(2)})$ $\geq\frac{1}{1-\gamma_{k-1}/2}\mathrm{Var}_{\pi_{\tau,1}}(f_{\tau}^{(1)})$
$\geq 2\gamma_{k-1}\mathrm{Var}_{\pi_{\tau,1}}(f_{\tau}^{(1)}).$ (71)
We can now complete the proof of Eq. 58:
$\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)})$ $=\mathrm{Var}_{\pi_{k+1}}(f^{(k+1)})-\mathrm{Var}_{\pi_{k-1}}(f^{(k-1)})$ by Lemma 6.4
$=\sum_{\tau\in\mathcal{P}_{k-1}}\pi_{k-1}(\tau)\mathrm{Var}_{\pi_{\tau,2}}(f_{\tau}^{(2)})$ by Claim 6.6
$\geq 2\gamma_{k-1}\sum_{\tau\in\mathcal{P}_{k-1}}\pi_{k-1}(\tau)\mathrm{Var}_{\pi_{\tau,1}}(f_{\tau}^{(1)})$ by Eq. 71
$=2\gamma_{k-1}\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)}),$ by Claim 5.7
which proves that $\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})\geq(2\gamma_{k-1}-1)\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f^{(k)})$, and hence proves Eq. 58 and Lemma 6.5.
If we used the bound $1/(1-\gamma_{k-1}/2)$ in (71) instead of $2\gamma_{k-1}$ then the statement of Lemma 6.5 has $\gamma_{k-1}/(2-\gamma_{k-1})$ in place of $(2\gamma_{k-1}-1)$; this matches the statement of [CLV21a, Fact A.8 and Theorem A.9]. ∎
6.8 Optimal Mixing Time via Entropy Decay
In the previous section we established Theorem 1.5 which says that spectral independence implies optimal $O(n)$ relaxation time of the Glauber dynamics. In contrast, Theorem 1.8 says that spectral independence and marginal boundedness implies optimal $O(n\log{n})$ mixing time. Both Theorems 1.5 and 1.8 were proved by Chen, Liu, and Vigoda [CLV21a], and our proofs of these theorems follow their proof approach. To obtain Theorem 1.8, we follow the same proof approach as in the proof of Theorem 1.5 with the general objective of replacing the variance functional $\mathrm{Var}$ by the entropy functional $\mathrm{Ent}$.
We first recall the definition of entropy of a functional.
Definition 6.7 (Entropy).
The entropy of a function $f:\Omega\rightarrow\mathbb{R}$ with respect to $\mu$ is the following:
$\mathrm{Ent}[f]:=\mu[f\log f]-\mu[f]\log\left(\mu[f]\right)=\sum_{\sigma\in\Omega}\mu(\sigma)f(\sigma)\log(f(\sigma))-\sum_{\sigma\in\Omega}\mu(\sigma)f(\sigma)\log[\sum_{\eta\in\Omega}\mu(\eta)f(\eta)].$
For a vertex $x$, denote the expected entropy at $x$ by:
$\mathrm{Ent}_{x}[f]=\sum_{\tau}\mu(\tau)\mathrm{Ent}^{\tau}_{x}[f]=\sum_{\tau}\mu(\tau)\mathrm{Ent}[F\mid\sigma(V\setminus\{x\})=\tau(V\setminus\{x\}].$
We utilize entropy tensorization, which is the analog of approximate tensorization of variance, see Definition 3.4.
Definition 6.8 (Entropy Tensorization).
A distribution $\mu$ with support $\Omega\subseteq\{0,1\}^{V}$ satisfies approximate tensorization of entropy with constant $C>0$, if for all $f:\Omega\to\mathbb{R}_{\geq 0}$ we have that
$\mathrm{Ent}(f)\leq C\sum_{v\in V}\mu\left(\mathrm{Ent}_{v}(f)\right).$
Note, the corresponding notion of variance tensorization is equivalent to the spectral gap of the (heat-bath) Glauber dynamics as the RHS in the above is equivalent to the Dirichlet form, see Corollary 3.5.
Entropy tensorization with constant $C$ yields an optimal mixing time bound for the Glauber dynamics by bounding the decay rate of entropy of the Glauber dynamics with respect to the Gibbs distribution, see [CMT15, Cap23]:
$T_{\mathrm{mix}}\leq Cn\log(\log(1/\mu^{*})).$ (72)
As suggested above, to establish approximate tensorization of entropy we aim to replace $\mathrm{Var}$ by $\mathrm{Ent}$ in the proof of Theorem 1.5. The first is that the proof of Theorem 1.5 uses the spectral gap of the Glauber dynamics. Recall the formulation of the spectral gap in the Poincaré inequality, see Eq. 15, where the spectral gap $\gamma=\min_{f}\mathcal{E}(f)/\mathrm{Var}(f)$. Moreover, for the down-up chain $\mathrm{P}_{i,j}^{\vee}$ one can replace the $\mathcal{E}_{\mathrm{P}_{i,j}^{\vee}}(f)$ by $\mathrm{Var}_{\pi_{i}}(f)-\mathrm{Var}_{\pi_{j}}(f)$, see Lemma 6.4. Notice that in Section 6 for the proof of Theorem 1.5 we only considered the down-up chain and hence we can rephrase the proof only in terms of $\mathrm{Var}(f)$, without use of $\mathcal{E}(f)$. Therefore we can aim to replace all occurrences of $\mathrm{Var}(f)$ with $\mathrm{Ent}(f)$ in order to establish entropy decay instead of variance decay.
The only technical obstacle in replacing $\mathrm{Var}$ by $\mathrm{Ent}$ is in the application of Theorem 1.4 in the proof of Theorem 1.5 in Section 6.5; in particular, see Eq. 54. Theorem 1.4 established a lower bound on the spectral gap of the Glauber dynamics but it does not establish a bound on its entropy decay. To get an analgous bound on entropy decay we need to use an additional property known as marginal boundedness, see Definition 1.7. Using $b$-marginal boundedness, [CLV21a, Lemma 4.2] proved a corresponding bound on $\mathrm{Ent}_{T}(f)$ in terms of $\sum_{v\in T}\mu(\mathrm{Ent}_{v}(f))$ which depends on $b$ and $|T|$:
$\mathrm{Ent}_{T}(F|\tau)\leq C^{\prime}\sum_{v\in T}\mu_{T}^{\tau}(\mathrm{Ent}_{v}[F|\tau]).$ (73)
In [CLV21a, Lemma 4.2] they established $C^{\prime}=C^{\prime}(k,b)=\frac{3k^{2}\log(1/b)}{2b^{2k+2}}$ where $k=|T|$ (note $b$ typically depends on $\Delta$, and also depends on $\lambda$ or $\eta$). Below, we provide a simple proof that $C^{\prime}=C^{\prime}(|T|,b,\Delta,\eta)=2C(\eta,\Delta)k\log(1/b)$ where $C(\eta,\Delta)$ is the constant from Theorem 1.5. The bound in Eq. 73 (with either constant $C^{\prime}$) can be used in the analogous step of Eq. 54 to instead establish entropy decay, and then conclude approximate tensorization of entropy of the Gibbs distribution (see Definition 6.8 above) which implies an optimal $O(n\log{n})$ bound on the mixing time, via Eq. 72.
This outlines the approach of [CLV21a] to obtain optimal mixing time bounds. Further improvements are obtained in [CFYZ22, CE22]; these works upper bound the mixing time by $C(\eta)n\log{n}$ instead of $C(\eta,\Delta)n\log{n}$ based on a stronger form of spectral independence. The work of Chen, Feng, Yin, and Zhang [CFYZ22] introduce the field dynamics which is presented in Section 10. In contrast, the work of Chen and Eldan [CE22] apply the framework of stochastic localization.
Proof of Eq. 73.
Consider the graph $G=(V,E)$ where $T\subseteq V$. Let $k=|T|$. We are running the Glauber dynamics on $T$ where for every $v\notin T$ the configuration at $v$ is fixed to $\tau(v)$. Let $\mu=\mu_{\tau}$ denote the Gibbs distribution on $T$ with fixed configuration $\tau$ outside $T$.
From Diaconis and Saloff-Coste [DS96] (see the proof of Corollary A.4), we know the following inequality:
$\mathrm{Ent}_{\mu}(f)\leq\left[\frac{\log(1/\mu^{*})-1}{1-2\mu^{*}}\right]\mathrm{Var}_{\mu}(\sqrt{f}),$ (74)
where $\mu^{*}=\min_{\sigma\in\Omega}\mu(\sigma)$. From Theorem 1.5 and Corollary 3.5, we have approximation tensorization of variance:
$\mathrm{Var}_{\mu}(f)\leq C(\eta,\Delta)\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)].$ (75)
We will use the following elementary fact (e.g., see [DS96]), which we will prove momentarily:
$\mathrm{Var}_{v}(\sqrt{f})\leq\mathrm{Ent}_{v}(f).$ (76)
Therefore, by combining Eq. 74 and Eq. 75, and since $\mu^{*}\geq b^{k}$ we have:
$\mathrm{Ent}_{\mu}(f)$ $\leq C(\eta,\Delta)\left[\frac{\log(1/\mu^{*})-1}{1-2\mu^{*}}\right]\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(\sqrt{f})]$ by Eq. 74 (77)
$\leq 2C(\eta,\Delta)k\log(1/b)\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(\sqrt{f})]$ by Eq. 75 (78)
$\leq 2C(\eta,\Delta)k\log(1/b)\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Ent}_{v}(f)]$ by Eq. 76. (79)
This completes the proof of Eq. 73 with constant $C^{\prime}=2C(\eta,\Delta)k\log(1/b)$.
For completeness we now prove Eq. 76. Without loss of generality, we assume $f$ satisfies $\mathbb{E}_{\mu}[f]=1$. We then have the following:
$\mathrm{Ent}_{\mu}(f)$ $=\mathbb{E}_{\mu}[f\log(f)]$
$=2\mathbb{E}_{\mu}[f\log(\sqrt{f})]$
$\geq 2\mathbb{E}_{\mu}\left[f\left(1-\frac{1}{\sqrt{f}}\right)\right]$ (80)
$=2(1-\mathbb{E}_{\mu}[\sqrt{f}])$
$\geq 1-(\mathbb{E}_{\mu}[\sqrt{f}])^{2}$ (81)
$=\mathrm{Var}_{\mu}(\sqrt{f}),$
where Eq. 80 follows from the fact that $\log(x)=\log(1+x-1)\leq x-1$ and hence $\log(1/x)\geq 1-x$, and Eq. 81 follows from $2(1-x)\geq 1-x^{2}$. ∎
7 Matroids
We now utilize the spectral independence approach to establish fast mixing of a Markov chain for generating a random basis of a matroid. The main result for matroids is the work of Anari, Liu, Oveis Gharan, and Vinzant [ALOV19]. In fact, the work of [ALOV19] for matroids inspired the idea of spectral independence as we presented earlier and introduced in [ALO20], however we present these results in the opposite order.
The general approach to prove fast mixing of the chain on bases of a matroid is to apply the Random Walk Theorem (Theorem 4.1). To bound the spectral gap of the local walks we will use a result of Oppenheim [Opp18] known as the Trickle-Down Theorem. The Trickle-Down Theorem bounds the spectral gap of the local walk $Q_{S}$ for a pinning $S$ by induction on $|S|$. The base case will be a bound on the gap of the local walk $Q_{S}$ when k:=|S|=n2k:=|S|=n-2 which will correspond to rank-2 matroids; this is a significantly easier case to analyze as it corresponds to an unweighted walk. We will then use the Trickle-Down Theorem to deduce a bound on the spectral gap of $Q_{S}$ for pinnings of size k1k-1 and hence any pinning $S$ by induction.
7.1 Matroids: Definitions
A matroid $\mathcal{M}=(E,\mathcal{I})$ consists of the following:
  • Ground Set denoted by $E$;
  • Collection $\mathcal{I}$ of subsets of the ground set. We refer to each $S\in\mathcal{I}$ as an independent set.
To be a matroid we require the following properties:
P1: Downward closure:
Every subset of an independent set is also an independent set. This is means that $\mathcal{I}$ is downward closed, or equivalently that $\mathcal{I}$ forms a simplicial complex. To be clear, this property says that for every $S\in\mathcal{I}$, if $T\subset S$ then $T\in\mathcal{I}$.
P2: Exchange property:
If $S\in\mathcal{I}$ and $T\in\mathcal{I}$, and if $|S|<|T|$ then there exists $e\in S\setminus T$ where $S\cup e\in I$.
Note, the empty set $\emptyset$ is always an independent set by the downward closure property. A set $R\subset E$ which is not an independent set is called dependent.
One important consequence of the exchange axiom is that all maximal independent sets are of the same size (maximal independent set means it is not contained in another independent set).
Definition 7.1.
Consider a matroid $\mathcal{M}=(E,\mathcal{I})$. A basis is a maximal independent set. In words, a basis is an independent set $S\in\mathcal{I}$ where for all $e\not\in S$, we have that $S\cup\{e\}\not\in\mathcal{I}$. Note that the set of bases uniquely determines a matroid. We will use $\mathcal{F}$ to denote the collection of bases of a matroid.
All bases of a matroid have the same size which is called the rank of the matroid $\mathcal{M}$.
Definition 7.2 (Truncation).
Let $M$ be a matroid. Let $k$ be a positive integer. The set $\{S\in\mathcal{I}:|S|\leq k\}$ is called the truncation of $M$. Note that the truncation of a matroid is also a matroid.
We state a few exercises that detail important, fundamental properties of matroids.
Exercise 7.3.
Let $\mathcal{F}$ be a set of bases of a matroid $\mathcal{M}$. Let $\mathcal{F}^{*}$ be the set of complements, that is, $\mathcal{F}^{*}=\{[n]-B:B\in\mathcal{F}\}$. Show that $\mathcal{F}^{*}$ are the bases of a matroid $\mathcal{M}^{*}$ (called the dual of the matroid $\mathcal{M}$).
Exercise 7.4.
Let $\mathcal{I}$ be a set of independent sets of a matroid $M=(E,\mathcal{I})$. Let $S\subseteq[n]$. Let $\mathcal{I}^{\prime\prime}$ be the set of independent sets contained in $S$, that is, $\{T:T\in\mathcal{I}\ \mbox{and}\ T\subseteq S\}$. Show that $\mathcal{M}^{\prime\prime}=(E,\mathcal{I}^{\prime\prime})$ is a matroid (called the restriction of the matroid $\mathcal{M}$; denoted by $\mathcal{M}|S$).
Exercise 7.5.
Let $\mathcal{I}$ be a set of independent sets of a matroid $\mathcal{M}=(E,\mathcal{I})$. Let $S\subseteq[n]$. Let $\mathcal{I}^{\prime}$ be the set of independent sets containing $S$ with $S$ removed, that is, $\{T\setminus S:T\in\mathcal{I}\ \mbox{and}\ S\subseteq T\}$. Show that $\mathcal{M}^{\prime}=(E,\mathcal{I}^{\prime})$ is a matroid (called the contraction of the matroid $\mathcal{M}$; denoted by $\mathcal{M}/S$).
Remark 7.6.
A matroid $\mathcal{M}^{\prime}$ that can be obtained from $\mathcal{M}$ by contraction and restriction is called a minor of $\mathcal{M}$.
7.2 Matroids: Examples
Here are some natural examples of matroids.
Graphic matroid:
Let $G=(V,E)$ be a connected graph. Let $\mathcal{I}=\{S\subset E:(V,S)\mbox{ is acyclic}\}$ be the collection of all forests of $G$. The collection $\mathcal{M}=(E,\mathcal{I})$ is a matroid. The bases of $\mathcal{M}$ are maximal acyclic subgraphs. Since $G$ is assumed to be connected then the bases of $\mathcal{M}$ are the spanning trees of $G$ since $G$, and hence this is also called the spanning tree matroid.
Transversal Matroid:
Let $G=(V,E)$ be a bipartite graph where $V=L\cup R$ is the bipartition. Let $\mathcal{I}$ denote the subsets of $L$ that are the endpoints of some matching of $G$. Note, $S\in\mathcal{I}$ is a set of vertices and they can correspond to multiple matchings.
The collection $\mathcal{M}=(L,\mathcal{I})$ is a matroid. The bases are the subsets of $L$ covered by a maximum matching.
Linear matroid:
Let $V=\{v_{1},\ldots,v_{n}\}$ be vectors in some vector space. Let
$\mathcal{I}=\{S\subseteq V:S\text{ is linear independent}\}.$
Then $\mathcal{M}=(V,\mathcal{I})$ is a matroid.
7.3 Bases-exchange Walk
Our central question is, for a given matroid $\mathcal{M}=(E,\mathcal{I})$, can we generate a basis, uniformly at random from the set $\mathcal{F}$ of all bases, in time polynomial in $n=|E|$. We are given the matroid in implicit form, so we can efficiently check whether a given subset $S\subset E$ is independent or dependent. We will use the following natural Markov chain known as the bases-exchange walk to generate a random bases.
The bases-exchange walk is the following Markov chain $(B_{t})$ on the set of bases of a given matroid $\mathcal{M}$. Let $\mathcal{F}$ denote the collection of bases of a matroid $M$. From $B_{t}\in\mathcal{F}$, the transitions $B_{t}\rightarrow B_{t+1}$ are defined as follows:
  • Choose an element $e\in B_{t}$ uniformly at random.
  • Let $F=\{f\in E:B_{t}\cup\{f\}\setminus\{e\}\in\mathcal{I}\}$ denote the set of edges that we can add to $B_{t}\setminus e$ while remaining independent.
  • Choose an element $f\in F$ uniformly at random. Let $B_{t+1}=B_{t}\cup\{f\}\setminus\{e\}$.
Note the edge $e$ is in $F$ and hence $P(B_{t},B_{t})>0$; thus, the chain is aperiodic.
Exercise 7.7.
Using the exchange-axiom, prove by induction that the chain is irreducible.
Therefore, the bases-exchange walk is an ergodic Markov chain. Since the chain is symmetric then the unique stationary distribution is uniform over the set $\mathcal{F}$ of bases of the matroid $M$. In the terminology from the spectral independence section, the bases-exchange walk is equivalent to the down-up walk on the bases of a matroid.
7.4 Main Theorem: Fast Mixing of Bases-Exchange Walk
The main result is a bound on the spectral gap of the bases-exchange walk.
See 1.15
By proving entropy contraction, [CGM21] improves the mixing time bound to $O(r(M)\log{\log(|\mathcal{F}|)})=O(r(M)\log r(M)+r(M)\log{\log{n}})$. Furthermore, the $\log\log n$ term was subsequently removed in [ALO+21], achieving $O(r(M)\log r(M))$ mixing time. This is tight because we need $\Omega(r(M)\log r(M))$ time to replace every element due to coupon collector.
7.5 Application: Reliability
Before proving fast convergence of the bases-exchange walk we discuss an important application for sampling/approximate counting bases of a matroid.
An interesting application is to compute reliability. Let $X\subseteq[n]$ be a random set where every element is included independently with probability $p$. The reliability is defined as the $\mathrm{Pr}(X\supseteq\text{basis})$, and the unreliability is defined as $\mathrm{Pr}(X\not\supseteq\text{basis})$.
For a graphical matroid these problems are known as network reliability/unreliability as the problem corresponds to whether $X$ is connected/disconnected. Karger [Kar99] presented an $\mathsf{FPRAS}$ for network unreliability, and Guo and Jerrum [GJ19] presented an $\mathsf{FPRAS}$ for network reliability.
Theorem 1.15 yields an $\mathsf{FPRAS}$ for the reliability of a matroid. In particular, rapid mixing of the bases-exchange walk yields an efficient sampling algorithm for generating a basis of a matroid (almost) uniformly at random; the classical result of [JVV86] then yields an $\mathsf{FPRAS}$ for computing the number of bases.
Consider a matroid $\mathcal{M}$ for which we would like to compute the reliability, which is the probability that the random subset $X$ contains a basis of $\mathcal{M}$. Let $\mathcal{M}^{*}$ denote the dual of $\mathcal{M}$, see 7.3; note $\mathcal{M}^{*}$ is a matroid. Note,
[…]
and we can obtain an $\mathsf{FPRAS}$ for the number of independent sets (of a matroid $\mathcal{M}^{*}$) of size $k$ since the truncation of a matroid is a matroid. This yields an $\mathsf{FPRAS}$ for the reliability of the matroid $\mathcal{M}$.
8 Trickle-Down Theorem
Here we present Oppenheim’s Trickle-Down Theorem [Opp18] which is the new tool needed to prove Theorem 1.15. We present the theorem in the general context of simplicial complexes; this enables one to apply it both in the context of spin systems and for bases of a matroid.
8.1 Simplical Complexes
Let $\Lambda$ be a finite set. Let $\Omega$ denote a collection of subsets of $\Lambda$ that is 1) closed under taking subsets (that is, if $S\in\Omega$ and $Z\subseteq S$ then $Z\in\Omega$), and 2) all maximal (under inclusion) sets in $\Lambda$ have the same size $r$. Such a pair $(\Lambda,\Omega)$ is called a pure abstract simplicial complex.
For the example of spectral independence on a (binary) spin system, the set $\Lambda=V\times\{0,1\}$ for an input graph $G=(V,E)$. Here $r=|V|$ while $|\Lambda|=2r$. The distribution $\mu$ is the Gibbs distribution on assignments $\sigma\in\{0,1\}^{V}$. Now we construct $\Omega$. The maximal sets in $\Omega$ are the support of $\mu$. Defining the distributions $\pi_{k}$ as in Section 5.1, then the union of the supports of $\pi_{0},\dots,\pi_{r}$ yields $\Omega$.
For the example of bases of a matroid, let $\Lambda=E$, the ground set of the matroid $\mathcal{M}=(E,\mathcal{I})$. Then let $\Omega=\mathcal{I}$ be the independent sets of $\mathcal{M}$. The maximal sets in $\Omega$ are the bases of $\mathcal{M}$ and hence $r$ is the rank of the matroid. All bases in $\Omega$ have the same weight, thus we set $\mu$ to be the uniform distribution over the set of bases $\mathcal{F}$. Defining the distributions $\pi_{k}$ as in Section 5.1, then the union of the supports of $\pi_{0},\dots,\pi_{r}$ is $\Omega$.
More generally, let $\mu$ be a distribution on sets of size $r$. Define the distributions $\pi_{k}$ and the up/down chains exactly as in Section 5.1. The union of the supports of distributions $\pi_{0},\pi_{1},\dots,\pi_{r}$ is a pure abstract simplicial complex.
8.2 Trickle-Down Statement
Define the distributions $\pi_{k}$ and the up/down chains exactly as in Section 5.1, and the local walks $Q_{S}$ the same as in Section 4.3. Let $(\Lambda,\Omega)$ be the simplicial complex that is the support of $\pi_{k}$’s. We can now state the Trickle-Down theorem.
Theorem 8.1 ([Opp18]).
Consider $S\in\Omega$. Let $i=|S|$ and assume $0\leq i<r-2$. Let $\gamma_{i}=\gamma(Q_{S})$. Assume that $Q_{S}$ is irreducible and hence $\gamma_{i}>0$. Suppose there exists $\gamma_{i+1}>0$ where for all $Z\in\Omega$ such that $S\subset Z$ and $|Z|=|S|+1$,
$\gamma(Q_{Z})\geq\gamma_{i+1}.$ (82)
Then the following holds:
$\gamma_{i}\geq 2-\frac{1}{\gamma_{i+1}}.$
As a consequence, if there is optimal spectral gap, namely, $\gamma(Q_{S^{\prime}})\geq 1$, for all $S^{\prime}\in\Omega$ where $|S^{\prime}|=r-2$, then by induction we obtain, via the Trickle-Down Theorem, that there is optimal spectral gap of $\gamma(Q_{S})\geq 1$ for all $S\in\Omega$.
Corollary 8.2 (Trickle-Down Without Loss).
If for all $S^{\prime}\in\Omega$ with $|S^{\prime}|=r-2$ we have $\gamma(Q_{S^{\prime}})\geq 1$. Further assume that for all $S\in\Omega$ we have that $Q_{S}$ is irreducible. Then for all $S\in\Omega$ it holds that
$\gamma(Q_{S})\geq 1.$
8.3 Rapid Mixing of Bases-Exchange Walk: Proof of Theorem 1.15
The proof of Theorem 1.15 uses the Trickle-down Theorem (Theorem 8.1), which bounds the spectral gap of the local walks, and the Random Walk Theorem (Theorem 4.1), which is the local to global theorem.
Our goal is to bound the spectral gap of the bases-exchange walk using the Random Walk Theorem. To that end we need to bound the spectral gap of the local walk (this is the up-down walk $\mathrm{P}_{1}^{\wedge}$) for every pinning (or link in the terminology of simplicial complexes). The pinnings are defined for every independent set $S$. In particular, for an independent set $S$ the pinning is defined as $\{B\setminus S:S\subseteq B,B\in\mathcal{I}\}$, which is a contraction of the matroid. 7.5 establishes that the contraction of a matroid is also a matroid.
These local walks for a particular pinning are weighted as these are the projections in the original simplicial complex. To handle these weighted local walks we utilize the Trickle-down Theorem. The Trickle-down Theorem says that if we have an “optimal” bound on the spectral gaps for all local walks on (r2)(r-2)-dimensional pinning/link then we obtain an optimal bound on the spectral gap for the local walks for all pinnings/links, and then we can apply the Random Walk Theorem. This greatly simplifies matters because the local walks on (r2)(r-2)-dimensional pinnings/links are unweighted, and hence they simply correspond to (unweighted) rank 2 matroids.
Lemma 8.3.
For a rank 2 matroid $\mathcal{M}$, the spectral gap of the local walk satisfies
$\gamma(Q_{\mathcal{M}})\geq 1.$
Proof.
Consider a rank 2 matroid $\mathcal{M}$. Consider a graph $G_{\mathcal{M}}$ whose vertex set are the elements of the matroid $\mathcal{M}$ and whose edges are the bases of the matroid $\mathcal{M}$. We will argue that this graph consists of a complete multipartite graph plus some isolated vertices. The isolated vertices $v$ are those for which $\{v\}$ is not independent. Remove the isolated vertices. Now, by the Exchange Axiom, for any three vertices $i,j,k$ with $\{j,k\}\in E$, either $\{i,j\}\in E$ or $\{i,k\}\in E$. Equivalently, if $\{i,j\}\not\in E$ and $\{i,k\}\not\in E$ then $\{j,k\}\not\in E$. This means that we have a multipartite graph (where the partitions are maximal sets of pairwise dependent elements).
The adjacency matrix $A$ of a complete multipartite graph is the all ones matrix (that has only one non-zero eigenvalue) minus a block diagonal matrix, so it has the second eigenvalue $\leq 0$. (Adding the isolated vertices does not change the non-zero part of the spectrum.)
The local walk $Q_{\mathcal{M}}$ is equivalent to the random walk on the graph $G_{\mathcal{M}}$. The transition matrix of the random walk on $G_{\mathcal{M}}$ is the matrix $D^{-1}A$ where $D$ is the diagonal matrix with the degrees. Note, $D^{-1}A$ is similar to $D^{-1/2}AD^{-1/2}$. Hence, by Sylvester’s law of inertia, see Lemma 2.1 (which states that a change of basis does not change the number of eigenvalues of a given sign) we have that $\lambda_{2}(A)\leq 0\iff\lambda_{2}(D^{-1/2}AD^{-1/2})\leq 0$. This completes the proof. ∎
Lemma 8.4.
Let $\mathcal{M}=(E,\mathcal{I})$ be a matroid of rank $r$. For any $S\in\mathcal{I}$, $|S|\leq r-2$, we have that the spectral gap of the local walk $Q_{S}$ satisfies $\gamma(Q_{S})\geq 1$.
Proof of Lemma 8.4.
This lemma follows from Lemma 8.3 and Corollary 8.2 (since the (r2)(r-2)-links are rank $2$ matroids). ∎
We can now prove the main theorem establishing rapid mixing of the bases-exchange walk.
Proof of Theorem 1.15.
The random walk theorem Theorem 4.1 implies that the spectral gap $\gamma$ for the transition matrix $P$ of the bases-exchange walk satisfies:
$\gamma(P)\geq\frac{1}{r}\prod_{k=0}^{r-2}\gamma_{k},$ (83)
where
$\gamma_{k}=\min_{S\in\mathcal{I}:|S|=k}\gamma(Q_{S})$
and $\gamma(Q_{S})$ is the spectral gap for the local walk $Q_{S}$, which by Lemma 8.4 we have that $\gamma_{k}\geq 1$ for every $k$. ∎
8.4 Connections to Spectral Independence and Log-Concavity
Recall, the definition of the modified influence matrix $\widetilde{\Psi}$ in Definition 2.7, which is slightly different than the influence matrix $\Psi$ in Definition 1.1; the former $\widetilde{\Psi}$ leads to the notion of weak spectral independence (see Definition 2.9) and the latter $\Psi$ leads to spectral independence (see Definition 1.2). Both weak spectral independence and spectral independence upper bound the maximum eigenvalue $\lambda_{\max}$ of the corresponding matrix but do so for all valid pinnings (which corresponds to all induced subgraphs in the case of the hard-core model on independent sets).
We first observe that the local walks having an optimal spectral gap is equivalent to optimal spectral independence.
Remark 8.5.
The following are equivalent by Lemma 4.2:
  1. 1.
    For every pinning $S$, the local walk $Q_{S}$ satisfies $\gamma(Q_{S})\geq 1$.
  2. 2.
    0-spectral independence holds.
Recall, 0-spectral independence implies 0-weak spectral independence. Now we relate 0-weak spectral independence with log-concavity.
Given a distribution $\mu$ on $\{0,1\}^{n}$ consider the following generating polynomial
$f(x_{1},\dots,x_{n})=\sum_{\sigma\in\{0,1\}^{n}}\mu(\sigma)\prod_{i=1}^{n}x_{i}^{\sigma_{i}}.$
Definition 8.6 (log-concavity at a point).
We say that $f$ is log-concave at a point $x$ if $(\nabla^{2}\log f)(x)\preceq 0$.
As noted above, 0-weak spectral independence is defined as $\lambda_{\max}(\widetilde{\Psi_{\tau}})\leq 1$ for all pinnings $\tau$. The following lemma shows that $\lambda_{\max}(\widetilde{\Psi})\leq~1$ is equivalent to log-concavity of $f$.
Theorem 8.7.
Log-concavity of $f$ at the point $(1,\dots,1)$ is equivalent to the influence matrix $\widetilde{\Psi}$ having $\lambda_{\max}(\widetilde{\Psi})\leq~1$.
Proof.
For $1\leq i\neq j\leq n$ we have
$x_{i}x_{j}(\nabla^{2}\log f(1,\dots,1))_{ij}$ $=x_{i}x_{j}\left(\frac{(\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f)}{f}(1,\dots,1)-\frac{(\frac{\partial}{\partial x_{i}}f)(\frac{\partial}{\partial x_{i}}f)}{f^{2}}(1,\dots,1)\right)$ (84)
$=\mathrm{Pr}_{\sigma\sim\mu}(\sigma(i)=\sigma(j)=1)-\mathrm{Pr}_{\sigma\sim\mu}(\sigma(i)=1)\mathrm{Pr}_{\sigma\sim\mu}(\sigma(j)=1),$ (85)
and for $1\leq i\leq n$ we have
$x_{i}^{2}(\nabla^{2}\log f(1,\dots,1))_{ii}$ $=-x_{i}^{2}\left(\frac{\frac{\partial}{\partial x_{i}}f}{f}(1,\dots,1)\right)^{2}=-\mathrm{Pr}_{\sigma\sim\mu}(\sigma(i)=1)^{2}.$ (86)
Recall from Eq. 4 the definition of the covariance matrix:
$\mathrm{Cov}_{\mu}(i,j)$ $=\mathrm{Pr}_{\sigma\sim\mu}[\sigma(i)=\sigma(j)=1]-\mathrm{Pr}_{\sigma\sim\mu}[\sigma(i)=1]\cdot\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=1].$
Let $\widetilde{D}={\mathrm{diag}}(\mathbb{E}(\mu))$ be the diagonal matrix with entries $\widetilde{D}(i,i)=\mu(\sigma(i)=1)$. Hence $(\nabla^{2}f)(1,\dots,1)\preceq 0$ is equivalent to
$\mathrm{Cov}_{\mu}-\widetilde{D}=\mathrm{diag}(x_{i})(\nabla^{2}f(1,\dots,1))\mathrm{diag}(x_{i})\preceq 0,$ (87)
where in the first equality we used Eq. 84 and (86).
Recall from Eq. 13 that $\lambda_{\max}(\widetilde{\Psi})\leq 1$ is equivalent to $\mathrm{Cov}_{\mu}\preceq\widetilde{D}$. Therefore, Eq. 87 establishes the condition Eq. 12 in the definition of 0-weak spectral independence, see Definition 2.9. ∎
We refer interested readers to the works [ALOV19, BH20] for more discussion on (strongly) log-concave polynomials.
8.5 Proof of the Trickle-Down Theorem: Proof of Theorem 8.1
In this section, we complete the proof of the Trickle-Down Theorem (Theorem 8.1); this is the only remaining task to complete the proof of fast mixing of the bases-exchange walk (Theorem 1.15).
Remark 8.8.
For an alternative proof of the Trickle-Down Theorem (Theorem 8.1) in the language of covariance matrices (see Eq. 10), see [Koe24, AKV24]. We highlight that [AKV24] established a trickle-down theorem for general localization schemes, which can be applied to the analysis of the field dynamics for the hardcore model; see Section 11 for more details.
Fix $S\subset E$. Let $\pi_{S}=\pi_{S,1}$ and hence $\pi_{S}$ is a distribution over $a\in E\setminus S$. Recall that the local walk $Q_{S}$ has stationary distribution $\pi_{S}$.
We will use the following technical lemma.
Lemma 8.9.
Let $S\subset E$ where $i=|S|$. For any function $f:E\rightarrow\mathbb{R}$,
$\mathcal{E}_{Q_{S}}(f)$ $=\sum_{a\in E\setminus S}\pi_{S}(a)\mathcal{E}_{Q_{S\cup a}}(f)$ (88)
$\mathbb{E}_{\pi_{S}}(f)$ $=\sum_{a\in E\setminus S}\pi_{S}(a)\mathbb{E}_{\pi_{S\cup a}}(f).$ (89)
We can now prove the Trickle-Down Theorem.
Proof of Theorem 8.1.
Our goal is to bound $\gamma_{i}=\gamma(Q_{S})$ which is the spectral gap for the local walk $Q_{S}$ with pinning $S\subset E$. Recall, for all $f:E\rightarrow\mathbb{R}$,
$\gamma_{i}=\min_{f}\frac{\mathcal{E}_{Q_{S}}(f)}{\mathrm{Var}_{\pi_{S}}(f)},$ (90)
where the minimization is over all $f:E\rightarrow\mathbb{R}$ where $f$ is not constant on $E\setminus S$ and hence $\mathrm{Var}_{\pi_{S}}\neq 0$. Let $f^{*}$ be the function $f$ which achieves the minimum in Eq. 90, and hence
$\gamma_{i}\mathrm{Var}_{\pi_{S}}(f)=\mathcal{E}_{Q_{S}}(f).$
The minimizer $f^{*}$ of (90) has to be an “eigenvector” in the following sense (this follows more directly by viewing Dirichlet forms through eigenvalues; we include a self-contained proof).
Lemma 8.10.
For any $a\in E\setminus S$,
$\mathbb{E}_{\pi_{S\cup a}}(f^{*})=(1-\gamma_{i})f^{*}(a).$
Without loss of generality, in Eq. 90 we can restrict attention to functions $f$ where $\mathbb{E}_{\pi_{S}}[f]=0$ (since translating $f$ by a constant function does not change the numerator and the denominator), and hence for $f^{*}$ we can also assume that $\mathbb{E}_{\pi_{S}}[f^{*}]=0$.
$\gamma_{i}\mathrm{Var}_{\pi_{S}}(f^{*})$ $=\mathcal{E}_{Q_{S}}(f^{*})$
$=\sum_{a\in E\setminus S}\pi_{S}(a)\mathcal{E}_{Q_{S\cup a}}(f^{*})$ by Eq. 88
$\geq\gamma_{i+1}\sum_{a\in E\setminus S}\pi_{S}(a)\mathrm{Var}_{\pi_{S\cup a}}(f^{*})$ by Eq. 82
$=\gamma_{i+1}\sum_{a}\pi_{S}(a)\mathbb{E}_{\pi_{S\cup a}}[(f^{*})^{2}]-\gamma_{i+1}\sum_{a}\pi_{S}(a)[\mathbb{E}_{\pi_{S\cup a}}(f^{*})]^{2}$
$=\gamma_{i+1}\sum_{a}\pi_{S}(a)\mathbb{E}_{\pi_{S\cup a}}[(f^{*})^{2}]-\gamma_{i+1}(1-\gamma_{i})^{2}\sum_{a}\pi_{S}(a)(f^{*}(a))^{2}$ by Lemma 8.10
$=\gamma_{i+1}\sum_{a}\pi_{S}(a)\mathbb{E}_{\pi_{S\cup a}}[(f^{*})^{2}]-\gamma_{i+1}(1-\gamma_{i})^{2}\mathbb{E}_{\pi_{S}}[(f^{*})^{2}]$
$=\gamma_{i+1}\mathbb{E}_{\pi_{S}}[(f^{*})^{2}]-\gamma_{i+1}(1-\gamma_{i})^{2}\mathbb{E}_{\pi_{S}}[(f^{*})^{2}]$ by Eq. 89
$=\gamma_{i+1}\mathrm{Var}_{\pi_{S}}(f^{*})-\gamma_{i+1}(1-\gamma_{i})^{2}\mathrm{Var}_{\pi_{S}}(f^{*}).$
Now, since $\mathrm{Var}_{\pi_{S}}(f^{*})>0$ we have
$\gamma_{i}\geq\gamma_{i+1}(1-(1-\gamma_{i})^{2})=\gamma_{i+1}\gamma_{i}(2-\gamma_{i}).$
Since $\gamma_{i}>0$ (we assumed that $Q_{S}$ is irreducible) we have
$1\geq\gamma_{i+1}(2-\gamma_{i}),$
which is equivalent to:
$\gamma_{i}\geq 2-\frac{1}{\gamma_{i+1}},$
which completes the proof. ∎
8.6 Proofs of Technical Lemmas
It remains to prove Lemma 8.9 and Lemma 8.10.
Proof of Lemma 8.9.
Recall $i=|S|$, and $\mu$ is the uniform distribution over subsets $S\subset E$ which are bases of the matroid. Let us begin with two basic facts. Observe that
$\pi_{S}(a)=\pi_{S,1}(a)=\frac{\mu(S\cup a)}{(n-i)\mu(S)}=\frac{\pi_{i+1}(S\cup a)\binom{n}{i+1}}{(n-i)\pi_{i}(S)\binom{n}{i}}=\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}.$ (91)
Applying Eq. 41 we have the following:
$\sum_{a\in E\setminus T}\pi_{k+1}(T\cup a)=(k+1)\pi_{k}(T).$ (92)
We can now proceed with the proof of Eq. 89:
$\mathbb{E}_{\pi_{S}}(f)$ $=\sum_{b\in E\setminus S}\pi_{S}(b)f(b)$
$=\sum_{b\in E\setminus S}\frac{\pi_{i+1}(S\cup b)}{(i+1)\pi_{i}(S)}f(b)$ by Eq. 91
$=\sum_{b\in E\setminus S}\sum_{a\in E\setminus(S\cup b)}\frac{\pi_{i+2}(S\cup b\cup a)}{(i+1)(i+2)\pi_{i}(S)}f(b)$ by Eq. 92 with $T=S\cup b$
$=\sum_{a\in E\setminus S}\sum_{b\in E\setminus(S\cup a)}\frac{\pi_{i+2}(S\cup b\cup a)}{(i+2)\pi_{i+1}(S\cup a)}\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}f(b)$
$=\sum_{a\in E\setminus S}\pi_{S}(a)\sum_{b\in E\setminus(S\cup a)}\frac{\pi_{i+2}(S\cup b\cup a)}{(i+2)\pi_{i+1}(S\cup a)}f(b)$ by Eq. 92
$=\sum_{a\in E\setminus S}\pi_{S}(a)\sum_{b\in E\setminus(S\cup a)}\pi_{S\cup a}(b)f(b)$ by Eq. 91
$=\sum_{a\in E\setminus S}\pi_{S}(a)\mathbb{E}_{\pi_{S\cup a}}(f).$
We can establish Eq. 88 in a similar manner:
$\mathcal{E}_{Q_{S}}$ $(f)=\sum_{b,c\in E\setminus S}\pi_{S}(b)Q_{S}(b,c)(f(b)-f(c))^{2}$
$=\sum_{b,c\in E\setminus S}\pi_{S}(b)(i+2)\mathrm{P}_{i+1}^{\wedge}(S\cup b,S\cup c)(f(b)-f(c))^{2}$ by Eq. 44
$=\sum_{b,c\in E\setminus S}\pi_{S}(b)\frac{\pi_{i+2}(S\cup b\cup c)}{(i+2)\pi_{i+1}(S\cup b)}(f(b)-f(c))^{2}$ by Eq. 43
$=\sum_{b,c\in E\setminus S}\frac{\pi_{i+1}(S\cup b)}{(i+1)\pi_{i}(S)}\frac{\pi_{i+2}(S\cup b\cup c)}{(i+2)\pi_{i+1}(S\cup b)}(f(b)-f(c))^{2}$ by Eq. 91
$=\sum_{b,c\in E\setminus S}\frac{\pi_{i+2}(S\cup b\cup c)}{(i+1)(i+2)\pi_{i}(S)}(f(b)-f(c))^{2}$
$=\sum_{b,c\in E\setminus S}\sum_{a\in E\setminus(S\cup b\cup c)}\frac{\pi_{i+3}(S\cup a\cup b\cup c)}{(i+3)(i+2)(i+1)\pi_{i}(S)}(f(b)-f(c))^{2}$ by Eq. 92
$=\sum_{a\in E\setminus S}\sum_{b,c\in E\setminus(S\cup a)}\frac{\pi_{i+3}(S\cup a\cup b\cup c)}{(i+3)(i+2)\pi_{i+1}(S\cup a)}\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}(f(b)-f(c))^{2}$
$=\sum_{a\in E\setminus S}\pi_{S}(a)\sum_{b,c\in E\setminus(S\cup a)}\frac{\pi_{i+3}(S\cup a\cup b\cup c)}{(i+3)(i+2)\pi_{i+1}(S\cup a)}(f(b)-f(c))^{2}$ by Eq. 91
$=\sum_{a\in E\setminus S}\pi_{S}(a)\sum_{b,c\in E\setminus(S\cup a)}\pi_{S\cup a}(b)\frac{\pi_{i+3}(S\cup a\cup b\cup c)}{(i+3)\pi_{i+2}(S\cup a\cup b)}(f(b)-f(c))^{2}$ by Eq. 91
$=\sum_{a\in E\setminus S}\pi_{S}(a)\sum_{b,c\in E\setminus(S\cup a)}\pi_{S\cup a}(b)(i+3)\mathrm{P}_{i+2}^{\wedge}(S\cup a\cup b,S\cup a\cup c)(f(b)-f(c))^{2}$ by Eq. 43
$=\sum_{a\in E\setminus S}\pi_{S}(a)\sum_{b,c\in E\setminus(S\cup a)}\pi_{S\cup a}(b)Q_{S\cup a}(b,c)(f(b)-f(c))^{2}$ by Eq. 44
$=\sum_{a\in E\setminus S}\pi_{S}(a)\mathcal{E}_{Q_{S\cup a}}(f).\qed$
We now prove Lemma 8.10, which will complete the proof of the Trickle-Down Theorem (Theorem 8.1).
Proof of Lemma 8.10.
It will be more convenient to work with the following minimization problem equivalent to (90):
$\gamma_{i}=\min_{f;\mathrm{Var}_{\pi_{S}}(f^{*})=1}\mathcal{E}_{Q_{S}}(f).$ (93)
By Lagrange multipliers a critical point of (93) satisfies for every $a\in E\setminus S$
$\sum_{b\in E\setminus S}\frac{\pi_{i+2}(S\cup a\cup b)}{(i+2)(i+1)\pi_{i}(S)}(f(a)-f(b))=\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}\lambda f(a),$ (94)
for some multiplier $\lambda$. Plugging (94) and $\mathrm{Var}_{\pi_{S}}(f^{*})=1$ into $\mathcal{E}_{Q_{S}}(f)$ we obtain $\lambda=\gamma_{i}$.
Note that we can rewrite (94) as follows.
$\sum_{b\in E\setminus S}\frac{\pi_{i+2}(S\cup a\cup b)}{(i+2)(i+1)\pi_{i}(S)}f(b)$ $=f(a)\sum_{b\in E\setminus S}\frac{\pi_{i+2}(S\cup a\cup b)}{(i+2)(i+1)\pi_{i}(S)}-\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}\lambda f(a)$
$=f(a)\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}-\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}\lambda f(a)$ by Eq. 41
$=(1-\lambda)f(a)\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}.$
Dividing both sides by $\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}$ we obtain
$\sum_{b\in E\setminus S}\frac{\pi_{i+2}(S\cup a\cup b)}{(i+2)\pi_{i+1}(S\cup a)}f(b)=(1-\lambda)f(a).$
Since
$\mathbb{E}_{\pi_{S\cup a}}(f^{*})=\sum_{b\in E\setminus S}\frac{\pi_{i+2}(S\cup a\cup b)}{(i+2)\pi_{i+1}(S\cup a)}f(b),$
this completes the proof of the lemma. ∎
9 Methods for Establishing Spectral Independence
In this section, we present general techniques for establishing spectral independence, using the hard-core model over independent sets as an example and thereby establish Theorem 1.14.
As we saw previously, spectral independence immediately implies $O(n\log{n})$ mixing time of the Glauber dynamics for constant-degree graphs. We focus on three approaches. In Section 9.2, we show that correlation decay approaches as used in Weitz’s algorithm [Wei06] imply spectral independence. In Section 9.3, we show that stability of the partition function, so-called zero-freeness, also implies spectral independence; such conditions were used in the approximate counting algorithm introduced by Barvinok [Bar16a]. Finally, in Section 9.4 we discuss the coupling independence method introduced in [CZ23] which gives a simple and intuitive way to establish spectral independence. In Section 9.5, we show that an optimal relaxation time for a local chain implies spectral independence, implying the necessity of spectral independence for optimal mixing of the Glauber dynamics.
9.1 Preliminaries
9.1.1 Hard-core model
For a graph $G$, recall $\Omega(G)$ is the collection of all independent sets of $G$. For an independent set $I\in\Omega(G)$, we say a vertex $v\in V$ is occupied if $v\in I$, and unoccupied otherwise; we often denote $v\in I$ as $v$ and $v\notin I$ as $\overline{v}$. Consequently, for the hard-core model we can simplify the notation in the following manner. For a pair of vertices $u,v\in V$, the influence of $u$ on $v$ is denoted as follows:
$\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)$ $=\mu_{G,\lambda}(v\in I\mid u\in I)-\mu_{G,\lambda}(v\in I\mid u\notin I)$
$=\mu_{G,\lambda}(v\mid u)-\mu_{G,\lambda}(v\mid\bar{u}).$
We define the occupancy ratio $R_{G,\lambda}(v)$ at $v$ as follows:
$R_{G,\lambda}(v)=\frac{\mu_{G,\lambda}(v)}{\mu_{G,\lambda}(\bar{v})}.$ (95)
9.1.2 Tree-uniqueness threshold
Fix an integer $d\geq 2$ and a real $\lambda>0$. Consider the hard-core model on a complete $d$-ary tree of height $h$ denoted by $T_{h}$; this is the complete $\Delta$-regular tree where $d=\Delta-1$. For $v\in V$, define the occupancy ratio at $v$ as
$R_{G,\lambda}(v)=\frac{\lambda Z_{G}^{v}(\lambda)}{Z_{G}^{\bar{v}}(\lambda)},$ (96)
where “vv” represents the pinning $\tau(v)=1$ (i.e., $v$ is occupied), and $\bar{v}$” represents the pinning $\tau(v)=0$ ($v$ is unoccupied).
The tree recursion is a function $F=F_{d,\lambda}$ that can be used to compute the occupancy ratio at the root, defined as
$F:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0},\quad F(R)=\frac{\lambda}{(1+R)^{d}}.$ (97)
Denote by $R_{h}=R_{d,\lambda,h}$ the root occupancy ratio for $T_{h}$; e.g., $R_{0}=\lambda$, $R_{1}=\lambda/(1+\lambda)^{d}$. Then one can easily show that $R_{h}=F(R_{h-1})$. A natural and important question is whether the sequence $\{R_{h}\}$ converges when $h$ tends to infinity, which is closely related to the Gibbs measure on the infinite $d$-ary tree. The answer to this question is determined by whether the (unique) fixed point of $F$ is attractive or repulsive. Denote the unique positive fixed point of $F$ by $R^{*}$, i.e., $R^{*}(1+R^{*})^{d}=\lambda$. Define the critical fugacity by
$\lambda_{c}(\Delta)=\frac{(\Delta-1)^{\Delta-1}}{(\Delta-2)^{\Delta}},$ (98)
where $\Delta=d+1$ is the maximum degree of complete $d$-ary trees. It can be shown that if $\lambda\leq\lambda_{c}(\Delta)$ then the fixed point $R^{*}$ is attractive and $R_{h}\to R^{*}$ as $h\to\infty$, and if instead $\lambda>\lambda_{c}(\Delta)$ then the fixed point $R^{*}$ is repulsive and $R_{2h-1}\to R^{\prime}$, $R_{2h}\to R^{\prime\prime}$ as $h\to\infty$ for some $R^{\prime}<R^{*}<R^{\prime\prime}$.
Let $\Delta\geq 3$ be an integer. The critical fugacity $\lambda_{c}(\Delta)$ captures phase transitions for the hard-core model in multiple aspects.
  • When $\lambda\leq\lambda_{c}(\Delta)$ there exists a unique Gibbs measure on the infinite $d$-array tree; meanwhile, when $\lambda>\lambda_{c}(\Delta)$ there are multiple Gibbs measures. For this reason the critical value $\lambda_{c}(\Delta)$ is called the tree-uniqueness threshold.
  • When $\lambda<\lambda_{c}(\Delta)$, for complete $d$-ary trees we have $|R_{h}-R_{h-1}|=\exp(-\Theta(h))$, which can be viewed as the difference of root occupancy ratios on $T_{h+1}$ between fixing all leaves to be unoccupied (corresponding to $R_{h}$ on $T_{h}$) and fixing all leaves to be occupied (corresponding to $R_{h-1}$ on $T_{h-1}$). This describes a spatial mixing/correlation decay property with exponential decay rate, which fails when $\lambda>\lambda_{c}(\Delta)$.
    More generally, when $\lambda<\lambda_{c}(\Delta)$, for any graph $G$ of maximum degree at most $\Delta$, for any vertex $v\in V$ and any two pinnings $\sigma,\tau$ on a subset of vertices $\Lambda\subseteq V\setminus v$, it holds $|R^{\sigma}(v)-R^{\tau}(v)|=\exp(-\Omega(\ell))$ where $\ell$ is the distance from $v$ to a closest vertex $u\in\Lambda$ such that $\sigma(u)\neq\tau(u)$. This is known as the strong spatial mixing property with exponential decay rate; see [Wei06].
  • There exists an open set $\Gamma$ of complex numbers containing the interval $0,\lambda_{c}(\Delta))$ such that, for all graph $G$ of maximum degree at most $\Delta$, one has $Z_{G}(\lambda)\neq 0$ whenever $\lambda\in\Gamma$. Meanwhile, the (complex) zeros of $Z_{G}(\lambda)$ can be arbitrarily close to $\lambda_{c}(\Delta)$ for a graph $G$ of maximum degree at most $\Delta$. See [PR19].
  • When $\lambda<\lambda_{c}(\Delta)$, there exists a fully polynomial-time approximation scheme (FPTAS) for the partition function $Z_{G}(\lambda)$ for all graphs $G$ of maximum degree at most $\Delta$ [Wei06, Bar15, PR17], and the Glauber dynamics for sampling from $\mu_{G,\lambda}$ converges in $O(n\log n)$ steps (see Theorem 9.1 below). Meanwhile, when $\lambda>\lambda_{c}(\Delta)$ there is no $\mathsf{FPTAS}/\mathsf{FPRAS}$ for estimating the partition function for a graph $G$ of maximum degree at most $\Delta$ assuming $\mathsf{RP}\neq\mathsf{NP}$ [Sly10, SS14, GŠV16], see Theorem 1.12; this hardness reduction utilizes as a gadget that the Gibbs distribution on random $\Delta$-regular bipartite graphs is bimodal when $\lambda>\lambda_{c}(\Delta)$ and hence the Glauber dynamics has exponential mixing time.
9.1.3 Spectral independence
Consider the hard-core model on a graph $G=(V,E)$ with fugacity $\lambda>0$. Recall, for two distinct vertices $u,v\in V$, the (pairwise) influence of $u$ on $v$ is defined by:
$\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)=\mu_{G,\lambda}(v\mid u)-\mu_{G,\lambda}(v\mid\bar{u}).$
And for the diagonal entries we set $\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}u)=1$. For a pinning $\tau$, we also define the influence matrix $\Psi_{G,\lambda}^{\tau}$ for the conditional Gibbs distribution $\mu_{G,\lambda}^{\tau}$, where we let $\Psi_{G,\lambda}^{\tau}(u\text{\scriptsize{$\,\rightarrow\,$}}v)=0$ if $\tau$ forces $u$ to be unoccupied (note that in this case $\Psi_{G,\lambda}^{\tau}(v\text{\scriptsize{$\,\rightarrow\,$}}u)=0$ by definition). Finally, recall the Gibbs distribution $\mu_{G,\lambda}$ is said to be $\eta$-spectrally independent if for any pinning $\tau$, the maximum eigenvalue of the influence matrix $\Psi_{G,\lambda}^{\tau}$ is at most $1+\eta$.
The main aim in this section is to establish the following spectral independence result for the hard-core model in the tree-uniqueness regime.
Theorem 9.1.
For all integer $\Delta\geq 3$ and real $\delta\in(0,1)$, there exists $\eta=\eta(\delta)=O(1/\delta)$, such that for any graph $G=(V,E)$ of maximum degree at most $\Delta$, any vertex $u\in V$, and any $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, it holds
$\sum_{v\in V}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|\leq 1+\eta.$ (99)
As a consequence, for any graph $G$ of maximum degree at most $\Delta$ the Gibbs distribution $\mu_{G,\lambda}$ is $\eta$-spectrally independent, and the mixing time of the Glauber dynamics is $O(n\log n)$.
Note, Theorem 1.14 is a corollary of Theorem 9.1. In this monograph, we prove Theorem 9.1 with a constant $1+\eta\leq 16\sqrt{2}/\delta$. A better bound was given by [CCYZ25, Lemma 3.18] which shows that spectral independence holds with $1+\eta\leq 4e(1+1/(\Delta-2))/\delta$. The constant here matters as it determines the mixing time upper bound of the Glauber dynamics for the critical hard-core model, i.e., when $\lambda=\lambda_{c}(\Delta)$; see [CCYZ25].
9.1.4 Relating influences and occupancy ratios
Here we give a lemma relating the influences of a vertex $u$ and the occupancy ratio at $u$. It is helpful to consider a more general setting where every vertex $v$ has a distinct fugacity $\lambda_{v}$. Let $\boldsymbol{\lambda}=(\lambda_{v})_{v\in V}$ be a vector of fugacities, and the hard-core distribution is then defined by, for $\sigma\in\Omega(G)$,
$\mu_{G,\boldsymbol{\lambda}}(\sigma)=\frac{\prod_{v\in\sigma}\lambda_{v}}{Z_{G}(\boldsymbol{\lambda})},$ (100)
where the multivariate partition function (independence polynomial) is defined as
$Z_{G}(\boldsymbol{\lambda})=\sum_{\sigma\in\Omega(G)}\prod_{v\in\sigma}\lambda_{v}.$ (101)
Viewing the influences and the occupancy ratios as rational functions of $\boldsymbol{\lambda}$, we have the following relationship.
Claim 9.2.
For two distinct vertices $u,v\in V$, we have
$\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v;\boldsymbol{\lambda})=\frac{\partial\log R_{G}(u;\boldsymbol{\lambda})}{\partial\log\lambda_{v}}=\frac{\lambda_{v}}{R_{G}(u;\boldsymbol{\lambda})}\frac{\partial R_{G}(u;\boldsymbol{\lambda})}{\partial\lambda_{v}}.$
Proof.
Similarly as Eq. 96, we have
$R_{G}(u;\boldsymbol{\lambda})=\frac{\lambda_{u}Z_{G}^{u}(\boldsymbol{\lambda})}{Z_{G}^{\bar{u}}(\boldsymbol{\lambda})},$ (102)
and hence
$\frac{\partial\log R_{G}(u;\boldsymbol{\lambda})}{\partial\log\lambda_{v}}=\frac{\partial}{\partial\log\lambda_{v}}\log\left(Z_{G}^{u}(\boldsymbol{\lambda})\right)-\frac{\partial}{\partial\log\lambda_{v}}\log\left(Z_{G}^{\bar{u}}(\boldsymbol{\lambda})\right).$
We compute that
$\frac{\partial}{\partial\log\lambda_{v}}\log\left(Z_{G}^{u}(\boldsymbol{\lambda})\right)$ $=\frac{\lambda_{v}}{Z_{G}^{u}(\boldsymbol{\lambda})}\frac{\partial}{\partial\lambda_{v}}Z_{G}^{u}(\boldsymbol{\lambda})$
$=\frac{\lambda_{v}}{Z_{G}^{u}(\boldsymbol{\lambda})}\sum_{I\in\mathcal{I}(G):\,u\in I}\frac{\partial}{\partial\lambda_{v}}\prod_{w\in I\setminus\{u\}}\lambda_{w}$
$=\frac{\lambda_{v}}{Z_{G}^{u}(\boldsymbol{\lambda})}\sum_{I\in\mathcal{I}(G):\,u,v\in I}\prod_{w\in I\setminus\{u,v\}}\lambda_{w}$
$=\frac{\lambda_{v}Z_{G}^{uv}(\boldsymbol{\lambda})}{Z_{G}^{u}(\boldsymbol{\lambda})}=\mu_{G,\boldsymbol{\lambda}}(v\mid u).$
Similarly, we have
$\frac{\partial}{\partial\log\lambda_{v}}\log\left(Z_{G}^{\bar{u}}(\boldsymbol{\lambda})\right)=\mu_{G,\boldsymbol{\lambda}}(v\mid\bar{u}).$
Therefore, we conclude that
$\frac{\partial\log R_{G}(u;\boldsymbol{\lambda})}{\partial\log\lambda_{v}}=\mu_{G,\boldsymbol{\lambda}}(v\mid u)-\mu_{G,\boldsymbol{\lambda}}(v\mid\bar{u})=\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v;\boldsymbol{\lambda}),$
as claimed. ∎
Since the occupancy ratios were intensively studied in previous works for establishing properties like correlation decay or zero-freeness, by Claim 9.2 we can transform these properties or their proof approaches into results for influences and thus establish spectral independence.
9.2 Spectral Independence via Correlation Decay
In this section, we show Theorem 9.1 establishing $\eta$-spectral independence with $\eta=O(1/\delta)$ when $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$ using an approach based on the strong spatial mixing (correlation decay) property, which appeared in [ALO20, CLV20] and was based on techniques in [Wei06, LLY13].
9.2.1 Proof approach
  • We need to show that for any graph $G=(V,E)$ of maximum degree at most $\Delta$ and any vertex $u\in V$ it holds
    $\sum_{v\in V}\left|\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|=O(1/\delta).$
  • Fix a graph $G=(V,E)$ of maximum degree at most $\Delta$ and a vertex $u\in V$. We construct a tree rooted at $u$ called the self-avoiding walk tree $T=T_{\textsc{saw}}(G,u)$, which enumerates all self-avoiding walks (in $G$) starting from $u$. The tree $T$ is, in general, exponentially large in $n=|V|$, and each vertex of $G$ can appear multiple times in $T$. The maximum degree of $T$ is at most $\Delta$.
    We will define a hard-core model on $T$, such that the following properties hold:
    • The occupancy ratio at $u$ is preserved:
      $R_{G}(u)=R_{T}(u).$
    • The influence from $u$ to another vertex $v$ is preserved:
      $\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v)=\sum_{w\in\mathcal{C}_{T}(v)}\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w),$
      where $\mathcal{C}_{T}(v)$ denotes the set of all copies of $v$ in $T$.
  • Then, it suffices to show that for any tree $T=(V,E)$ of maximum degree at most $\Delta$ and any vertex $u\in V$ it holds
    $\sum_{v\in V}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|=O(1/\delta).$
    This will be proved via the so-called potential function method [RST+13, LLY13].
9.2.2 Self-avoiding walk tree
We now define the self-avoiding walk tree more formally. Fix an arbitrary total order “<<” on the vertices of $G$.
Figure 1: An example of the self-avoiding walk tree and the hard-core model on it. Solid red vertices are those fixed to be occupied in the pinning $\xi$, and hollow blue vertices are those fixed to be unoccupied.
Definition 9.3 (Self-avoiding walk tree).
Let $G=(V,E)$ be a connected graph and $u\in V$ be a vertex. The self-avoiding walk (SAW) tree $T=T_{\textsc{saw}}(G,u)$ of $G$ rooted at $u$ is a tree consisting of all self-avoiding walks starting from $u$, defined as follows.
  • The root of $T$ is $u$;
  • Every path in $T$ from the root $u$ to a leaf corresponds to a “maximal” self-avoiding walk in $G$. More precisely, if $u=v_{0}$-$v_{1}$-$\cdots$-$_{\ell-1}-v=v$$_{\ell}=v$ is a path from $u$ to a leaf $v$, then it corresponds to a walk in $G$ (i.e., $\{v_{i-1},v_{i}\}\in E$) such that:
    • either $u=v_{0},v_{1},\dots,v_{\ell-1},v_{\ell}=v$ are all distinct vertices (so they form a self-avoiding walk), and $\deg_{G}(v)=1$ (so the self-avoiding walk is maximal);
    • or $u=v_{0},v_{1},\dots,v_{\ell-1}$ are all distinct vertices (so they form a self-avoiding walk), and $v=v_{\ell}=v_{i}$ for some $i\leq\ell-2$ (so the last vertex $v$ makes a cycle, and the self-avoiding walk is “maximal” in some sense).
Remark 9.4.
Observe the following basic properties of $T=T_{\textsc{saw}}(G,u)$:
  1. (1)
    The maximum degree of $T$ is the same as that of $G$.
  2. (2)
    Leaves in $T$ correspond to either “pendant vertices” (those of degree $1$) in $G$ or to the completion of a cycle.
  3. (3)
    Each vertex of $G$ possibly appears multiple times in $T$. We denote the set of all copies in $T$ of a vertex $v\in V$ by $\mathcal{C}_{T}(v)$.
  4. (4)
    If $G$ itself is a tree, then $T=G$. However, in general $T$ can be exponentially larger than $G$.
Now for the hard-core model defined on $G$ with the fugacity vector $\boldsymbol{\lambda}=(\lambda_{v})_{v\in V}$, we define an associated hard-core model on the self-avoiding walk tree with a specific pinning on some leaves.
Definition 9.5 (Hard-core model on $T_{\textsc{saw}}(G,u)$).
Let $G=(V,E)$ be a connected graph and $u\in V$ be a vertex. Let $T=T_{\textsc{saw}}(G,u)$ be the SAW tree of $G$ rooted at $u$. Define the hard-core model on $T$ with a pinning $\xi$ as follows.
  • For each vertex $v\in V$, every copy $w\in\mathcal{C}_{T}(v)$ of $v$ has the same fugacity $\lambda_{w}=\lambda_{v}$.
  • We define a pinning $\xi$ on a subset of leaves in the following way. Let $u=v_{0}$-$v_{1}$-$\cdots$-$_{\ell-1}-v=v$$_{\ell}=v$ be a path in $T$ from $u$ to a leaf $v$.
    • If $\deg_{G}(v)=1$, then $v$ is not pinned;
    • Otherwise $v=v_{\ell}=v_{i}$ for some $i\leq\ell-2$, and we fix $v_{i}$ to be occupied if $v_{i+1}<v_{\ell-1}$, and unoccupied if instead $v_{i+1}>v_{\ell-1}$.
Remark 9.6.
[…]
For each leaf $v$ pinned to be occupied we can remove $v$ and its parent, and for each leaf $v$ pinned to be unoccupied we can remove $v$. In this way we obtain an equivalent hard-core model on a tree $T^{\prime}$ without any pinned vertices, and the root vertex $u$ only appears once in $T^{\prime}$ since any additional occurences of $u$ in $T$ must have been as leaf vertices since they corresponded to closing a cycle.
Below we still use $\boldsymbol{\lambda}$ to denote the fugacity vector for the hard-core model on $T$. The following lemma relates the hard-core models on $G$ and on the corresponding SAW tree $T$.
Lemma 9.7 ([CLV20]).
Let $T=T_{\textsc{saw}}(G,u)$ be the SAW tree of $G$ rooted at $u$. Consider the hard-core model on $T$ with the pinning $\xi$ as defined in Definition 9.5.
  1. (1)
    $Z_{G}(\boldsymbol{\lambda})$ divides $Z_{T}^{\xi}(\boldsymbol{\lambda})$. Moreover, there exists a polynomial $P(\boldsymbol{\lambda})$ independent of $\lambda_{u}$, such that
    $Z_{T}^{\xi}(\boldsymbol{\lambda})=Z_{G}(\boldsymbol{\lambda})P(\boldsymbol{\lambda}).$
  2. (2)
    [Wei06] The occupancy ratio at $u$ is preserved:
    $R_{G}(u;\boldsymbol{\lambda})=R_{T}^{\xi}(u;\boldsymbol{\lambda}).$
  3. (3)
    The influence of $u$ on another vertex $v$ is preserved:
    $\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v;\boldsymbol{\lambda})=\sum_{w\in\mathcal{C}_{T}(v)}\Psi_{T}^{\xi}(u\text{\scriptsize{$\,\rightarrow\,$}}w;\boldsymbol{\lambda}).$
Proof of “(1)”.
We will prove (1) by induction on the lexicographic ordering of the following pair of numbers: number of vertices of degree more than one, number of vertices of degree one.
First suppose $u$ has degree one. Let $w$ be the neighbor of $u$. Let $G^{\prime}$ be $G$ with vertex $u$ removed. Let $T^{\prime}=T_{\textsc{saw}}(G^{\prime},w)$. By the induction hypothesis there exists a polynomial $P^{\prime}(\boldsymbol{\lambda})$ independent of $\lambda_{w}$ such that
$Z_{T^{\prime}}^{\xi}(\boldsymbol{\lambda})=Z_{G^{\prime}}(\boldsymbol{\lambda})P^{\prime}(\boldsymbol{\lambda}).$ (103)
Let $\boldsymbol{\lambda}^{\prime}$ be $\boldsymbol{\lambda}$ with $\lambda_{w}$ replaced by $\lambda_{w}/(1+\lambda_{u})$. Note that
$Z_{T}^{\xi}(\boldsymbol{\lambda})=(1+\lambda_{u})Z_{T^{\prime}}^{\xi}(\boldsymbol{\lambda}^{\prime})\quad\mbox{and}\quad Z_{G}(\boldsymbol{\lambda})=(1+\lambda_{u})Z_{G^{\prime}}(\boldsymbol{\lambda}^{\prime}).$ (104)
Plugging (104) into (103) we obtain (1) for $G$ and $T$ (since $P^{\prime}(\boldsymbol{\lambda})$ is independent of $\lambda_{w}$ we have $P^{\prime}(\boldsymbol{\lambda})=P^{\prime}(\boldsymbol{\lambda}^{\prime})$).
Now suppose $u$ has degree $d>1$. Let $w_{1},\dots,w_{d}$ be the neighbors of $v$. Let $G^{\prime}$ be the graph obtained by removing $u$ form $G$ and by adding $u_{1},\dots,u_{d}$ where $u_{i}$ is connected to $w_{i}$ for $i\in[d]$. For $i\in\{0,\dots,d+1\}$ let $\xi^{\prime}_{i}$ be a partial assignment of spins to $u_{1},\dots,u_{d}$ where $u_{j}$ for $j<i$ are assigned unoccupied and $u_{j}$ for $j>i$ are assigned occupied. Let $G_{i}$ be the graph that corresponds to the assignment $\xi^{\prime}_{i}$ in $G^{\prime}$ (we obtain $G_{i}$ from $G^{\prime}$ by removing $u_{j}$ for $j<i$ and removing $u_{j}$ and $N(u_{j})=\{w_{j}\}$ for $j>i$). Let $T_{i}=T_{\textsc{saw}}(G^{\prime},u_{i})$. Note that $T_{i}$ is a subtree of $T$ (it includes the root, one child of the root, and the entire subtree rooted at the child) and the pinning $\xi_{i}$ on $T_{i}$ defined by Definition 9.5 is compatible with $\xi^{\prime}_{i}$. By the induction hypothesis there exist polynomials $P_{i}$ independent of $\lambda_{u_{i}}$ such that
$Z_{T_{i}}^{\xi_{i}}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi_{i}^{\prime}}(\boldsymbol{\lambda})P_{i}(\boldsymbol{\lambda})=Z_{G_{i}}(\boldsymbol{\lambda})P_{i}(\boldsymbol{\lambda}).$ (105)
Recall that $Z_{G}^{u}(\boldsymbol{\lambda})$ is $(\partial/\partial\lambda_{u})Z_{G}(\boldsymbol{\lambda})$, that is, the part of $Z_{G}(\boldsymbol{\lambda})$ with $u$ occupied and $Z_{G}^{\bar{u}}(\boldsymbol{\lambda})$ is $Z_{G}(\boldsymbol{\lambda})|_{\lambda_{u}=0}$, that is, the part of $Z_{G}(\boldsymbol{\lambda})$ with $u$ unoccupied. Note that we have
$Z_{G}^{u}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi^{\prime}_{0}}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi^{\prime}_{1},u_{1}}(\boldsymbol{\lambda})\quad\mbox{and}\quad Z_{G}^{\bar{u}}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi^{\prime}_{d+1}}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi^{\prime}_{d},\bar{u_{d}}}(\boldsymbol{\lambda}).$ (106)
We also have
$Z_{T}^{\xi,u}=\prod_{i=1}^{d}Z_{T_{i}}^{\xi_{i},u_{i}}(\boldsymbol{\lambda}),\quad Z_{T}^{\xi,\bar{u}}=\prod_{i=1}^{d}Z_{T_{i}}^{\xi_{i},\bar{u_{i}}}(\boldsymbol{\lambda}).$ (107)
Since $P_{i}$ does not depend on $\lambda_{u_{i}}$ from (105) we have
$Z_{T_{i}}^{\xi_{i},u_{i}}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi_{i}^{\prime},u_{i}}(\boldsymbol{\lambda})P_{i}(\boldsymbol{\lambda})\quad\mbox{and}\quad Z_{T_{i}}^{\xi_{i},\bar{u_{i}}}(\boldsymbol{\lambda})=Z_{G^{\prime}}^{\xi_{i}^{\prime},\bar{u_{i}}}(\boldsymbol{\lambda})P_{i}(\boldsymbol{\lambda}).$ (108)
Plugging (108) into (107) we obtain
$Z_{T}^{\xi,u}=\prod_{i=1}^{d}Z_{G^{\prime}}^{\xi_{i},u_{i}}(\boldsymbol{\lambda})P_{i}(\boldsymbol{\lambda})=Z_{G}^{u}\prod_{i=2}^{d}Z_{G^{\prime}}^{\xi_{i},u_{i}}(\boldsymbol{\lambda})\prod_{i=1}^{d}P_{i}(\boldsymbol{\lambda}).$
Similarly we have
$Z_{T}^{\xi,\bar{u}}=\prod_{i=1}^{d}Z_{G^{\prime}}^{\xi_{i},\bar{u_{i}}}(\boldsymbol{\lambda})P_{i}(\boldsymbol{\lambda})=Z_{G}^{\bar{u}}\prod_{i=1}^{d-1}Z_{G^{\prime}}^{\xi_{i},\bar{u_{i}}}(\boldsymbol{\lambda})\prod_{i=1}^{d}P_{i}(\boldsymbol{\lambda})=Z_{G}^{\bar{u}}\prod_{i=2}^{d}Z_{G^{\prime}}^{\xi_{i},u_{i}}(\boldsymbol{\lambda})\prod_{i=1}^{d}P_{i}(\boldsymbol{\lambda}).$
Letting $P(\boldsymbol{\lambda})=\prod_{i=2}^{d}Z_{G^{\prime}}^{\xi_{i},u_{i}}(\boldsymbol{\lambda})\prod_{i=1}^{d}P_{i}(\boldsymbol{\lambda})$ we have
$Z_{T}^{\xi}=Z_{T}^{\xi,\bar{u}}+\lambda_{u}Z_{T}^{\xi,u}=\left(Z_{G}^{\bar{u}}+\lambda_{u}Z_{G}^{u}\right)P(\boldsymbol{\lambda})=Z_{G}P(\boldsymbol{\lambda}).$
Proof of “(1) $\Rightarrow$ (2)”.
Note that
$R_{G}(u;\boldsymbol{\lambda})=\lambda_{u}\frac{Z_{G}^{u}(\boldsymbol{\lambda})}{Z_{G}^{\bar{u}}(\boldsymbol{\lambda})}=\lambda_{u}\frac{Z_{G}^{u}(\boldsymbol{\lambda})P(\boldsymbol{\lambda})}{Z_{G}^{\bar{u}}(\boldsymbol{\lambda})P(\boldsymbol{\lambda})}=\lambda_{u}\frac{Z_{T}^{\xi,u}(\boldsymbol{\lambda})}{Z_{T}^{\xi,\bar{u}}(\boldsymbol{\lambda})}=R_{T}^{\xi}(u;\boldsymbol{\lambda}).$
Proof of “(2) $\Rightarrow$ (3)”.
We deduce from Claim 9.2 and the chain rule that
$\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v;\boldsymbol{\lambda})$ $=\frac{\partial\log R_{G}(u;\boldsymbol{\lambda})}{\partial\log\lambda_{v}}$ (Claim 9.2)
$=\frac{\partial\log R_{T}^{\xi}(u;\boldsymbol{\lambda})}{\partial\log\lambda_{v}}$ (Part (2))
$=\sum_{w\in\mathcal{C}_{T}(v)}\frac{\partial\log R_{T}^{\xi}(u;\boldsymbol{\lambda})}{\partial\log\lambda_{w}}\frac{\partial\log\lambda_{w}}{\partial\log\lambda_{v}}$ (Chain rule)
$=\sum_{w\in\mathcal{C}_{T}(v)}\Psi_{T}^{\xi}(u\text{\scriptsize{$\,\rightarrow\,$}}w;\boldsymbol{\lambda}),$ (Claim 9.2)
which shows Part (3). ∎
Note, for a vertex $w$ in $G$, if $w$ is fixed by $\xi$, then $\Psi_{T}^{\xi}(u\text{\scriptsize{$\,\rightarrow\,$}}w;\boldsymbol{\lambda})=0$ by definition.
Corollary 9.8.
By the triangle inequality, we deduce that
$\sum_{v\in V(G)}\left|\Psi_{G}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|\leq\sum_{w\in V(T)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|.$
Hence, we reduce the problem to trees (albeit a possibly exponentially large tree).
9.2.3 Bounding influences on trees
In this subsection, we bound the absolute sum of influences of the root on bounded-degree trees. In particular, we show the following result of Chen, Liu, and Vigoda [CLV20].
Lemma 9.9.
Let $T=(V,E)$ be a tree of maximum degree at most $\Delta$ rooted at $u$. For an integer $k\in\mathbb{N}^{+}$ and a vertex $v\in V$, let $L_{v}(k)$ denote the set of all descendants at distance $k$ from $v$. Then for all $k\in\mathbb{N}^{+}$ we have
$\sum_{v\in L_{u}(k)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|\leq 2\sqrt{2}(1-\delta/8)^{k}.$
The following claim is helpful to us.
Claim 9.10.
Let $T=(V,E)$ be a tree and $u,v\in V$ be two distinct vertices. If $w$ is a vertex on the unique path from $u$ to $v$, then
$\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)=\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\,\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v).$
Proof.
Using the Markov property of the hard-core model (i.e., conditional on the value of $w$, the two vertices $u$ and $v$ are independent), we deduce that
$\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)={}$ $\mu_{T}(v\mid u)-\mu_{T}(v\mid\bar{u})$
$={}$ $\mu_{T}(w\mid u)\mu_{T}(v\mid w)+\mu_{T}(\bar{w}\mid u)\mu_{T}(v\mid\bar{w})$
$-\mu_{T}(w\mid\bar{u})\mu_{T}(v\mid w)-\mu_{T}(\bar{w}\mid\bar{u})\mu_{T}(v\mid\bar{w})$
$={}$ $\left(\mu_{T}(w\mid u)-\mu_{T}(w\mid\bar{u})\right)\mu_{T}(v\mid w)-\left(\mu_{T}(\bar{w}\mid\bar{u})-\mu_{T}(\bar{w}\mid u)\right)\mu_{T}(v\mid\bar{w})$
$={}$ $\left(\mu_{T}(w\mid u)-\mu_{T}(w\mid\bar{u})\right)\left(\mu_{T}(v\mid w)-\mu_{T}(v\mid\bar{w})\right)$
$={}$ $\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\,\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v),$
as claimed. ∎
For an integer $k\geq 1$ and a vertex $u$, recall $L_{u}(k)$ is the set of all descendants at distance $k$ from $u$. We can apply Claim 9.10 to build an inductive approach for bounding the sum of the influences from a vertex $u$ as follows:
$\sum_{v\in L_{u}(k)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ $=\sum_{w\in L_{u}(k-1)}\sum_{v\in L_{w}(1)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$
$=\sum_{w\in L_{u}(k-1)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|\sum_{v\in L_{w}(1)}\left|\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v)\right|.$ (109)
Hence, an easy approach to establish Lemma 9.9 is to show for all $w$,
$\sum_{v\in L_{w}(1)}\left|\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v)\right|\leq(1-\delta/8).$ (110)
Then Lemma 9.9 follows by induction. But such an approach does not work since Eq. 110 is not true for all $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, though it is easy to check that Eq. 110 holds when $\lambda\leq\frac{1-\delta}{\Delta-2}$ which is below the uniqueness threshold. Note, for the Ising model (see Section 12.1), the analog of Eq. 110 holds for all $\beta$ where $|\beta|<\beta_{c}(\Delta)$, and hence we do not need to consider a potential function for the Ising model.
To prove Lemma 9.9 we use the potential function method, which establishes a generalization of Eq. 110.
Proof of Lemma 9.9.
The multivariate tree recursion is a function $F=F_{d,\lambda}:\mathbb{R}_{\geq 0}^{d}\to\mathbb{R}_{\geq 0}$ given by
$R=F(R_{1},\dots,R_{d}):=\frac{\lambda}{\prod_{i=1}^{d}(1+R_{i})},$
which means that for a tree rooted at $w$ with $d$ children $v_{1},\dots,v_{d}$, the occupancy ratio $R_{T}(w)$ at the root is given by $R_{T}(w)=F(R_{T_{1}}(v_{1}),\dots,R_{T_{d}}(v_{d}))$ where $T_{i}$ is the subtree rooted at $v_{i}$ and $R_{T_{i}}(v_{i})$ is the root occupancy ratio of $T_{i}$. It would be helpful to consider the logarithm of occupancy ratios in the spirit of Claim 9.2. Writing $x=\log R$ and $x_{i}=\log R_{i}$, we define a multivariate function $H=H_{d,\lambda}:\mathbb{R}^{d}\to\mathbb{R}$ by
$x=H(x_{1},\dots,x_{d}):=\log\lambda-\sum_{i=1}^{d}\log(1+e^{x_{i}}).$
One can check that
$\frac{\partial H}{\partial x_{i}}=\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v_{i})$
which is similar to Claim 9.2, and therefore
$\left\lVert\nabla H\right\rVert_{1}=\sum_{i=1}^{d}\left|\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v_{i})\right|.$
Let $\varphi:\mathbb{R}\to\mathbb{R}$ be a suitable potential function that is monotone increasing, and we consider the tree recursion composed with $\varphi$. That is, let $y=\varphi(x)$ and $y_{i}=\varphi(x_{i})$, and then for $H^{\varphi}=\varphi\circ H\circ\varphi^{-1}$ we have
$y=H^{\varphi}(y_{1},\dots,y_{d}).$
Moreover, it is easy to check that
$\left\lVert\nabla H^{\varphi}\right\rVert_{1}=\sum_{i=1}^{d}\frac{\varphi^{\prime}(x)}{\varphi^{\prime}(x_{i})}\left|\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v_{i})\right|.$ (111)
Therefore, we have the following:
$\sum_{v\in L_{u}(k)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ $=\sum_{v\in L_{u}(k)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{v})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$
$=\sum_{w\in L_{u}(k-1)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{w})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|\sum_{v\in L_{w}(1)}\frac{\varphi^{\prime}(x_{w})}{\varphi^{\prime}(x_{v})}\left|\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ by Eq. 109 (112)
It remains to choose a suitable potential function. We use an ingenious potential function given in [LLY13]:
$\varphi(x)=\log\left(e^{x/2}+\sqrt{e^{x}+1}\right).$
Note that we have
$\varphi^{\prime}(x)=\sqrt{\frac{e^{x/2}}{e^{x/2}+1}}\in\left[\frac{1}{\sqrt{2}},1\right].$ (113)
We will prove the following contraction for this potential function.
Claim 9.11.
For any integer $d\leq\Delta-1$, and for any $x_{i}$ and $x=H(x_{1},\dots,x_{d})$, for any $\delta>0$, for $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$,
$\left\lVert\nabla H^{\varphi}\right\rVert_{1}\leq(1-\delta/8),$ (114)
and when $d=\Delta$ then we have $\left\lVert\nabla H^{\varphi}\right\rVert_{1}\leq 2(1-\delta/8)$.
Using Claim 9.11 we will prove by induction that for every integer $k\geq 1$:
$\sum_{v\in L_{u}(k)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{v})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|\leq 2(1-\delta/8)^{k}$ (115)
For the base case $k=1$ we have the following bound by Claim 9.11:
$\sum_{v\in L_{u}(1)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{v})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|=\left\lVert\nabla H^{\varphi}\right\rVert_{1}\leq 2(1-\delta/8).$ (116)
Now suppose Eq. 115 holds for k1k-1, and then we have:
$\sum_{v\in L_{u}(k)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{v})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$
$={}$ $\sum_{w\in L_{u}(k-1)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{w})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|\sum_{v\in L_{w}(1)}\frac{\varphi^{\prime}(x_{w})}{\varphi^{\prime}(x_{v})}\left|\Psi_{T}(w\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ by Claim 9.11
$\leq{}$ $\sum_{w\in L_{u}(k-1)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{w})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|\cdot(1-\delta/8)$ by Eq. 116
$\leq{}$ $2(1-\delta/8)^{k}$ by induction.
We can now complete the proof of the lemma as follows:
$\sum_{v\in L_{u}(k)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ $=\sum_{v\in L_{u}(k)}\frac{\varphi^{\prime}(x_{w})}{\varphi^{\prime}(x_{u})}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{w})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$
$\leq\frac{\max_{x}\varphi^{\prime}(x)}{\min_{x}\varphi^{\prime}(x)}\sum_{v\in L_{u}(k)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{w})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$
$\leq\sqrt{2}\sum_{v\in L_{u}(k)}\frac{\varphi^{\prime}(x_{u})}{\varphi^{\prime}(x_{w})}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ by Eq. 113
$\leq 2\sqrt{2}(1-\delta/8)^{k},$ by Eq. 115
as claimed. ∎
It remains to prove Claim 9.11 which we do so now.
Proof of Claim 9.11.
Note that the root $u$ can have $\Delta$ children while any other vertex has at most $\Delta-1$ children.
First, observe that:
$\left\lVert\nabla H^{\varphi}\right\rVert_{1}=\sum_{i=1}^{d}\sqrt{\frac{e^{x}}{e^{x}+1}}\sqrt{\frac{e^{x_{i}}}{e^{x_{i}}+1}}.$ (117)
Recalling $x=\log R$, $x_{i}=\log R_{i}$ and writing $p=R/(R+1)$, $p_{i}=R_{i}/(R_{i}+1)$ for the occupancy probability, and letting $q=\frac{1}{d}\sum_{i=1}^{d}p_{i}$, then we deduce by the Cauchy-Schwarz inequality that
$\left\lVert\nabla H^{\varphi}\right\rVert_{1}^{2}=\left(\sum_{i=1}^{d}\sqrt{pp_{i}}\right)^{2}\leq dp\sum_{i=1}^{d}p_{i}=d^{2}pq.$ (118)
The tree recursion, in terms of the occupancy probability, is given by
$\frac{p}{1-p}=\lambda\prod_{i=1}^{d}(1-p_{i})\leq\lambda\left(1-\frac{1}{d}\sum_{i=1}^{d}p_{i}\right)^{d}=\lambda(1-q)^{d},$ (119)
where the inequality is by the arithmetic-geometric mean inequality (AM-GM). Combining Eqs. 118 and 119 yields:
$\left\lVert\nabla H^{\varphi}\right\rVert_{1}^{2}\leq d^{2}pq\leq d^{2}q\cdot\frac{\lambda(1-q)^{d}}{\lambda(1-q)^{d}+1},$ (120)
where we used the fact that $p/(1-p)\leq x$ is equivalent to $p\leq x/(1+x)$.
To bound Eq. 120 we use the following property of the fixpoints of the tree recursions.
Claim 9.12.
Let $p^{*}$ be the fixed point of the (univariate) tree recursion such that $\lambda=p^{*}/(1-p^{*})^{d+1}$. For $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$ we have
$p^{*}\leq\frac{1-\delta/4}{d}.$
Now we prove that
$q\cdot\frac{\lambda(1-q)^{d}}{\lambda(1-q)^{d}+1}\leq\frac{p^{*}}{d}.$ (121)
We observe Eq. 121 is equivalent to
$\lambda(1-q)^{d}\leq\frac{p^{*}}{dq-p^{*}}\quad\Leftrightarrow\quad(dq-p^{*})(1-q)^{d}\leq(1-p^{*})^{d+1}.$
By the AM–GM inequality, we have
$(dq-p^{*})(1-q)^{d}\leq\left(\frac{(dq-p^{*})+d(1-q)}{d+1}\right)^{d+1}=\left(\frac{d-p^{*}}{d+1}\right)^{d+1}\leq(1-p^{*})^{d+1},$
where the last inequality follows from $p^{*}\leq 1/d$, which is implied by Claim 5.7. This establishes Eq. 121.
Finally, Eq. 114, and hence Claim 9.11, follows from plugging the bound from Eq. 121 and Claim 9.12 into Eq. 120.
$\left\lVert\nabla H^{\varphi}\right\rVert_{1}^{2}$ $\leq d^{2}q\cdot\frac{\lambda(1-q)^{d}}{\lambda(1-q)^{d}+1}.$ by Eq. 120
$\leq dp^{*}$ by Eq. 121
$\leq(1-\delta/4)$ by Claim 9.12.
Since $\left\lVert\nabla H^{\varphi}\right\rVert_{1}^{2}\leq(1-\delta/4)$ then we have that $\left\lVert\nabla H^{\varphi}\right\rVert_{1}\leq(1-\delta/8)$, which completes the proof of Lemma 9.9. ∎
Using Lemma 9.9 we can complete the proof of the main result (Theorem 9.1) of this section.
Proof of Theorem 9.1.
We can now conclude the main technical result Eq. 99 in Theorem 9.1.
$\sum_{v\in V(G)}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$ $\leq\sum_{w\in V(T)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|.$ by Corollary 9.8
$=\sum_{k}\sum_{w\in L_{u}(k)}\left|\Psi_{T}(u\text{\scriptsize{$\,\rightarrow\,$}}w)\right|$
$\leq 2\sqrt{2}\sum_{k}(1-\delta/8)^{k}$ by Lemma 9.9
$\leq\frac{16\sqrt{2}}{\delta},$
which proves Eq. 99 in Theorem 9.1 with $1+\eta=16\sqrt{2}/\delta$.
To see why Eq. 99 implies spectral independence, recall from Lemma 2.3 in Section 2.2 that for any graph $G$ of maximum degree at most $\Delta$, the maximum eigenvalue of $\Psi_{G,\lambda}$ is upper bounded by $\left\lVert\Psi_{G,\lambda}\right\rVert_{\infty}=\max_{u\in V}\sum_{v\in V}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|$, which is at most $1+\eta$ by Eq. 99. Meanwhile, for any pinning $\tau\in\{0,1\}^{\Lambda}$, the conditional Gibbs distribution corresponds to the hard-core model on a smaller graph (removing $\Lambda$ and neighbors of occupied vertices in $\Lambda$) which also has maximum degree at most $\Delta$. Since Eq. 99 applies to all graphs of maximum degree at most $\Delta$, the maximum eigenvalue of $\Psi_{G,\lambda}^{\tau}$ is also at most $1+\eta$. Hence, $\eta$-spectral independence follows and then optimal mixing of the Glauber dynamics follows from Theorem 1.8. ∎
Proof of Claim 9.12.
It is enough to show that for $p=\frac{1-\delta/4}{d}$ and $\lambda=(1-\delta)\frac{(\Delta-1)^{\Delta-1}}{(\Delta-2)^{\Delta}}$ we have
$p>\lambda(1-p)^{\Delta},$ (122)
since the fixpoint cannot occur in $p,\infty)$ (note that the LHS of (122) is increasing and the RHS of (122) is decreasing). Plugging in the values of $p$ and $\lambda$ in (122) and simplifying we obtain the following equivalent inequality.
$1-\delta/4>(1-\delta)\left(1+\frac{\delta/4}{\Delta-2}\right)^{\Delta}$ (123)
The RHS of (123) is maximized for $\Delta=3$ and hence it is enough to show
$1-\delta/4>(1-\delta)\left(1+\delta/4\right)^{3},$
which is equivalent to $\delta^{2}(\delta^{2}+11\delta+36)>0$. ∎
Remark 9.13.
Another choice of the potential function is from [RST+13]:
$\varphi(x)=\log\left(e^{x}+\frac{1}{\Delta}\right).$
9.3 Spectral Independence via Zero-Freeness
In this section, we give a proof of a weaker version Theorem 9.1 using the zero-freeness of the partition function; the version proved here is weaker in the sense that it does not have an explicit bound on $\eta$. The approach here is based on [AASV21] with appropriate modifications and generalizations; see also [CLV21b] for a more general setting.
9.3.1 Some preliminaries
For a complex number $\zeta\in\mathbb{C}$ and a real number $r>0$, let
$\mathcal{D}(\zeta,r)=\{z\in\mathbb{C}:|z-\zeta|<r\}$
be the open disk around $\zeta$ of radius $r$. Furthermore, for a subset $A\subseteq\mathbb{C}$ of complex numbers define
$\mathcal{D}(A,r)=\bigcup_{\zeta\in A}\mathcal{D}(\zeta,r).$
Let $\overline{\mathcal{D}}(\zeta,r)$ and $\overline{\mathcal{D}}(A,r)$ denote their closure.
Consider the multivariate independence polynomial defined by Eq. 101. The following stability (zero-freeness) result was established by Peters and Regts [PR19] and is the basis of our approach.
Theorem 9.14 ([PR19, Theorem 4.2]).
Let $\Delta\geq 3$ be an integer. For any $\delta\in(0,1)$, there exists $\varepsilon=\varepsilon(\delta)>0$ such that for any graph $G$ of maximum degree at most $\Delta$, we have $Z(\boldsymbol{\lambda})\neq 0$ whenever $\lambda_{v}\in\mathcal{D}([0,(1-\delta)\lambda_{c}(\Delta)],\varepsilon)$ for each vertex $v$.
The zero-freeness result of Theorem 9.14 yielded an $\mathsf{FPTAS}$ for partition function for all graphs of maximum degree $\Delta$ when $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$ for $\delta>0$ via the algorithm of Patel and Regts [PR17] which builds upon the Taylor polynomial interpolation method introduced by Barvinok [Bar15, Bar16b, Bar16a, PR22].
We also need the following lemma from complex analysis. Recall, a holomorphic function on an open set $U\subseteq\mathbb{C}$ is differentiable at every point in $U$.
Lemma 9.15 (Schwarz–Pick Lemma).
Let $f:\mathcal{D}(0,1)\to\overline{\mathcal{D}}(0,1)$ be a holomorphic function. Then
$|f^{\prime}(0)|\leq 1-|f(0)|^{2}\leq 1.$
9.3.2 Proof approach
We need to show that for any graph $G$ of maximum degree at most $\Delta$ and any vertex $u\in V$, it holds
$\sum_{v\in V}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|=O(1).$
At a high level, the proof proceeds in the following manner:
  1. 1.
    Consider the multivariate case where every vertex has its own fugacity. For a complex number $\zeta\in\mathbb{C}$, define $\boldsymbol{\lambda}(\zeta)$ to be some perturbation of the fugacity vector such that:
    1. (a)
      $\boldsymbol{\lambda}(0)=\lambda\mathbf{1}$ is the uniform fugacity vector.
    2. (b)
      Consider the complex function
      $f(\zeta)=\lambda\log\big{(}R_{G}\left(u;\boldsymbol{\lambda}(\zeta)\right)\big{)},$
      where $R_{G}$ is the occupancy ratio at $u$ defined in Eq. 96. The derivative of $f$ is a rational function and Theorem 9.14 implies that its denominator does not vanish in $\mathcal{D}(0,\varepsilon)$ and hence $f$ is holomorphic on $\mathcal{D}(0,\varepsilon)$.
    3. (c)
      For a suitable choice of $\boldsymbol{\lambda}(\zeta)$ we will show, using Claim 9.2, that we have
      $f^{\prime}(0)=\sum_{v\in V}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|.$
    In the actual proof we define $f$ differently from above, so that it is easier to describe the image of $f$ as needed in the next step.
  2. 2.
    We need to show the image of $f$ is contained in $\overline{\mathcal{D}}(0,B)$ where $B>0$ is a constant. Then, Lemma 9.15 applied to the function $g(z)=\frac{1}{B}f(\varepsilon z)$, implies that
    $\sum_{v\in V}|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)|=f^{\prime}(0)=\frac{B}{\varepsilon}g^{\prime}(0)\leq\frac{B}{\varepsilon}.$
9.3.3 Proofs
Now we present the proof of Theorem 9.1; our bound here on spectral independence depends on the radius of the zero-free region that appears in Theorem 9.14 which implicitly depends on $\delta$.
Proof of (the weaker version of) Theorem 9.1.
Fix the graph $G$ and the vertex $u$. For each $v\neq u$ define
$s_{v}=\mathrm{sgn}\left(\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right):=\begin{cases}1,&\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\geq 0;\\ -1,&\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)<0.\end{cases}$
Note that $|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)|=s_{v}\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)$. We then define the perturbed fugacity vector $\boldsymbol{\lambda}(\zeta)$ by $\lambda_{v}(\zeta)=\lambda+s_{v}\zeta$ for $v\neq u$ and $\lambda_{u}(\zeta)=\lambda$. Consider the complex function
$f(\zeta)=\frac{\lambda}{R_{G,\lambda}(u)}R_{G}\big{(}u;\boldsymbol{\lambda}(\zeta)\big{)}.$
Claim 9.16.
For $\varepsilon$ given by Theorem 9.14, the complex function $f$ is holomorphic in $\mathcal{D}(0,\varepsilon)$.
Proof.
As in Eqs. 96 and 102, we can write
$R_{G}\big{(}u;\boldsymbol{\lambda}(\zeta)\big{)}=\frac{\lambda Z_{G}^{u}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}}{Z_{G}^{\bar{u}}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}}.$
For any $\zeta\in\mathcal{D}(0,\varepsilon)$, we observe that $\lambda_{v}(\zeta)=\lambda+s_{v}\zeta\in\mathcal{D}(\lambda,\varepsilon)\subseteq\mathcal{D}([0,(1-\delta)\lambda_{c}(\Delta)],\varepsilon)$ for any $v\neq u$, and hence, by Theorem 9.14, $Z_{G}^{\bar{u}}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}\neq 0$ (note that $Z_{G}^{\bar{u}}$ is the independence polynomial for the graph $G\setminus u$ of maximum degree at most $\Delta$). A rational function is holomorphic if the denominator is zero-free. Thus, $R_{G}\big{(}u;\boldsymbol{\lambda}(\zeta)\big{)}$ is holomorphic in $\mathcal{D}(0,\varepsilon)$ and so is $f$ (since the term $\lambda/R_{G,\lambda}(u)$ is a real number). ∎
Claim 9.17.
We have
$f^{\prime}(0)=\sum_{v\in V}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|.$
Proof.
By the chain rule, we have
$f^{\prime}(0)$ $=\frac{\lambda}{R_{G,\lambda}(u)}\frac{\,\mathrm{d}}{\,\mathrm{d}\zeta}R_{G}\big{(}u;\boldsymbol{\lambda}(\zeta)\big{)}\bigg{|}_{\zeta=0}$
[…]
$=\sum_{v\in V\setminus u}\left|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)\right|,$
as claimed. ∎
Claim 9.18.
The image of $f$ is contained in $\overline{\mathcal{D}}\big{(}0,\lambda^{2}/(\varepsilon R_{G,\lambda}(u))\big{)}$.
Proof.
Observe that
$R_{G}\big{(}u;\boldsymbol{\lambda}(\zeta)\big{)}=y$
$\Longleftrightarrow\qquad$ $\frac{\lambda Z_{G}^{u}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}}{Z_{G}^{\bar{u}}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}}=y$
$\Longleftrightarrow\qquad$ $\left(-\frac{\lambda}{y}\right)Z_{G}^{u}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}+Z_{G}^{\bar{u}}\big{(}\boldsymbol{\lambda}(\zeta)\big{)}=0$
$\Longleftrightarrow\qquad$ $Z_{G}\big{(}\boldsymbol{\rho}(\zeta)\big{)}=0,$
where
$\rho_{v}(\zeta)=\begin{cases}\lambda_{v}(\zeta),&v\neq u;\\ -\frac{\lambda}{y},&v=u.\end{cases}$
Hence, we deduce from Theorem 9.14 that
$\frac{\lambda}{y}\notin\mathcal{D}([0,(1-\delta)\lambda_{c}(\Delta)],\varepsilon).$
In particular,
$\frac{\lambda}{y}\notin\mathcal{D}(0,\varepsilon)\quad\Longrightarrow\quad y\in\overline{\mathcal{D}}\left(0,\frac{\lambda}{\varepsilon}\right).$
The claim then follows. ∎
Let $B=\lambda^{2}/(\varepsilon R_{G,\lambda}(u))$, and let $g(z)=\frac{1}{B}f(\varepsilon z)$. We can now conclude the proof of (the weaker version of) Theorem 9.1 as follows:
$\sum_{v\in V}|\Psi_{G,\lambda}(u\text{\scriptsize{$\,\rightarrow\,$}}v)|$ $=f^{\prime}(0)$ by Claim 9.17
$=\frac{B}{\varepsilon}g^{\prime}(0)$ by the definition of $g(z)$
$\leq\frac{B}{\varepsilon}$ by Lemma 9.15
$=\frac{\lambda^{2}}{\varepsilon^{2}R_{G,\lambda}(u)}$ by the definition of $B$
$=O(\lambda/\varepsilon^{2}),$
where the last line follows from $R_{G,\lambda}(u)\geq\lambda/(1+\lambda)^{\Delta}=\Omega(\lambda)$ when $\lambda=O(1/\Delta)$, which is the case when $\lambda$ is in the tree-uniqueness regime. This completes the proof of (the weaker version of) Theorem 9.1 as intended. ∎
Remark 9.19.
Note the above proof establishes $\eta$-spectral independence for $\eta=O(\lambda/\varepsilon^{2})$ where $\varepsilon=\varepsilon(\delta)$ is the zero-free radius that appears in Theorem 9.14.
9.4 Spectral Independence via Coupling Independence
The coupling independence approach was introduced in Chen and Zhang [CZ23], though it was influenced by earlier works utilizing disagreement percolation to establish Gibbs uniqueness [BS94] and recursive coupling to prove strong spatial mixing [GMP05]. Since its introduction, coupling independence has been widely applied due to its simplicity and convenience. Roughly speaking, it aims to construct a coupling between two conditional Gibbs distributions $\mu_{G,\lambda}^{u}$ where we fix a vertex $u$ to be occupied, and $\mu_{G,\lambda}^{\bar{u}}$ where we fix $u$ to be unoccupied, such that the expected Hamming distance of the two samples is as small as possible.
For $\sigma,\tau\in\{0,1\}^{V}$, the Hamming distance between them is
$d_{\mathrm{H}}(\sigma,\tau):=\sum_{v\in V}\mathbbm{1}\{\sigma(v)\neq\tau(v)\}=\left\lVert\sigma-\tau\right\rVert_{1}.$
The Wasserstein distance between two distributions $\mu,\nu$, denoted as $W_{1}(\mu,\nu)$, is the expected Hamming distance under the optimal coupling between $\mu$ and $\nu$. Hence, to upper bound the Wasserstein distance it suffices to provide a coupling and upper bound the expected Hamming distance under this coupling.
Definition 9.20 (Wasserstein Distance $W_{1}$).
For two distributions $\mu,\nu$ over $\{0,1\}^{V}$, the $1$-Wasserstein distance between $\mu$ and $\nu$ with respect to the Hamming distance is defined as
$W_{1}(\mu,\nu):=\inf_{\pi\in\mathcal{P}(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\pi}\left[d_{\mathrm{H}}(X,Y)\right],$
where $\mathcal{P}(\mu,\nu)$ denotes the family of all couplings of $\mu$ and $\nu$.
We can now define coupling independence.
Definition 9.21 (Coupling Independence (without pinning)).
Let $\mu$ be a distribution over $\{0,1\}^{V}$ where $V$ is a finite set. For each $u\in V$, let $\mu^{u}(\cdot)=\mu(\cdot\mid\sigma(u)=1)$ be the conditional distribution where $\sigma(u)$ is fixed to be $1$ and $\mu^{\bar{u}}(\cdot)=\mu(\cdot\mid\sigma(u)=0)$ be the conditional distribution where $\sigma(u)$ is fixed to be 0. We say $\mu$ is $C_{\mathsf{CI}}$-coupling independent if for every $u\in V$ such that $\min\{\mu(u),\mu(\bar{u})\}>0$, it holds
$W_{1}\left(\mu^{u},\mu^{\bar{u}}\right)\leq C_{\mathsf{CI}}.$ (124)
Remark 9.22.
Rigorously, just like spectral independence, one needs to define coupling independence under arbitrary pinnings, that is, Eq. 124 holds true for any conditional Gibbs distribution $\mu_{\tau}$ where $\tau:\Lambda\to\{0,1\}$ for $\Lambda\subset V$ is a valid pinning. For simplicity, we only consider the version without pinning in this section, which makes the connection clearer between coupling independence and the maximum eigenvalue of the influence matrix.
It is immediate that coupling independence implies spectral independence, as can be seen by the following lemma.
Lemma 9.23.
Let $\mu$ be a distribution over $\{0,1\}^{V}$ where $V$ is a finite set. If $\mu$ is $C_{\mathsf{CI}}$-coupling independent, then
$\lambda_{\max}(\Psi_{\mu})\leq C_{\mathsf{CI}}.$
Proof.
Fix $u\in V$. Since $V$ is finite, there is a coupling $\pi$ achieving $W_{1}\left(\mu^{u},\mu^{\bar{u}}\right)$, such that
$\mathbb{E}_{(X,Y)\sim\pi}\left[d_{\mathrm{H}}(X,Y)\right]\leq C_{\mathsf{CI}}.$
It follows that
$\sum_{v\in V}|\Psi_{\mu}(u\text{\scriptsize{$\,\rightarrow\,$}}v)|$ $=\sum_{v\in V}|\mu^{u}(v)-\mu^{\bar{u}}(v)|$
$=\sum_{v\in V}\left|\mathbb{E}_{\pi}[X(v)]-\mathbb{E}_{\pi}[Y(v)]\right|$
$\leq\sum_{v\in V}\mathbb{E}_{\pi}\left[\left|X(v)-Y(v)\right|\right]$
$=\mathbb{E}_{\pi}\left[d_{\mathrm{H}}(X,Y)\right]\leq C_{\mathsf{CI}}.$
Hence, we conclude that $\lambda_{\max}(\Psi_{\mu})\leq\left\lVert\Psi_{\mu}\right\rVert_{\infty}\leq C_{\mathsf{CI}}$ by Lemma 2.3. ∎
Thus, to establish spectral independence, it suffices to construct a coupling between $\mu^{u}$ and $\mu^{\bar{u}}$ which has small Hamming distance in expectation.
For the hard-core model, Chen and Feng [CF25] showed that coupling independence holds up to the uniqueness threshold, and hence yields an alternative proof of Theorem 9.1.
Theorem 9.24 ([CF25]).
Let $\Delta\geq 3$ and $\delta\in(0,1)$. For any graph $G=(V,E)$ of maximum degree at most $\Delta$, any vertex $u\in V$, and any $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, it holds
$W_{1}\left(\mu_{G,\lambda}^{u},\mu_{G,\lambda}^{\bar{u}}\right)=O\left(\frac{1}{\delta}\right).$
The above result of [CF25] utilizes correlation-decay arguments based on self-avoiding walk trees, similar to what is done in Section 9.2.
Since its introduction, coupling independence has seen great success in proving spectral independence for general multi-spin systems due to its simple combinatorial nature. Besides the hard-core model, example applications also include the ferromagnetic Ising model [CZ23], $k$-SAT [CGG+24], $b$-matchings [CG24], vertex colorings [CLMM23, CF25, CFG+24], and edge colorings [CWZZ25].
In addition to implying spectral independence and hence optimal mixing of the Glauber dynamics, coupling independence has also found applications in deterministic approximate counting algorithms [CFG+24]. In particular, Chen, Feng, Guo, Zhang, and Zou [CFG+24, Theorem 4] showed that if $C$-coupling independence (under arbitrary pinnings) and $b$-marginal boundedness hold then there is an $\mathsf{FPTAS}$ for the partition function where the polynomial exponent in the running time depends on the constants $C$ and $b$ and also on the maximum degree $\Delta$ of the input graph.
In the following theorem, we prove coupling independence holds for the hard-core model for sufficiently small $\lambda$ as an illustrative example. Unlike Theorem 9.1 which we proved earlier, the following Theorem 9.25 works for fugacity strictly below the tree-uniqueness threshold $\lambda_{c}(\Delta)$.
Theorem 9.25.
Let $\Delta\geq 3$ and $\delta\in(0,1)$. For any graph $G=(V,E)$ of maximum degree at most $\Delta$, any vertex $u\in V$, and any $\lambda\leq(1-\delta)/(\Delta-2)$, it holds
$W_{1}\left(\mu_{G,\lambda}^{u},\mu_{G,\lambda}^{\bar{u}}\right)\leq\frac{1+2\lambda}{\delta}.$
Proof.
Throughout the proof, we omit $\lambda$ in the subscript of the Gibbs distribution for simplicity. We first prove that for any graph $G$ of maximum degree at most $\Delta$ and any vertex $u\in V$ of degree $\leq\Delta-1$, it holds
$W_{1}\left(\mu_{G}^{u},\mu_{G}^{\bar{u}}\right)\leq\frac{1+\lambda}{\delta}.$ (125)
We prove Eq. 125 by induction on the number of vertices. If $|V|=1$ then Eq. 125 is trivial. Now suppose Eq. 125 holds true for any graph $G$ of maximum degree at most $\Delta$ and size at most n1n-1. Fix an $n$-vertex graph $G$ of maximum degree at most $\Delta$ and a vertex $u\in V$ of degree $\leq\Delta-1$. Notice that if $\sigma(u)=1$ then $\sigma(v)=0$ for all $v\in N$ where $N=N(u)=\{v\in V:\{u,v\}\in E(G)\}$ is the set of neighbors of $u$. Meanwhile, if $\sigma(u)=0$ then we can decompose the conditional distribution $\mu_{G}^{\bar{u}}$ as
$\mu_{G}^{\bar{u}}(\cdot)=\sum_{\tau\in\{0,1\}^{N}}\mu_{G}^{\bar{u}}(\tau)\mu_{G}^{\bar{u},\tau}(\cdot),$
where $\mu_{G}^{\bar{u}}(\tau)$ is the probability that the neighborhood $N$ of $u$ receives the configuration $\tau$, and $\mu_{G}^{\bar{u},\tau}(\cdot)$ is the conditional Gibbs distribution where we fix $\sigma(u)=0$ and $\sigma(N)=\tau$. Let $H=G\setminus\{u\}$ denote the induced subgraph obtained by removing the vertex $u$. We then deduce that
$W_{1}\left(\mu_{G}^{u},\mu_{G}^{\bar{u}}\right)$ $\leq\sum_{\tau\in\{0,1\}^{N}}\mu_{G}^{\bar{u}}(\tau)\cdot W_{1}\left(\mu_{G}^{u,\mathbf{0}},\mu_{G}^{\bar{u},\tau}\right)$
$=1+\sum_{\tau\in\{0,1\}^{N}}\mu_{G}^{\bar{u}}(\tau)\cdot W_{1}\left(\mu_{H}^{\mathbf{0}},\mu_{H}^{\tau}\right),$
where we write $\mu_{G}^{u,\mathbf{0}}=\mu_{G}^{u}$ to emphasize that the neighborhood $N$ receives the all-zero configuration $\mathbf{0}$. Now for every feasible $\tau\in\{0,1\}^{N}$, we construct a sequence of configurations on $N$ denoted as $\mathbf{0}=\tau^{(0)},\tau^{(1)},\dots,\tau^{(k)}=\tau$ by flipping 0’s to $1$’s one at a time, where $k=d_{\mathrm{H}}(\mathbf{0},\tau)=|\tau|$. Hence, we get
$W_{1}\left(\mu_{H}^{\mathbf{0}},\mu_{H}^{\tau}\right)\leq\sum_{i=1}^{k}W_{1}\left(\mu_{H}^{\tau^{(i-1)}},\mu_{H}^{\tau^{(i)}}\right)\leq\frac{(1+\lambda)k}{\delta},$
where the last inequality follows from the induction hypothesis since $\tau^{(i-1)}$ and $\tau^{(i)}$ differ by one vertex of degree $\leq\Delta-1$ in the subgraph $H=G\setminus\{u\}$. Therefore, we conclude that
$W_{1}\left(\mu_{G}^{u},\mu_{G}^{\bar{u}}\right)\leq 1+\sum_{\tau\in\{0,1\}^{N}}\mu_{G}^{\bar{u}}(\tau)\cdot\frac{(1+\lambda)|\tau|}{\delta}\leq 1+\frac{1+\lambda}{\delta}\cdot(\Delta-1)\cdot\frac{\lambda}{1+\lambda}\leq\frac{1+\lambda}{\delta}.$
Finally, if the degree of $u$ is $\Delta$, then the same argument as above shows
$W_{1}\left(\mu_{G}^{u},\mu_{G}^{\bar{u}}\right)\leq 1+\frac{1+\lambda}{\delta}\sum_{\tau\in\{0,1\}^{N}}\mu_{G}^{\bar{u}}(\tau)\cdot|\tau|\leq 1+\frac{1+\lambda}{\delta}\cdot\Delta\cdot\frac{\lambda}{1+\lambda}\leq\frac{1+2\lambda}{\delta},$
as claimed. ∎
We end this subsection by showing that if there exists a contractive coupling of a Markov chain, then coupling independence and thus spectral independence hold. The following lemma states that for any Markov chain, not necessarily the Glauber dynamics, a contractive coupling implies an upper bound on the Wasserstein distance.
Lemma 9.26 ([CFG+24, Lemma 19]).
Let $\mu,\nu$ be distributions supported on $\Omega\subseteq[q]^{V}$ where $q\in\mathbb{N}^{+}$ and $V$ is finite. Let $\rho>0$ and $\kappa\in(0,1)$ be reals. Denote by $P,Q\in\mathbb{R}^{|\Omega|\times|\Omega|}$ the transition matrices of (not necessarily irreducible) Markov chains with stationary distributions $\mu,\nu$ respectively. Suppose the following two conditions are satisfied:
  1. 1.
    For any $\sigma\in\Omega$, it holds
    $W_{1}\left(P(\sigma,\cdot),Q(\sigma,\cdot)\right)\leq\rho;$ (126)
  2. 2.
    ($\kappa$-contractive coupling of $Q$) For any $\sigma,\tau\in\Omega$, it holds
    $W_{1}\left(Q(\sigma,\cdot),Q(\tau,\cdot)\right)\leq(1-\kappa)d_{\mathrm{H}}(\sigma,\tau).$ (127)
Then, we have
$W_{1}\left(\mu,\nu\right)\leq\frac{\rho}{\kappa}.$
Proof.
Since the Wasserstein distance is a metric, we deduce from the triangle inequality that
$W_{1}\left(\mu,\nu\right)=W_{1}\left(\mu P,\nu Q\right)\leq W_{1}\left(\mu P,\mu Q\right)+W_{1}\left(\mu Q,\nu Q\right).$ (128)
For the first term, we observe that
$W_{1}\left(\mu P,\mu Q\right)\leq\sum_{\sigma\in\Omega}\mu(\sigma)W_{1}\left(P(\sigma,\cdot),Q(\sigma,\cdot)\right)\leq\rho,$ (129)
which follows from Eq. 126. For the second term, $\kappa$-contractive coupling Eq. 127 implies that
$W_{1}\left(\mu Q,\nu Q\right)\leq\sum_{\sigma,\tau}\pi(\sigma,\tau)W_{1}\left(Q(\sigma,\cdot),Q(\tau,\cdot)\right)\leq(1-\kappa)\sum_{\sigma,\tau}\pi(\sigma,\tau)d_{\mathrm{H}}(\sigma,\tau),$
where $\pi$ is an arbitrary coupling of $\mu$ and $\nu$. Taking infimum over $\pi$, we thus obtain
$W_{1}\left(\mu Q,\nu Q\right)\leq(1-\kappa)W_{1}\left(\mu,\nu\right).$ (130)
Combining Eqs. 128, 129 and 130, the lemma then follows. ∎
We remark that Lemma 9.26 holds more generally for any metric space $(\Omega,d)$ with the Wasserstein distance defined with respect to the metric $d$; see [CFG+24].
We provide here a simple illustration of the above techniques. We show that there is a contractive coupling for the hard-core model when $\lambda$ is sufficiently small, and then we can apply Lemma 9.26 to conclude coupling independence.
Lemma 9.27.
Let $\Delta\geq 3$ and $\delta\in(0,1)$. For any $n$-vertex graph $G=(V,E)$ of maximum degree at most $\Delta$ and any $\lambda\leq(1-\delta)/(\Delta-1)$, there is a $\kappa$-contractive coupling of the Glauber dynamics with stationary distribution $\mu_{G,\lambda}$ where $\kappa=\delta/((1+\lambda)n)$. Specifically, for any two independent sets $\sigma,\tau$ of $G$ it holds
$W_{1}\left(P_{\textsc{GD}}(\sigma,\cdot),P_{\textsc{GD}}(\tau,\cdot)\right)\leq\left(1-\frac{\delta}{(1+\lambda)n}\right)d_{\mathrm{H}}(\sigma,\tau).$
Proof.
We use path coupling to establish a contractive coupling; we refer the reader to [Jer03, Chapter 4] and [LP17, Chapter 14.2] for an introduction to path coupling.
We first consider the case where $d_{\mathrm{H}}(\sigma,\tau)=1$, i.e., $\sigma$ and $\tau$ differ at only one vertex which we denote as $z$. We construct an identity coupling $\pi$ for the transition $(\sigma,\tau)\to(X,Y)$ where $X\sim P_{\textsc{GD}}(\sigma,\cdot)$ and $Y\sim P_{\textsc{GD}}(\tau,\cdot)$. More specifically, from $(\sigma,\tau)$, we choose a random vertex $v$ for update in both chains, and then with probability $\lambda/(1+\lambda)$ both chains attempt to add $v$ to the current independent set (it might succeed in both chains, only one of the chains, or in neither chain), and with probability $1/(1+\lambda)$ we remove $v$ from both chains.
Note that if the updated vertex $v=z$ then $X=Y$ with probability 1, since $\sigma(w)=\tau(w)$ for all $w\in N(z)$. If $v\in V\setminus(\{z\}\cup N(z))$, then similarly $d_{\mathrm{H}}(X,Y)=1$ with probability 1. Finally, if $v\in N(z)$ then with probability $\leq\lambda/(1+\lambda)$ the Hamming distance increases by 1, and otherwise remains the same. Therefore, under this identity coupling $\pi$ we deduce that
$W_{1}\left(P_{\textsc{GD}}(\sigma,\cdot),P_{\textsc{GD}}(\tau,\cdot)\right)-1\leq\frac{1}{n}\left(\frac{\Delta\lambda}{1+\lambda}-1\right)=\frac{(\Delta-1)\lambda-1}{(1+\lambda)n}\leq-\frac{\delta}{(1+\lambda)n},$
which establishes a $\delta/((1+\lambda)n)$-contraction when $d_{\mathrm{H}}(\sigma,\tau)=1$.
For the general case, we construct a sequence of independent sets $\sigma=\sigma^{(0)},\sigma^{(1)},\dots,\sigma^{(k)}=\tau$ where $k=d_{\mathrm{H}}(\sigma,\tau)$, obtained by first removing vertices in $\sigma\setminus\tau$ one by one to reach $\sigma\cap\tau$, and then adding vertices in $\tau\setminus\sigma$ one by one. Note that $d_{\mathrm{H}}(\sigma^{(i-1)},\sigma^{(i)})=1$ for each $i$. It follows from the triangle inequality that
$W_{1}\left(P_{\textsc{GD}}(\sigma,\cdot),P_{\textsc{GD}}(\tau,\cdot)\right)\leq\sum_{i=1}^{k}W_{1}\left(P_{\textsc{GD}}(\sigma^{(i-1)},\cdot),P_{\textsc{GD}}(\sigma^{(i)},\cdot)\right)\leq\sum_{i=1}^{k}(1-\kappa)=(1-\kappa)d_{\mathrm{H}}(\sigma,\tau),$
as claimed. ∎
By Lemmas 9.26 and 9.27, we deduce the following coupling independence result, where the regime of $\lambda$ is slightly worse than Theorem 9.25.
Theorem 9.28.
Let $\Delta\geq 3$ and $\delta\in(0,1)$. For any graph $G=(V,E)$ of maximum degree at most $\Delta$, any vertex $u\in V$, and any $\lambda\leq(1-\delta)/(\Delta-1)$, it holds
$W_{1}\left(\mu_{G,\lambda}^{u},\mu_{G,\lambda}^{\bar{u}}\right)\leq\frac{1+\lambda}{\delta}.$
Proof.
Let $\Omega$ denote the set of all independent sets of $G$. Let $P,Q\in\mathbb{R}^{\Omega\times\Omega}$ be the transition matrices of the Glauber dynamics with stationary distributions $\mu_{G,\lambda}^{u}$ and $\mu_{G,\lambda}^{\bar{u}}$ respectively. Note that $P,Q$ are reducible; their transition rules are the same as the Glauber dynamics for $\mu_{G,\lambda}$ when updating a vertex $v\neq u$, while $P$ forces $\sigma(u)=1$ and $Q$ forces $\sigma(u)=0$ when $u$ is picked to be updated. Therefore, for any independent set $\sigma\in\Omega$, we have $W_{1}(P(\sigma,\cdot),Q(\sigma,\cdot))\leq 1/n$, establishing Eq. 126. Meanwhile, Lemma 9.27 shows a $\kappa$-contractive coupling Eq. 127 for $Q$ with $\kappa=\delta/((1+\lambda)n)$. The theorem then follows from Lemma 9.26. ∎
9.5 Optimal relaxation time implies SI
In this section, we show the result of Anari, Jain, Koehler, Pham, and Vuong [AJK+24] that an optimal relaxation time for the Glauber dynamics, namely $O(n)$ relaxation time, implies spectral independence; this was stated in Theorem 1.6 of Section 1.7, which we restate now for convenience.
See 1.6
Proof.
Recall (see equation Eq. 10) that (C1)(C-1)-spectral independence is equivalent to the following inequality:
$\mathrm{Cov}_{\mu}\preceq CD,$ (131)
where $D$ is diagonal matrix with $D(i,i)=\mu(\sigma(i)=1)\mu(\sigma(i)=0)$. That is, for any $y\in\mathbb{R}^{n}$ we need to show $y^{T}\mathrm{Cov}_{\mu}y\leq Cy^{T}Dy$.
Recall (see Definition 3.4 and Corollary 3.5) that relaxation time being bounded by $Cn$ implies that for any $f:\{0,1\}^{n}\rightarrow\mathbb{R}$ we have
$\mathrm{Var}_{\mu}(f)\leq C\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)].$ (132)
Let $f(\sigma)=\sum_{v\in V}y_{v}\sigma(v)$. Note that we have
$\mathrm{Var}_{\mu}(f)=y^{T}\mathrm{Cov}_{\mu}y$ (133)
and
$\mathrm{Var}_{v}(f)=y_{v}^{2}\,\mathbb{E}_{\mu}[\chi(\mathbb{E}[X_{v}|X_{1},\dots,X_{v-1},X_{v+1},\dots,X_{n}])],$
where $\chi(x)=x(1-x)$. Since x(1x)x(1-x) is concave we have
$\mathbb{E}_{\mu}[\chi(\mathbb{E}[X_{v}|X_{1},\dots,X_{v-1},X_{v+1},\dots,X_{n}]]$ $\leq\chi(\mathbb{E}_{\mu}[\mathbb{E}[X_{v}|X_{1},\dots,X_{v-1},X_{v+1},\dots,X_{n}]])$
$=\mu(\sigma(v)=1)\mu(\sigma(v)=0),$
and hence $\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]\leq y_{v}^{2}\mu(\sigma(v)=1)\mu(\sigma(v)=0)$. Plugging this into (132) (and applying Eq. 133) we obtain $y^{T}\mathrm{Cov}_{\mu}y\leq Cy^{T}Dy$. Since we can take an arbitrary $y$, we established (131), and hence $C$-spectral independence. ∎
9.5.1 Contractive coupling implies relaxation time
Here, we show an example application of Theorem 1.6. We show that a contractive coupling of an arbitrary Markov chain not only implies an upper bound on the mixing time but also provides a lower bound on the spectral gap. Consequently, for a local Markov chain one obtains spectral independence, which provides an alternative proof approach compared with Lemma 9.26 which utilizes coupling independence.
Lemma 9.29.
Suppose that for a Markov chain on $\Omega$ with transition matrix $P$, we have a contractive coupling with contraction factor $\kappa\in(0,1)$; that is, we have a distance function $d:\Omega\times\Omega\rightarrow\mathbb{R}_{\geq 0}$ such that $d(\sigma,\tau)=0$ iff $\sigma=\tau$ and, for any $\sigma,\tau\in\Omega$, a coupling $\pi$ of $P(\sigma,\cdot)$ and $P(\tau,\cdot)$ such that
$\mathbb{E}_{(X,Y)\sim\pi}[d(X,Y)]\leq(1-\kappa)d(\sigma,\tau).$ (134)
Then the absolute spectral gap of $P$ satisfies $\gamma=1-\lambda_{*}\geq\kappa$, and hence the relaxation time of $P$ satisfies $T_{\mathrm{relax}}\leq 1/\kappa$.
For the Glauber dynamics we typically obtain $\kappa=\Omega(1/n)$. If a Markov chain is local in the sense that the transitions update $O(1)$ vertices in each step then optimal relaxation time for this local chain implies optimal relaxation time for the Glauber dynamics, see Remark 9.31 for further discussion and associated references.
Proof.
Let $\varphi:\Omega\rightarrow\mathbb{R}$ be the eigenvector corresponding to eigenvalue $\lambda$, which is the second largest in absolute value and hence corresponds to $\lambda_{*}$. Denote
$L=\max_{\begin{subarray}{c}\sigma,\tau\in\Omega\\ \sigma\neq\tau\end{subarray}}\frac{|\varphi(\sigma)-\varphi(\tau)|}{d(\sigma,\tau)},$
and let $\sigma,\tau$ be the maximizers; hence, $|\varphi(\sigma)-\varphi(\tau)|=L\cdot d(\sigma,\tau)$. Therefore, we have the following:
$|\varphi(X)-\varphi(Y)|\leq L\cdot d(X,Y).$ (135)
Note that since $\phi$ is an eigenvector with eigenvalue $\lambda$, we have
$\mathbb{E}[\varphi(X)]=\sum_{\xi\in\Omega}P(\sigma,\xi)\varphi(\xi)=(P\varphi)(\sigma)=\lambda\varphi(\sigma),$
and hence
$|\lambda|\cdot|\varphi(\sigma)-\varphi(\tau)|$ $=\big{|}\mathbb{E}[\varphi(X)-\varphi(Y)]\big{|}$
$\leq\mathbb{E}\left[\big{|}\varphi(X)-\varphi(Y)\big{|}\right]$
$\leq L\cdot\mathbb{E}[d(X,Y)]$ by Eq. 135
$\leq L\cdot(1-\kappa)d(\sigma,\tau)$ by Eq. 134
$=(1-\kappa)|\varphi(\sigma)-\varphi(\tau)|,$
and thus $|\lambda|\leq 1-\kappa$ which proves the lemma. ∎
Remark 9.30.
As noted above, for the Glauber dynamics, in the context of Lemma 9.29 we typically obtain $\kappa=\alpha/n$ for some constant $\alpha>0$; see e.g. Lemma 9.27. Note, a contractive coupling with such a contraction factor of $\kappa=\alpha/n$ implies a mixing time bound of $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$. The more general statement is that $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$ implies $T_{\mathrm{relax}}=O(n)$ holds; we refer the interested reader to Levin and Peres [LP17, Chapter 14 notes] for a proof of the more general implication.
Remark 9.31.
Recall, Lemma 9.29 applies to any Markov chain. If the Markov chain is local in the sense that only $O(1)$ vertices are updated at each step, then $O(n)$ relaxation time for this local chain implies $O(n)$ relaxation time for the Glauber dynamics, see [BCC+22, Section 4.3] for further details. By then applying Theorem 1.6 we obtain spectral independence and hence $O(n\log{n})$ mixing time for the Glauber dynamics for constant $\Delta$, via Theorem 1.4. This improves upon standard comparison arguments (comparing Dirichlet forms) which use $O(n)$ relaxation time of a local Markov chain to conclude $O(n^{2})$ mixing time of the Glauber dynamics.
10 Critical Point Analysis using the Field Dynamics
In this section we prove that the Glauber dynamics has polynomial mixing time at the critical point $\lambda=\lambda_{c}(\Delta)$ as stated in Theorem 1.11. This theorem was established by Chen, Chen, Yin, and Zhang [CCYZ25, Theorem 1.1] who proved that the exponent satisfies $C\leq 2+4e+4e/(\Delta-2)$; we simply show there exists a universal constant.
We restate Theorem 1.11 for convenience.
See 1.11
The above result follows as a corollary of the following theorem, which says that to establish polynomial mixing at a particular $\lambda^{*}$ we need two properties: (1) for all $\alpha<1$, $C_{SI}(\alpha)$-spectral independence holds for $\lambda=\alpha\lambda^{*}$, and (2) fast mixing at some $\lambda^{\prime}=\theta\lambda$ for some choice of $\theta<1$; then one obtains relaxation time $O(K(\theta)n)$ where $K(\theta)$ is defined in Eq. 136 below. Note to obtain a polynomial upper bound on $K(\theta)$ and deduce Theorem 1.11, we further use that $C_{SI}(\alpha)=O(1/\alpha)$.
Theorem 10.1.
Let $\lambda^{*}>0$. If the following hold:
  1. 1.
    For all  $0\leq\alpha<1$, there exists $C_{SI}(\alpha)$ such that the measure $\mu_{(1-\alpha)\lambda^{*}}$ is $C_{SI}(\alpha)$-spectrally independent,
  2. 2.
    there exists $0\leq\theta<1$ where the Glauber dynamics for the measure $\mu_{(1-\theta)\lambda^{*}}$ has relaxation time $\leq C_{Gl}n$ (on any graph with maximum degree $\Delta$),
then the relaxation time of the Glauber dynamics for the hard-core model at fugacity $\lambda^{*}$ for any graph of maximum degree $\Delta$ satisfies
$T_{\mathrm{relax}}\leq\left(K(\theta)C_{Gl}\right)n,$
where
$K(\theta)=\exp\left(\int_{0}^{\theta}\frac{C_{SI}(\alpha)}{1-\alpha}d\alpha\right).$ (136)
We will use Theorem 10.1 to prove the main theorem (Theorem 1.11) using Theorem 10.1. In Section 10.1 we will give intuition for Theorem 10.1 and then we will dive into its proof.
Proof of Theorem 1.11.
Applying Lemma 9.27 with Corollary 3.5 we obtain Item 2 in the hypotheses of Theorem 10.1 where $\theta$ is a constant such that $\theta\lambda_{c}(\Delta)=(1/2)/(\Delta-1)$ and use $\delta=1/2$ in Lemma 9.27. For Item 1 we use Theorem 9.1 in Section 9.2, which yields the following bound
$C_{SI}(\alpha)\leq\min\Big{\{}n,\frac{c^{*}}{\alpha}\Big{\}},$
where $c^{*}=16\sqrt{2}$ (see Theorem 9.1) and the trivial upper bound of $n$ follows from the maximum row sum of the influence matrix. Plugging in Eq. 136 we obtain the following bound
$\log K(\theta)=\int_{0}^{\theta}\frac{C_{SI}(\alpha)}{1-\alpha}d\alpha$ $\leq\int_{0}^{1/n}\frac{n}{1-\alpha}d\alpha+\int_{1/n}^{\theta}\frac{c^{*}}{\alpha(1-\alpha)}d\alpha$
$\leq\frac{1}{1-\theta}+\frac{c^{*}}{1-\theta}\ln n,$
and hence $K(\theta)=O(n^{C})$ where $C=\frac{c^{*}}{1-\theta}$. The result now follows from Theorem 10.1. ∎
Remark 10.2.
Note when $\lambda<(1-\delta)\lambda_{c}(\Delta)$ for $\delta>0$ then, using Theorem 10.1, one obtains $T_{\mathrm{relax}}\leq C(\delta)n$, hence the relaxation time is independent of $\Delta$. Let $\lambda^{*}=(1-\delta)\lambda_{c}(\Delta)$. Let $C_{SI}$ be the bound on spectral independence for $\lambda^{*}=\lambda_{c}$ and $C^{\prime}_{SI}$ be the bound on spectral independence for $\lambda^{*}=(1-\delta)\lambda_{c}$. In the application of Theorem 10.1 we have
$C^{\prime}_{SI}(\alpha)=C_{SI}(1-(1-\alpha)(1-\delta))\leq\frac{c^{*}}{\delta+\alpha-\alpha\delta}.$
And then we obtain:
$\log K(\theta)=\int_{0}^{\theta}\frac{C^{\prime}_{SI}(\alpha)}{1-\alpha}d\alpha$ $\leq C^{\prime}\log(1/\delta),$
for some constant $C^{\prime}>0$. This yields $T_{\mathrm{relax}}\leq C(\delta)n$ where $C(\delta)=\mathrm{poly}(1/\delta)$.
10.1 Field Dynamics
In the remainder of Section 10 we will prove Theorem 10.1. To establish mixing at our target fugacity $\lambda^{*}$ we will use a bootstrapping approach, which utilizes Item 2 in the hypotheses of Theorem 10.1 which says that we have fast mixing of the Glauber dynamics at $\lambda^{\prime}=\theta\lambda^{*}$ for $0<\theta<1$; we satisfy this hypothesis by using an easy argument such as the coupling argument in Lemma 9.27. Before formally defining the field dynamics, which is an important tool in the proof of Theorem 10.1, we first attempt to motivate its definition and how it enables a bootstrapping approach.
Given fast mixing of the Glauber dynamics at $\lambda^{\prime}=\theta\lambda^{*}$ we want to deduce fast mixing at $\lambda^{*}$. Additionally we know that spectral independence holds for all $\lambda<\lambda^{*}$, by Item 1 of the hypotheses of Theorem 10.1. As in the proof of Theorem 1.5 we look at the down-up walk. For fixed $\alpha>0$, the $\alpha n$-block dynamics at fugacity $\lambda^{*}$, which was used in the proof of Theorem 1.5, operates in the following manner: we first apply $\alpha n$ steps of the down-walk in which at each step we remove the configuration at a randomly chosen vertex $v$, then we apply $\alpha n$-steps of the up-walk which for each of these chosen vertices, we resample the configuration at $v$ from the Gibbs distribution at $\lambda^{*}$ conditional on the current configuration; this later resampling can equivalently be done simultaneously (for all $\alpha n$ vertices) or one-by-one.
We want to modify the above block dynamics so that in the resampling stage (up-walk steps) we are resampling from the Gibbs distribution at fugacity $\lambda^{\prime}=(1-\theta)\lambda^{*}$ for $\theta<1$. To this end, in the corresponding up-walk steps we will condition on a set of vertices which are occupied (in the independent set) and we will resample all of the remaining vertices (those dropped in the down-walk steps and those which are unoccupied initially). Consequently in the corresponding up-walk steps we will sample from the smaller $\lambda^{\prime}$ as we have conditioned on a set of occupied vertices (and have not conditioned on any unoccupied vertices) so we are biasing the distribution. The following Markov chain is known as the field dynamics and was introduced by Chen, Feng, Yin, and Zhang [CFYZ22].
In words, the field dynamics operates as follows. Consider an initial independent set $Y_{t}$; note, we are considering $Y_{t}\subseteq V$, we do not consider $Y_{t}$ to be in $\{0,1\}^{V}$. We first remove each $v\in Y_{t}$ with probability $(1-\theta)$. Let $Y^{\prime}$ be the remaining vertices, since $Y^{\prime}\subseteq Y_{t}$ thus $Y^{\prime}$ is an independent set. We then resample the configuration on $V\setminus Y^{\prime}$ from the Gibbs distribution at fugacity $\lambda^{\prime}=(1-\theta)\lambda^{*}$ conditional on all vertices in $Y^{\prime}$ being in the independent set. In the boundary case when $\theta=1$ then $Y^{\prime}=Y_{t}$ and in the resampling step we have $\lambda^{\prime}=0$, whereas when $\theta=0$ then $Y^{\prime}=\emptyset$ and we resample with $\lambda^{\prime}=\lambda^{*}$.
We now formally define the field dynamics following [CFYZ22, Example 2.1].
Definition 10.3.
The (discrete-time) field dynamics with parameters $(\lambda,\theta)$ where $\lambda>0$ and $0\leq\theta\leq 1$ has the following transitions $Y_{t}\rightarrow Y_{t+1}$. From an independent set $Y_{t}$,
  1. 1.
    Let $Y^{\prime}=\emptyset$.
  2. 2.
    Independently, for each $v\in Y_{t}$, with probability $\theta$ add $v$ to set $Y^{\prime}$.
  3. 3.
    Sample an independent set $\widehat{Y}$ on the set $V\setminus(Y^{\prime}\cup N(Y^{\prime}))$ from the hard-core measure with fugacity $\lambda^{\prime}=(1-\theta)\lambda$.
    (This is equivalent to conditioning on $Y^{\prime}$ being in the independent set and sampling the remaining configuration from $\mu_{(1-\theta)\lambda}$.)
  4. 4.
    Let $Y_{t+1}=Y^{\prime}\cup\widehat{Y}$.
We first state the observation that the field dynamics has the desired stationary distribution.
Lemma 10.4.
The field dynamics is reversible with respect to $\mu_{\lambda}$.
We will prove the above lemma momentarily.
Chen et al. [CFYZ22] utilize the field dynamics to establish a boosting (or inductive) statement, which expresses the spectral gap of the Glauber dynamics at $\lambda^{*}$ as (roughly) the product of spectral gap of the field dynamics and the spectral gap of the Glauber dynamics at $\lambda^{\prime}$; the first term (gap of the field dynamics) can be bounded using Item 1 of Theorem 10.1 and the second term (gap of the Glauber dynamics at $\lambda^{\prime}$) is given by Item 2.
To obtain mixing at the critical point we need a more refined approach. We want to consider the mixing properties of the field dynamics at an infinitesimally small time and then we can integrate over time. To enable us to take the derivative with respect to time we need to introduce a continuous-time analog of the field dynamics.
We now introduce two continuous-time processes, a down process and an up process, which together generalize the discrete-time field dynamics. In the following our inital configuration will be an independent set $X_{1}$. For $\theta$ where $0\leq\theta\leq 1$ we will define the down process $X_{1}\rightarrow X_{\theta}$ which is a generalization of the process $Y_{t}\rightarrow Y^{\prime}$ in the first two steps of the field dynamics. We then define the up process $X_{\theta}\rightarrow X^{\prime}_{1}$ which generalizes $Y^{\prime}\rightarrow Y_{t+1}$.
Definition 10.5.
Let $X_{1}$ be from $\mu_{\lambda}$.
  1. 1.
    Independently, for every vertex $v$ choose a uniformly random number $\theta_{v}$ in $[0,1]$.
  2. 2.
    For every $\theta\in[0,1]$, let $X_{\theta}$ be the set of vertices in $X$ that have $\theta_{x}<\theta$.
Let $0\leq\theta\leq\alpha\leq 1$. The distribution of $X_{\theta}$ conditioned on $X_{\alpha}$ is denoted by $D_{\alpha\rightarrow\theta}(X_{\alpha},\cdot)$ and is referred to as the down process. The distribution of $X_{\alpha}$ conditioned on $X_{\theta}$ is denoted by $U_{\theta\rightarrow\alpha}(X_{\theta},\cdot)$ and is referred to as the up process.
Note that the down process $D_{1\rightarrow\theta}$ is the following. For $Y^{\prime}\subseteq Y$
$\mathrm{Pr}(X_{\theta}=Y^{\prime}|X_{1}=Y)$ $\propto\mathrm{Pr}\big{(}\forall v\in Y\setminus Y^{\prime},\theta_{v}\in[\theta,1]\big{)}\times\mathrm{Pr}\big{(}\forall v\in Y^{\prime},\theta_{v}\in[0,\theta]\big{)}$
$=(1-\theta)^{|Y\setminus Y^{\prime}|}\theta^{|Y^{\prime}|},$
that is we drop each element of $Y$ independently with probability $1-\theta$. Hence, in the discrete-time field dynamics, starting from $Y_{t}=Y$ then the distribution of $Y^{\prime}$ is identical to $D_{1\rightarrow\theta}(Y,\cdot)$.
Similarly, for the up process, we have the following:
$\mathrm{Pr}(X_{1}=Y^{\prime}\cup\widehat{Y}|X_{\theta}=Y^{\prime})$ $=\frac{\mathrm{Pr}(X_{1}=Y^{\prime}\cup\widehat{Y}\wedge X_{\theta}=Y^{\prime})}{\mathrm{Pr}(X_{\theta}=Y^{\prime})}$
$\propto\lambda^{|Y^{\prime}|}\mathrm{Pr}\big{(}\forall v\in\widehat{Y},\theta_{v}\in[\theta,1]\big{)}\mathrm{Pr}\big{(}\forall v\in Y^{\prime},\theta_{v}\in[0,\theta]\big{)}$
$\propto(\lambda(1-\theta))^{|\widehat{Y}|}$.
Therefore, if we consider the discrete-time field dynamics, given $Y^{\prime}$ at the end of step 2 (after dropping each vertex $v\in Y_{t}$ with probability $1-\theta$), then the distribution of $Y_{t+1}$ for the discrete-time field dynamics is identical to $X_{1}$ conditional on $X_{\theta}=Y^{\prime}$.
In summary, a transition of the discrete-time field dynamics is a step of the $D_{1\rightarrow\theta}$ process followed by a step of the $U_{\theta\rightarrow 1}$ process (starting the discrete-time process from $Y_{t}$ we condition on $X_{1}=Y_{t}$ and use the down process to obtain $X_{\theta}$, then we resample $X^{\prime}_{1}$ using the up process from $X^{\prime}_{\theta}=X_{\theta}$, and let $Y_{t+1}=X^{\prime}_{1}$).
As promised earlier, the proof of Lemma 10.4 is now straightforward to establish.
Proof of Lemma 10.4.
Let $P_{FD}$ denote the transition matrix of the discrete-time field dynamics. We have the ergodic flow from $x$ to $y$ is
$\mu_{\lambda}(x)P_{FD}(x,y)$ $=\mu_{\lambda}(x)(D_{1\rightarrow\theta}U_{\theta\rightarrow 1})(x,y)$
$=\sum_{S}\mathrm{Pr}(X_{1}=x)\mathrm{Pr}(X_{\theta}=S|X_{1}=x)\mathrm{Pr}(X_{1}=y|X_{\theta}=S)$
$=\sum_{S}\frac{1}{\mathrm{Pr}(X_{\theta}=S)}\mathrm{Pr}(X_{\theta}=S\wedge X_{1}=x)\mathrm{Pr}(X_{1}=y\wedge X_{\theta}=S).$
Note the last expression does not change if we swap $x$ and $y$ and hence the ergodic flow from $y$ to $x$ is the same. ∎
10.2 SI implies Fast Mixing of Field Dynamics
Recall, the definition of $C$-approximate tensorization of variance. Here we present the analog for the field dynamics, known as $K$-approximate conservation of variance. Whereas $C$-approximate tensorization of variance implies that the spectral gap of the Glauber dynamics is $\geq 1/C$, similarly, $K$-approximate conservation of variance implies that the field dynamics has spectral gap $\geq 1/K$.
The following definition is from [CCYZ25, Definiton 3.10 with $\phi=x^{2}$].
Definition 10.6.
Let $\lambda>0$, $0\leq\theta\leq 1$ and $K>0$. Let $\mu=\mu_{\lambda}$ and let $\mu^{\prime}$ denote the distribution of $X_{\theta}$ from the down-walk $D_{1\rightarrow\theta}$ where $X_{1}\sim\mu$. We say that $K$-approximate conservation of variance up to time $\theta$ holds if for every $f$ we have for
$\mathrm{Var}_{\mu}(f)$ $\leq K\mathbb{E}_{\sigma\sim\mu^{\prime}}\left[\mathrm{Var}_{U_{\theta\rightarrow 1}(\sigma,\cdot)}(f)\right]$ (137)
Let us decipher the right-hand side of Eq. 137. Note, that
$\mathbb{E}_{\sigma\sim\mu^{\prime}}\left[\mathrm{Var}_{U_{\theta\rightarrow 1}(\sigma,\cdot)}(f)\right]=\frac{1}{2}\sum_{\sigma\in\Omega}\mu^{\prime}(\sigma)\sum_{\eta_{1},\eta_{2}\in\Omega}U_{\theta\rightarrow 1}(\sigma,\eta_{1})U_{\theta\rightarrow 1}(\sigma,\eta_{2})\left(f(\eta_{1})-f(\eta_{2})\right)^{2}.$
In words, we first take a sample $\sigma$ from the down-walk of the field dynamics, then we consider the variance of $f$ from the up-walk starting from $X_{\theta}=\sigma$.
Let
$f_{\theta}(y):=\mathbb{E}[f(X_{1})|X_{\theta}=y],$
where the expectation is with respect to the up-walk $U_{\theta\rightarrow 1}(y,\cdot)$. Another possible way of writing Eq. 137 is
$\mathrm{Var}_{\mu}(f(X_{1}))\leq K\mathbb{E}_{X_{\theta}}\left[\mathrm{Var}[f(X_{1})|X_{\theta}]\right].$ (138)
Our main technical result in this section is the following theorem.
Theorem 10.7.
For $\lambda>0$, if for all $0\leq\alpha\leq 1$, the measure $\mu_{(1-\alpha)\lambda}$ is spectrally independent with constant $C(\alpha)$ then for any $0\leq\theta\leq 1$, variance conservation holds for the field dynamics at parameters $(\lambda,\theta)$ with constant $K=K(\theta)$ defined in Eq. 136.
Proof.
Let
$Q(\alpha):=\mathbb{E}[\mathrm{Var}[f(X_{1})|X_{\alpha}]],$
that is sample $S$ from $X_{\alpha}$ then look at the variance of $f(X_{1})$ where $X_{1}\sim U_{\alpha\rightarrow 1}(S,\cdot)$. Note, $Q(1)=0$, and $Q(0)=\mathrm{Var}_{\mu}(f)$ (since $X_{0}=\emptyset$). To prove conservation of variance, recall Eq. 138. To establish Eq. 138 the rough intuition is that we will show that $Q(\alpha)/Q(0)$ is large for all $\alpha<1$ and hence $K(\alpha)$-approximate conservation of variance holds with small $K(\alpha)$ (this is an increasing function as $\alpha\rightarrow 1$). Hence, we will prove that the log-derivative is large:
$(\log Q(\alpha))^{\prime}=\frac{Q^{\prime}(\alpha)}{Q(\alpha)}\geq-\frac{C(\alpha)}{1-\alpha}.$ (139)
We will now prove the theorem assuming Eq. 139 and then we will go back to prove Eq. 139.
Fix $0\leq\theta\leq 1$.
$(\log Q(\theta))$ $=(\log Q(0))+\int_{0}^{\theta}(\log Q(\alpha))^{\prime}\,{\mathrm{d}}\alpha\geq(\log Q(0))-\int_{0}^{\theta}\frac{C(\alpha)}{1-\alpha}\,{\mathrm{d}}\alpha.$
After exponentiating both sides, and using the fact that $Q(0)=\mathrm{Var}_{\mu}(f)$ we then have:
$\mathrm{Var}_{\mu}(f)\leq KQ(\theta),$
with $K$ given by Eq. 136. This completes the proof of the theorem modulo the proof of Eq. 139 which we now establish.
Recall, $f_{\alpha}(y):=\mathbb{E}[f(X_{1})|X_{\alpha}=y]$. Note that we have $f_{1}=f$. For $0\leq s\leq 1$, let $\nu_{s}$ denote the distribution of $X_{s}$ from the down-walk $D_{1\rightarrow s}$ where $X_{1}\sim\mu$. For $0\leq\alpha<s\leq 1$, by the law of total variance we have:
$\mathrm{Var}_{\nu_{s}}[f_{s}]=\mathrm{Var}_{\nu_{\alpha}}[f_{\alpha}]+\mathbb{E}_{\nu_{\alpha}}[\mathrm{Var}_{U_{\alpha\rightarrow s}}[f_{s}]]$ (140)
From Eq. 140 for $s=1$ we have:
$Q(\alpha)=\mathrm{Var}_{\mu}[f]-\mathrm{Var}[\mathbb{E}[f(X_{1})|X_{\alpha}]],$ (141)
For $h>0$, from Eq. 140 we have:
$\mathrm{Var}_{\nu_{\alpha+h}}[f_{\alpha+h}]=\mathrm{Var}_{\nu_{\alpha}}[f_{\alpha}]+\mathbb{E}_{S\sim\nu_{\alpha}}[\mathrm{Var}_{U_{\alpha\rightarrow\alpha+h}(S,\cdot)}[f_{\alpha+h}]].$
Note that $\mathbb{E}_{T\sim U_{\alpha\rightarrow\alpha+h}(S,\cdot)}[f_{\alpha+h}(T)]=f_{\alpha}(S)$ and hence for any $S\in\Omega$:
$\mathrm{Var}_{U_{\alpha\rightarrow\alpha+h}(S,\cdot)}[f_{\alpha+h}]$ $=\sum_{T:T\supseteq S}U_{\alpha\rightarrow\alpha+h}(S,T)\left(f_{\alpha+h}(T)-f_{\alpha}(S)\right)^{2}$
$=\frac{h}{1-\alpha}\sum_{v\in V\setminus S}\mathrm{Pr}_{R\sim U_{\alpha\rightarrow 1}(S,\cdot)}(v\in R)\left(f_{\alpha}(S\cup\{v\})-f_{\alpha}(S)\right)^{2}+O(h^{2}),$ (142)
where in Eq. 142 the factor $h/(1-\alpha)$ follows from the observation that each $v\in R$ is added at a uniformly chosen time in the interval $\alpha,1]$ (and more than one vertex is added with probability $O(h^{2})$), and the substitution of $f_{\alpha+h}(T)$ by $f_{\alpha}(T)$ with $T=S\cup\{v\}$ follows from $f_{\alpha+h}(T)=f_{\alpha}(T)+O(h)$ (see also [CCYZ25, (81) in Appendix A.1]).
Recall, $\mu=\mu_{\lambda}$. Taking the derivative of variance in Eq. 141 we obtain:
$-Q^{\prime}$ $(\alpha)=\frac{\partial}{\partial\alpha}\mathrm{Var}[f_{\alpha}]$
$=\lim_{h\rightarrow 0}\frac{1}{h}\left[\mathrm{Var}_{\nu_{\alpha+h}}[f_{\alpha+h}]-\mathrm{Var}_{\nu_{\alpha}}[f_{\alpha}]\right]$
$=\lim_{h\rightarrow 0}\frac{1}{h}\mathbb{E}_{S\sim\nu_{\alpha}}\left[\mathrm{Var}_{U_{\alpha\rightarrow\alpha+h}(S,\cdot)}[f_{\alpha+h}]\right],$ by Eq. 140,
$=\frac{1}{1-\alpha}\mathbb{E}_{S\sim\nu_{\alpha}}\left[\sum_{v\in V\setminus S}\mathrm{Pr}_{R\sim U_{\alpha\rightarrow 1}(S,\cdot)}(v\in R)\left(f_{\alpha}(S\cup\{v\})-f_{\alpha}(S)\right)^{2}\right]$ by Eq. 142
$\leq\frac{1}{1-\alpha}\mathbb{E}_{S\sim\nu_{\alpha}}\left[\sum_{v\in V\setminus S}\mathrm{Var}[\mathbb{E}[f|X_{1}[v]\,]|X_{\alpha}=S]\right],$ by Eq. 143 below,
$\leq\frac{C(\alpha)}{1-\alpha}\mathbb{E}_{S\sim\nu_{\alpha}}\left[\mathrm{Var}[f|X_{\alpha}=S]\right],$ by Eq. 21 and $C(\alpha)$-SI.
$\leq\frac{C(\alpha)}{1-\alpha}Q(\alpha),$
which establishes Eq. 139.
We will argue that the following is true (for any $v$, $S$, and $\alpha$):
$\mathrm{Pr}_{R\sim U_{\alpha\rightarrow 1}}(v\in R)\left(f_{\alpha}(S\cup\{v\})-f_{\alpha}(S)\right)^{2}\leq\mathrm{Var}[\mathbb{E}[f|X_{1}[v]\,]|X_{\alpha}=S].$ (143)
To simplify the notation let $A=\mathbb{E}[f|v\in X_{1}\,X_{\alpha}=S]$, $B=\mathbb{E}[f|v\not\in X_{1},X_{\alpha}=S\,]$, and $p=\mathrm{Pr}_{R\sim U_{\alpha\rightarrow 1}(S,\cdot)}(v\in R)$. Equation (143) follows since the LHS is
$p(A-(pA+(1-p)B))^{2}=p(1-p)^{2}(A-B)^{2},$
and the RHS is
$p(1-p)(A-B)^{2}.$
10.3 Fast Mixing of Glauber Dynamics from Field Dynamics
Proof of Theorem 10.1.
From Theorem 10.7, we have $K$-approximate conservation of variance which states the following:
$\mathrm{Var}_{\mu}[f]\leq K\,\mathbb{E}_{\nu_{\theta}}[\mathrm{Var}[f(X_{1})|X_{\theta}]]=K\,\mathbb{E}_{\nu_{\theta}}[\mathrm{Var}_{U_{\theta\rightarrow 1}(S,\cdot)}[f(X_{1})|X_{\theta}=S]].$ (144)
For $i\in[n]$, let $\mathcal{P}_{i}=X_{1}[[n]\setminus\{i\}]$ be the configuration on all vertices but $i$. We want to prove approximate tensorization of variance for $\mu$ with constant $C_{Gl}K$, that is, establish the following inequality:
$\mathrm{Var}_{\mu}[f]\leq C_{Gl}K\,\sum_{i=1}^{n}\mathbb{E}_{\mu}[\mathrm{Var}[f(X_{1})|\mathcal{P}_{i}]].$ (145)
By Item 2 in the hypothesis of Theorem 10.1 combined with Corollary 3.5, we know approximate tensorization of variance holds with constant $C_{Gl}$ for distribution $\mu^{\prime}=\mu_{\lambda^{\prime}}$ where $\lambda^{\prime}=(1-\theta)\lambda$ and $\theta>0$; hence we know the following:
$\mathrm{Var}_{\mu^{\prime}}[f]\leq C_{Gl}\sum_{i=1}^{n}\mathbb{E}_{\mu^{\prime}}[\mathrm{Var}[f(X_{1})|\mathcal{P}_{i}]].$ (146)
Moreover, Eq. 146 holds for any induced subgraph $G^{\prime}$ of $G$. For any $S\subseteq V$, considering the subgraph $G^{\prime}$ induced by conditioning on the independent set $S$ then $\mu^{\prime}$ on $G^{\prime}$ is identical to the distribution $U_{\theta\rightarrow 1}(S,\cdot)$. Hence, we have the following, for any $S\subseteq V$,
$\mathrm{Var}_{U_{\theta\rightarrow 1}(S,\cdot)}[f(X_{1})|X_{\theta}=S]\leq C_{Gl}\sum_{i=1}^{n}\mathbb{E}_{U_{\theta\rightarrow 1}(S,\cdot)}[\mathrm{Var}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]].$ (147)
Taking $S\sim\nu_{\theta}$, we have:
$\mathbb{E}_{S\sim\nu_{\theta}}[\mathrm{Var}_{U_{\theta\rightarrow 1}(S,\cdot)}[f(X_{1})|X_{\theta}=S]]\leq C_{Gl}\sum_{i=1}^{n}\mathbb{E}_{S\sim\nu_{\theta}}[\mathbb{E}_{U_{\theta\rightarrow 1}}[\mathrm{Var}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]]].$ (148)
We also need the following observation:
$\mathbb{E}_{S\sim\nu_{\theta}}[\mathbb{E}_{U_{\theta\rightarrow 1}(S,\cdot)}[\mathrm{Var}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]]]$
$=\mathbb{E}_{\mu}\Big{[}\mathbb{E}_{S\sim D_{1\rightarrow\theta}}\Big{[}\mathrm{Var}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]]\Big{|}\mathcal{P}_{i}\Big{]}\Big{]}$ (149)
$\leq\mathbb{E}_{\mu}\Big{[}\mathbb{E}_{S\sim D_{1\rightarrow\theta}}\Big{[}\mathrm{Var}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]]\Big{|}\mathcal{P}_{i}\Big{]}\Big{]}$
$\hskip 72.26999pt+\mathbb{E}_{\mu}\Big{[}\mathrm{Var}_{S\sim D_{1\rightarrow\theta}}\Big{[}\mathbb{E}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]]\Big{|}\mathcal{P}_{i}\Big{]}\Big{]}$
$=\mathbb{E}_{\mu}\Big{[}\mathrm{Var}[f(X_{1})|\mathcal{P}_{i}]]\Big{]},$ (150)
where Eq. 149 is from the observation that sampling $S$ from $\mu_{\theta}$ and then applying the up-walk $U_{\theta\rightarrow 1}$ is the same as sampling from $\mu$ and then applying the down-walk $D_{1\rightarrow\theta}$, and Eq. 150 is by the law of total variance.
Putting it all together we have:
$\mathrm{Var}_{\mu}[f]$ $\leq K\,\mathbb{E}_{\nu_{\theta}}[\mathrm{Var}_{U_{\theta\rightarrow 1}(S,\cdot)}[f(X_{1})|X_{\theta}=S]]$ by Eq. 144
$\leq C_{Gl}K\sum_{i=1}^{n}\mathbb{E}_{S\sim\nu_{\theta}}[\mathbb{E}_{U_{\theta\rightarrow 1}(S,\cdot)}[\mathrm{Var}[f(X_{1})|X_{\theta}=S,\mathcal{P}_{i}]]]$ by Eq. 148
[…]
which establishes Eq. 145 and completes the proof of the theorem. ∎
11 Random Regular Graphs via Field Dynamics
In this section we prove the result of Chen, Chen, Chen, Yin, and Zhang [CCC+25] establishing fast mixing of the Glauber dynamics on random $\Delta$-regular graphs beyond the uniqueness threshold. We restate the formal statement of their result for convenience.
See 1.13
11.1 Proof Overview
The proof of Theorem 1.13 uses a very nice connection for the hard-core model between spectral independence and the minimum eigenvalue of the adjacency matrix $A(G)$ for sufficiently small $\lambda$. Let $f(\lambda)$ be the optimal constant $\eta$ for $\eta$-weak spectral independence of $\mu_{\lambda}$. Namely, we will show:
$f(\lambda)\leq 1+\lambda\,|\lambda_{\min}(A_{G}+I)|+o(\lambda).$ (151)
We then use the following lemma, which is a continuous analog of trickle-down, to conclude the bound on spectral independence for all $\lambda<1/|\lambda_{\min}(A_{G}+I)|$.
Lemma 11.1.
For any $\lambda>0$ we have
$D^{-}f(\lambda)\leq\frac{f(\lambda^{2})-f(\lambda)}{\lambda},$ (152)
where $D^{-}f$ is the left derivative of $f$.
Intuitively, the lemma allows us to extend the bound on $f(\lambda)$ from Eq. 151 to larger values of $\lambda$. More precisely, we will show that:
$f(\lambda)\leq\frac{1}{1-\lambda|\lambda_{\min}(A_{G}+I)|}$
when $\lambda<1/|\lambda_{\min}(A_{G}+I)|$. We then utilize the classical result of Friedman [Fri03] that, with high probability, a random $\Delta$-regular graph $G$ has $\lambda_{\min}(A_{G})\geq-2\sqrt{\Delta}$.
We first prove Eq. 151 and then we utilize Eq. 151 and Lemma 11.1 to prove Theorem 1.13; we then prove Lemma 11.1 in Section 11.2.
Let us define $f(\lambda)$ more formally. For $\lambda>0$, let $\widetilde{D}_{\lambda}={\mathrm{diag}}(\mathbb{E}(\mu_{\lambda}))$ be the diagonal matrix with entries $\widetilde{D}_{\lambda}(u,u)=\mu_{\lambda}(\sigma(u)=1)$ for $u\in V$. Recall from Section 2.5, $\eta$-weak spectral independence is defined as $\lambda_{\max}(\widetilde{\Psi_{\lambda}})\leq 1+\eta$, which is equivalent to:
$\mathrm{Cov}_{\mu_{\lambda}}\preceq(1+\eta)\widetilde{D}_{\lambda}.$ (13)
For a PSD matrix $A$, let $\rho(A)=\lambda_{\max}(A)$ be the spectral radius of $A$, and then for $\lambda>0$, let
$f(\lambda)=\max_{\tau}\rho(\widetilde{D}_{\lambda}^{-1/2}\mathrm{Cov}_{\lambda}\widetilde{D}_{\lambda}^{-1/2})$ (153)
where $\tau$ is a pinning. Note, Eq. 13 holds with $\eta=f(\lambda)-1$ for all pinnings $\tau$ and with respect to $\mu_{\lambda}^{\tau}$.
Proof of Eq. 151.
Fix the graph $G$ under consideration, and hence for notational convenience let $\mu_{\lambda}=\mu_{G,\lambda}$. To establish $f(\lambda)$-weak spectral independence, recall from Section 2.5, we need to show $\mathrm{Cov}_{\mu_{\lambda}}\preceq f(\lambda)\widetilde{D}_{\lambda}$ where $\widetilde{D}_{\lambda}$ is the diagonal matrix with entries $\widetilde{D}_{\lambda}(u,u)=\mu_{\lambda}(\sigma(u)=1)$ for $u\in V$.
To establish the bound on $f(\lambda)$ note that for $\lambda$ sufficiently small (e.g., $\lambda<<1/n^{2}$) then for $X_{1}\sim\mu_{\lambda}$ we have $\mathrm{Pr}\left(|X_{1}|\geq 3\right)=o(\lambda^{2})$. Hence, we have the following for any $u\neq v\in V$:
$\mathrm{Cov}_{\mu_{\lambda}}(u,v)$ $=\mathrm{Pr}[\{u,v\}\subseteq X_{1}]-\mathrm{Pr}[u\in X_{1}]\cdot\mathrm{Pr}[v\in X_{1}]$
$=(\mathrm{Pr}[X_{1}=\{u,v\}]+o(\lambda^{2}))-(\mathrm{Pr}[X_{1}=\{u\}]+o(\lambda))(\mathrm{Pr}[X_{1}=\{v\}]+o(\lambda))$
$=\mathrm{Pr}[X_{1}=\{u,v\}]-\lambda^{2}+o(\lambda^{2}).$
Since $X_{1}$ is an independent set weighted by $\lambda^{|X_{1}|}$ we have:
$\mathrm{Pr}[X_{1}=\{u,v\}]=\begin{cases}0&\mbox{if $\{u,v\}\in E$}\\ \lambda^{2}+o(\lambda^{2})&\mbox{if $\{u,v\}\notin E$}\end{cases}$
In other words,
$\mathrm{Pr}[X_{1}=\{u,v\}]-\lambda^{2}=-\lambda^{2}A_{G}(u,v)+o(\lambda^{2}).$
Now for the diagonal entries of the covariance matrix we have:
$\mathrm{Cov}_{\mu_{\lambda}}(u,u)=\mathrm{Pr}[u\in X_{1}]-\mathrm{Pr}[u\in X_{1}]^{2}=\mathrm{Pr}[u\in X_{1}]-\lambda^{2}+o(\lambda^{2}).$
Since $\widetilde{D}_{\mu_{\lambda}}(u,u)=\mathrm{Pr}[u\in X_{1}]$, we have:
$\mathrm{Cov}_{\mu_{\lambda}}-\mathrm{diag}(\mathbb{E}[X_{1}])=-\lambda^{2}(A_{G}+I)+o(\lambda^{2}).$ (154)
Combining the above bounds we obtain the following:
$\mathrm{Cov}_{\mu_{\lambda}}$ $=\mathrm{Cov}[X_{1}]$ since $X_{1}\sim\mu_{\lambda}$ (155)
$=\mathrm{diag}(\mathbb{E}[X_{1}])-\lambda^{2}(A_{G}+I)+o(\lambda^{2})$ by Eq. 154
$\preceq\mathrm{diag}(\mathbb{E}[X_{1}])+\lambda^{2}I|\lambda_{\min}(A_{G}+I)|+o(\lambda^{2})$ by Eq. 157
$=\mathrm{diag}(\mathbb{E}[X_{1}])(1+\lambda\,|\lambda_{\min}(A_{G}+I)|+o(\lambda))$ $\mbox{since }\mathrm{diag}(\mathbb{E}[X_{1}])=\lambda I+o(\lambda)$
$=\widetilde{D}_{\lambda}(1+\lambda\,|\lambda_{\min}(A_{G}+I)|+o(\lambda)).$ (156)
This establishes the bound on $f(\lambda)$ in the case when there is no pinning $\tau$. For a pinning $\tau$, consider the corresponding induced subgraph $H$. Note, $\lambda_{\min}(A_{G})\geq-2\sqrt{\Delta}$ is equivalent to
$A_{G}\succeq\lambda_{\min}(A_{G})I,$
which implies:
$A_{G}\preceq|\lambda_{\min}(A_{G})|I$ (157)
Note, for any induced subgraph $H$ of $G$ we also have:
$A_{H}\preceq|\lambda_{\min}(A_{G})|I,$ (158)
this follows from $\lambda_{\min}(A_{G})\leq\lambda_{\min}(A_{H})$. Now we apply Eq. 158 in Eq. 156 and we obtain:
$f(\lambda)\leq 1+\lambda\,|\lambda_{\min}(A_{G}+I)|+o(\lambda).$
We can now prove the main result Theorem 1.13.
Proof of Theorem 1.13.
Note that for any $c$ for $g=\frac{1}{1-\lambda c}$ we have
$g^{\prime}(\lambda)=D^{-}g(\lambda)=\frac{g(\lambda^{2})-g(\lambda)}{\lambda}.$ (159)
Hence if $f(\lambda^{\prime})\geq\frac{1}{1-\lambda^{\prime}c}$ for some $c$ and $\lambda^{\prime}<1/c$ then from (152) and (159) we have that
$f(\lambda)\geq\frac{1}{1-\lambda c}\ \mbox{for all}\ \lambda\in[0,\lambda^{\prime}].$ (160)
Let $c:=|\lambda_{\min}(A_{G}+I)|$. If there exists $\lambda^{\prime}<1/c$ where $f(\lambda^{\prime})>1/(1-\lambda^{\prime}c)$ then for some $c^{\prime}>c$ we have $f(\lambda^{\prime})>1/(1-\lambda^{\prime}c^{\prime})$, and by Eq. 160, for all $\lambda\leq\lambda^{\prime}$ we have $f(\lambda)\geq 1/(1-\lambda c^{\prime})$. But Eq. 151 states that $f(\lambda)\leq 1+\lambda c+o(\lambda)$ and hence we need that $1/(1-\lambda c^{\prime})\leq 1+\lambda c+o(\lambda)$ which is false for all sufficiently small $\lambda$, and thus we have a contradiction for the assumption that $f(\lambda^{\prime})>1/(1-\lambda^{\prime}c)$ for some $\lambda^{\prime}<1/c$. Therefore, for all $\lambda<1/|\lambda_{\min}(A_{G}+I)|$, we have:
$f(\lambda)\leq\frac{1}{1-\lambda|\lambda_{\min}(A_{G}+I)|}.$ (161)
Now we will apply Theorem 1.8 to finish off the proof. By [Fri03], we know that for random $\Delta$-regular graphs, with probability 1o(1)1-o(1) over the choice of the graph $G$, we have that $\lambda_{\min}(A_{G})\geq-2\sqrt{\Delta}$.
Let $\lambda^{*}(\Delta):=\frac{1}{3\sqrt{\Delta}}$. From Eq. 161, we have that the measure $\mu_{\lambda^{*}}$ is $f(\lambda^{*})$-spectrally independent where
$f(\lambda^{*})=1/(1-\lambda^{*}|\lambda_{\min}(A_{G}+I)|)\leq 1/(1-2\lambda^{*}\sqrt{\Delta})\leq 3.$
Applying Theorem 1.8 (where $b$-marginal boundedness is easily satisfied with $b=b(\Delta)$ since $\Delta$ is a constant), establishes the theorem. ∎
11.2 Proof of the trickle-down for the field dynamics
Here we prove the main technical lemma which is a trickle-down result for the field dynamics.
Proof of Lemma 11.1.
Let $X_{1}\sim\mu_{\lambda}$. We first consider the first order expansion of $\mathrm{Cov}$ for the field dynamics, following [CCC+25, Proposition 4.1]. Recall the definition of the field dynamics in Section 10.1, and, in particular, the definition of $X_{\theta}$ for $0\leq\theta\leq 1$ in Definition 10.5. For $h>0$, we have the following, which we will prove at the end of this section:
$\mathbb{E}[\mathrm{Cov}[X_{1}|X_{h}]]=\mathrm{Cov}[X_{1}]-h\mathrm{Cov}[X_{1}]\mathrm{diag}(\mathbb{E}[X_{1}])^{-1}\mathrm{Cov}[X_{1}]+o(h).$ (162)
By the definition of $f$ and the fact that $X_{1}$ conditioned on $X_{h}$ is according to the up-process $U_{h\rightarrow 1}(X_{h},\cdot)$ which is the hard-core model with fugacity $(1-h)\lambda$, hence we have for any $0<h<1$:
$\mathrm{Cov}[X_{1}|X_{h}]\preceq f((1-h)\lambda)\mathrm{diag}(\mathbb{E}[X_{1}|X_{h}]-\mathbbm{1}_{X_{h}}),$ (163)
where $1_{X_{h}}$ is one for coordinates $i\in X_{h}$ and zero otherwise. Taking the expectation over $X_{h}$ for both sides in Eq. 163 yields:
$\mathbb{E}[\mathrm{Cov}[X_{1}|X_{h}]]\$$\preceq f((1-h)\lambda)\mathrm{diag}(\mathbb{E}[\mathbb{E}[X_{1}|X_{h}]]-h\mathbb{E}[X_{1}])$ (164)
$=f((1-h)\lambda)(1-h)\mathrm{diag}(\mathbb{E}[X_{1}]).$ (165)
Let $\tau$ be the maximizer in Eq. 153 for $\mu_{\lambda}$. Let $v$ be the eigenvector in (153) for eigenvalue $f(\lambda)$; w.l.o.g. assume that $\|v\|_{2}=1$. We can rewrite Eq. 162 as
$\mathrm{Cov}[X_{1}]=\mathbb{E}[\mathrm{Cov}[X_{1}|X_{h}]]+h\mathrm{Cov}[X_{1}]\mathrm{diag}(\mathbb{E}[X_{1}])^{-1}\mathrm{Cov}[X_{1}]+o(h)$ (166)
We have (combining (166) and (165))
$f(\lambda)$ $=v^{T}\,\mathrm{diag}(\mathbb{E}[X_{1}])^{-1/2}\mathrm{Cov}[X_{1}]\mathrm{diag}(\mathbb{E}[X_{1}])^{-1/2}v$ (167)
$\leq f((1-h)\lambda)(1-h)\|v\|_{2}^{2}+v^{T}(\mathrm{diag}(\mathbb{E}[X_{1}])^{-1/2}\mathrm{Cov}[X_{1}]\mathrm{diag}(\mathbb{E}[X_{1}])^{-1/2})^{2}v+o(h).$ (168)
Since $\|v\|_{2}=1$ and $\mathrm{diag}(\mathbb{E}[X_{1}])^{-1/2}\mathrm{Cov}[X_{1}]\mathrm{diag}(\mathbb{E}[X_{1}])^{-1/2})v=f(\lambda)v$ (because $v$ is the eigenvector in Eq. 153 with eigenvalue $f(\lambda)$), and hence we have
$f(\lambda)\leq(1-h)f((1-h)\lambda)+f(\lambda)^{2}h+o(h).$ (169)
Equation (169) implies the following lower bound on the left derivative of $f$ as stated in Lemma 11.1. ∎
Finally, it remains to prove Eq. 162. This corresponds to [CCC+25, Proposition 4.1], which is a special case of more general localization schemes presented in [AKV24]. We present a more pedestrian proof here.
Proof of Eq. 162.
With probability $1-O(h^{2})$ we have $|X_{h}|\leq 1$ and hence, in order to obtain an expression with error $O(h^{2})$, we only need to consider the options $|X_{h}|=0$ and $|X_{h}|=1$. We have
$\mathrm{Pr}(X_{h}=\{\ell\})=h\mathrm{Pr}(\ell\in X_{1})+O(h^{2})=h\mathbb{E}[X_{1}]_{\ell}+O(h^{2}).$
For any $S\subset V$ we have
$\mathrm{Pr}(X_{1}=S|X_{h}=\{\ell\})=\frac{\mathrm{Pr}(X_{1}=S,\ell\in X_{1})}{\mathrm{Pr}(\ell\in X_{1})}+O(h).$
Hence for $i,j$ we have
$\mathrm{Cov}[X_{1}|X_{h}=\{\ell\}]_{i,j}$ $=\frac{\mathrm{Pr}(\{i,j,\ell\}\subseteq X_{1})}{\mathrm{Pr}(\ell\in X_{1})}-\frac{\mathrm{Pr}(\{i,\ell\}\subseteq X_{1})}{\mathrm{Pr}(\ell\in X_{1})}\frac{\mathrm{Pr}(\{j,\ell\}\subseteq X_{1})}{\mathrm{Pr}(\ell\in X_{1})}+O(h).$ (170)
Aggregating (170) over $\ell$ weighted with probability $\mathrm{Pr}(X_{h}=\{\ell\})=h\mathrm{Pr}(\ell\in X_{1})+O(h^{2})$ we obtain
$\sum_{\ell}\mathrm{Pr}(X_{h}=\{\ell\})\mathrm{Cov}[X_{1}|X_{h}=\{\ell\}]_{i,j}$
$=h\sum_{\ell}\mathrm{Pr}(\{i,j,\ell\}\subseteq X_{1})-h\sum_{\ell}\mathrm{Pr}(\{i,\ell\}\subseteq X_{1})\frac{1}{\mathrm{Pr}(\ell\in X_{1})}\mathrm{Pr}(\{j,\ell\}\subseteq X_{1})+O(h^{2}).$ (171)
We also have
$\mathrm{Pr}(X_{1}=S|X_{h}=\emptyset)=\mathrm{Pr}(X_{1}=S)\Big{(}1+h\sum_{\ell}\mathrm{Pr}(\ell\in X_{1})-h|S|\Big{)}+O(h^{2}).$
Note that we can write
$\sum_{S}\mathrm{Pr}(X_{1}=S,i,j\in X_{1})|S|=\sum_{\ell}\mathrm{Pr}(\{i,j,\ell\}\subseteq X_{1}).$
and hence
$\mathrm{Pr}(\{i,j\}\subseteq X_{1}|X_{h}=\emptyset)=\mathrm{Pr}(\{i,j\}\subseteq X_{1})\Big{(}1+h\sum_{\ell}\mathrm{Pr}(\ell\in X_{1})\Big{)}-h\sum_{\ell}\mathrm{Pr}(\{i,j\}\subseteq X_{1})+O(h^{2}).$ (172)
Taking the analogous expressions to (172) for $\mathrm{Pr}(k\in X_{1}|X_{h}=\emptyset)$ for $k=i,j$ we obtain the following
$\mathrm{Cov}[X_{1}|X_{h}=\emptyset]_{i,j}$ $=\mathrm{Pr}(\{i,j\}\subseteq X_{1})\left(1+h\sum_{\ell}\mathrm{Pr}(\ell\in X_{1})\right)$ (173)
$-\mathrm{Pr}(i\in X_{1})\mathrm{Pr}(j\in X_{2})\left(1+2h\sum_{\ell}\mathrm{Pr}(\ell\in X_{1})\right)$
$+h\sum_{\ell}\left(\mathrm{Pr}(i\in X_{1})\mathrm{Pr}(\{j,\ell\}\subseteq X_{1})+\mathrm{Pr}(j\in X_{1})\mathrm{Pr}(\{i,\ell\}\subseteq X_{1})\right)+O(h^{2}).$
Note that
$\mathrm{Pr}(X_{h}=\emptyset)=1-h\sum_{\ell}\mathrm{Pr}(\ell\in X_{1})+O(h^{2}).$ (174)
Combining (171), (11.2), and (174) we obtain
$\mathbb{E}[\mathrm{Cov}[X_{1}|X_{h}]]_{i,j}=\mathrm{Cov}[X_{1}]_{i,j}-hB_{i,j}+O(h^{2}),$
where
$B_{i,j}=$ $\sum_{\ell}\frac{\mathrm{Pr}(\{i,\ell\}\subseteq X_{1})\mathrm{Pr}(\{j,\ell\}\subseteq X_{1})}{\mathrm{Pr}(\ell\in X_{1})}-\sum_{\ell}\mathrm{Pr}(i\in X_{1})\mathrm{Pr}(\{\ell,j\}\subseteq X_{1})$
$-\sum_{\ell}\mathrm{Pr}(j\in X_{1})\mathrm{Pr}(\{\ell,i\}\subseteq X_{1})+\sum_{\ell}\mathrm{Pr}(j\in X_{1})\mathrm{Pr}(i\in X_{1})\mathrm{Pr}(\ell\in X_{1}).$
Finally, we note that $B_{i,j}=(\mathrm{Cov}[X_{1}]\mathrm{diag}(E[X_{1}])^{-1}\mathrm{Cov}[X_{1}])_{i,j}$. ∎
12 Extensions to Other Spin Systems: Ising Model and Colorings
Here we detail the extension of the results to other models beyond the hard-core model on weighted independent sets. In Section 12.1 we survey results for the Ising model, and in Section 12.2 we survey results for colorings and edge colorings. The results stated in this section are not proved in this monograph.
12.1 Ising Model
The Ising model is the classical model from statistical physics. Here we state the Ising model analogs of the hard-core model results from Section 1.6.
The Ising model is defined on a graph $G=(V,E)$ with a parameter $\beta\in\mathbb{R}$ which is referred to as the inverse temperature. The state space of the model is $\Omega=\{+1,-1\}^{V}$. For a configuration $\sigma\in\Omega$, let $m(\sigma)=|\{\{v,w\}\in E:\sigma(v)=\sigma(w)\}|$ denote the number of monochromatic edges, and define its weight as
$w(\sigma)=\exp(2\beta m(\sigma))=\exp\left(\beta\sum_{\{v,w\}\in E}\left(\sigma(v)\sigma(w)+1\right)\right)=\exp(\beta|E|)\exp\left(\beta\sum_{\{v,w\}\in E}\sigma(v)\sigma(w)\right).$
The Gibbs distribution is defined on $\Omega$ where for $\sigma\in\Omega$ let $\mu(\sigma)=w(\sigma)/Z$ and $Z=\sum_{\eta\in\Omega}w(\eta)$ is the partition function.
The model is ferromagnetic (sometimes referred to as attractive) when $\beta\geq 0$ as neighboring spins prefer to align, and the model is antiferromagnetic (repulsive) when $\beta<0$. The generalization to $q\geq 3$ spins is known as the Potts model. The results stated in this section only hold for the Ising model (i.e., when $q=2$).
Recall, the tree uniqueness threshold for the hard-core model occurs at $\lambda_{c}(\Delta)=(\Delta-1)^{\Delta-1}/(\Delta-2)^{\Delta}$. In the Ising model the corresponding tree uniqueness threshold occurs at $\beta_{c}=\beta_{c}(\Delta)$ satisfying one of the following equivalent conditions:
$\exp(2|\beta_{c}|)=\frac{\Delta}{\Delta-2}\iff\tanh(|\beta_{c}|)=\frac{1}{\Delta-1}.$
Recall, the phase transition on $\Delta$-regular trees in the hard-core model corresponds to even-height versus odd-height trees. In contrast, in the Ising model the phase transition on $\Delta$-regular trees corresponds to fixing all leaves to spin +1+1 versus fixing all leaves to spin 1-1. For integer $\ell\geq 1$, recall $T_{\ell}$ denotes the complete $\Delta$-regular tree of height $\ell$. For $s\in\{+1,-1\}$, let $\mu_{\ell}^{s}$ denote the Gibbs distribution on $T_{\ell}$ conditional on all leaves having the fixed spin $s$, and let $p_{\ell}^{s}$ denote the marginal probability that the root receives spin +1+1 in the conditional Gibbs distribution $\mu^{s}_{\ell}$.
When $\beta$ satisfies $\exp(2|\beta|)\leq\frac{\Delta}{\Delta-2}$ then
$\lim_{\ell\rightarrow\infty}p^{+1}_{2\ell}=\lim_{\ell\rightarrow\infty}p^{-1}_{2\ell}$ (175)
and we say the model is in the tree uniqueness region, whereas when $\exp(2|\beta|)>\frac{\Delta}{\Delta-2}$ then the limits are different and we say the model is in the tree non-uniqueness region (as there are multiple Gibbs measures on the infinite $\Delta$-regular tree). Notice that in Eq. 175 we only considered trees of even height; for the ferromagnetic Ising model this is unnecessary, we can consider $\lim_{\ell\rightarrow\infty}p^{+1}_{\ell}=\lim_{\ell\rightarrow\infty}p^{-1}_{\ell}$ to obtain the same threshold. However, for the antiferromagnetic Ising model we need to restrict attention to even height trees; alternatively, we can consider even height vs. odd height trees (both with all +1+1-leaf configuration) and obtain the same threshold.
The analog of Theorem 1.9 establishing fast mixing of the Glauber dynamics on any graph of maximum degree $\Delta$ in the tree uniqueness region was proved in Mossel and Sly [MS13] for the ferromagnetic Ising model and in Chen, Liu, and Vigoda [CLV21a] for the antiferromagnetic Ising model.
Theorem 12.1.
For all $\Delta\geq 3$, all $\delta>0$, there exists $C(\Delta,\delta)$, for all $\beta$ where $\exp(|\beta|)\leq\frac{\Delta-\delta}{\Delta-2+\delta}$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the mixing time of the Glauber dynamics for the Ising model satisfies for all $\varepsilon>0$,
$T_{\mathrm{mix}}(\varepsilon)\leq C(\Delta,\delta)n\log(n/\epsilon).$
As in the case of the hard-core model, the independent works of Chen, Feng, Yin, and Zhang [CFYZ22] and Chen and Eldan [CE22] improved Theorem 12.1 to obtain $T_{\mathrm{mix}}(\epsilon)\leq C(\delta)n\log(n/\varepsilon)$ and thus extend the theorem to classes of graphs where the maximum degree $\Delta$ grows as a function of $n$.
At the critical point, Bauerschmidt and Dagallier [BD24] proved polynomial mixing time of the Glauber dynamics. The upper bound on the critical exponent was improved to $3+O(1/\Delta)$ by Chen, Chen, Yin, and Zhang [CCYZ25] as stated in the following theorem; they also established a lower bound on the critical exponent of $3/2$. More general results were subsequently established by Caputo, Chen, and Parisi [CCP25].
Theorem 12.2.
For all $\Delta\geq 3$, for any graph $G=(V,E)$ of maximum degree $\Delta$, for the Glauber dynamics for the Ising model with $\beta=\beta_{c}(\Delta)$ or $\beta=-\beta_{c}(\Delta)$, at fugacity $\lambda=\lambda_{c}(\Delta)$ the mixing time satisfies
$T_{\mathrm{mix}}=O(n^{3+O(1/\Delta)}\log{n}).$
For the antiferromagnetic Ising model, a computational hardness result in the tree non-uniqueness region (analogous to Theorem 1.12) was established in [SS14, GŠV16]. In contrast for the ferromagnetic Ising model there is a polynomial-time algorithm for approximating the partition function for every graph $G=(V,E)$ and every $\beta\geq 0$ due to Jerrum and Sinclair [JS93], which utilizes the so-called high-temperature expansion of the Ising model to consider a computationally equivalent model over even-degree subgraphs; an efficient sampler was obtained via various methods [RW99, GJ09, GJ17].
The analog of Theorem 1.13 for the antiferromagnetic Ising model was proved by Anari, Koehler, and Vuong [AKV24, Corollary 104] establishing fast mixing of the Glauber dynamics on random $\Delta$-regular graphs beyond the uniqueness threshold.
12.2 Multispin Systems: Colorings and Edge Colorings
In this section we outline the extensions of spectral independence to multispin systems; these are models where each vertex has $q\geq 3$ possible labels.
A prominent example of multispin systems is the colorings problem, namely proper vertex $k$-colorings. For simplicitly, we focus attention on the colorings problem instead of considering arbitrary multispin systems.
Given a graph $G=(V,E)$ of maximum degree $\Delta$, and an integer $k\geq 2$, let $\Omega$ denote the collection of proper vertex $k$-colorings of $G$; more formally, let
$\Omega=\{\sigma\in[k]^{V}:\mbox{for all }\{v,w\}\in E,\sigma(v)\neq\sigma(w)\},$
where $k]=\{1,\dots,k\}$. Then the Gibbs distribution $\mu$ is the uniform distribution over $\Omega$ and thus for all $\sigma\in\Omega,\mu(\sigma)=1/|\Omega|$.
The transitions of the Glauber dynamics for colorings are as follows. From a coloring $X_{t}\in\Omega$, we choose a vertex $v$ uniformly at random from $V$, and a color $c$ uniformly at random from $[k]$. If $c\notin X_{t}(N(v))$ then we set $X_{t+1}(v)=c$ and otherwise we set $X_{t+1}(v)=X_{t}(v)$. For all $w\neq v$, we set $X_{t+1}(w)=X_{t}(w)$. In words, we choose a random vertex $v$ for (attempted) recoloring and we choose a random color $c$, if we can recolor $v$ to color $c$ while maintaining a proper coloring then we do so, and otherwise we stay at the current coloring.
There are two natural generalizations of spectral influence to colorings which were introduced in the independent works of Chen, Galanis, Stefankovic, and Vigoda [CGŠV21] and Feng, Guo, Yin, and Zhang [FGYZ21]; both of these approaches extend to aribtrary multispin systems.
The two notions of spectral independence differ in their definition of the influence matrix. Both approaches generalize the defintion of the influence matrix in Definition 1.1 for the special case of $q=2$ spins.
In [CGŠV21] they consider a $qn\times qn$ modified influence matrix, which we denote as $\widehat{\Psi}$; it generalizes the definition of the modified influence matrix in Definition 2.7. For each pair vertices $v,w$, and pair of colors $c_{1},c_{2}\in[k]$, let
$\widetilde{\Psi}((v,c_{1}),(w,c_{2}))=\mu[\sigma(w)=c_{2}|\sigma(v)=c_{1}]-\mu[\sigma(w)=c_{2}],$ (176)
denote the influence of $(\mathrm{vertex},\mathrm{color})$ pair $(v,c_{1})$ on $(w,c_{2})$.
In contrast, [FGYZ21] considers a $n\times n$ influence matrix $\Psi$, which generalizes the definition of the influence matrix in Definition 1.1. For each pair of vertices $v,w$, let
$\Psi(v,w)=\max_{c,c^{\prime}\in[k]}\|\mu[\sigma(w)=\cdot|\sigma(v)=c]-\mu[\sigma(w)=\cdot|\sigma(v)=c^{\prime}]\|_{\mathrm{TV}},$ (177)
denote the total variation distance for the marginal distribution at $w$ conditional on the worst-case pair of colors at $v$.
We say a pinning is a fixed assignment $\tau$ to a subset of vertices $S\subset V$ (i.e., $\tau:S\rightarrow[k]$) such that there is a valid coloring $\sigma$ of the entire graph $G$ which is the same as $\tau$ on $S$. We then generalize the above definitions as $\widetilde{\Psi_{\tau}}$ and $\Psi_{\tau}$ where $\mu$ is replaced by $\mu_{\tau}$, which is the Gibbs measure conditional on the fixed assignment $\tau$.
We say that $\eta$-weak spectral independence holds if we have $\lambda_{\max}(\widetilde{\Psi_{\tau}})\leq 1+\eta$ for all pinnings $\tau$, and $\eta$-spectral independence holds if we have $\lambda_{\max}(\Psi_{\tau})\leq 1+\eta$ for all pinnings $\tau$.
As in the case of two-spin systems (see Remark 2.10), spectral independence implies weak spectral independence.
Remark 12.3.
[…]
We also have the following:
$|\mu[\sigma(w)=c_{2}|\sigma(v)=c_{1}]-\mu[\sigma(w)=c_{2}]|\leq\|\mu[\sigma(w)=\cdot|\sigma(v)=c]-\mu[\sigma(w)=\cdot|\sigma(v)=c^{\prime}]\|_{\mathrm{TV}}$
Therefore, if one transforms the $n\times n$ matrix $\Psi$ into a $kn\times kn$ matrix $\Psi_{k}$ where each entry $\Psi(v,w)$ is replaced by a $k\times k$ block in $\Psi_{k}$ with all entries in this block equal to $\Psi(v,w)$, then we obtain:
$\lambda_{\max}(\widetilde{\Psi})\leq\lambda_{\max}(|\widetilde{\Psi}|)\leq\lambda_{\max}(\Psi_{k})=k\lambda_{\max}(\Psi),$
which proves that $\eta$-spectral independence (with respect to $\Psi$) implies $(k\eta)$-weak spectral independence (with respect to $\widetilde{\Psi}$).
Chen, Liu, and Vigoda [CLV21a] established the analog of Theorem 1.8: spectral independence with respect to $\widetilde{\Psi}$ and marginal boundedness implies optimal mixing time bounds for the Glauber dynamics.
The big conjecture in the area is that whenever $k\geq\Delta+2$ then the mixing time of the Glauber dynamics is $O(n\log{n})$. We review the main results for colorings.
Theorem 12.4 ([Jer95]).
For $k\geq 2\Delta+1$, for any graph $G=(V,E)$ of maximum degree $\Delta$, the mixing time of the Glauber dynamics $T_{\mathrm{mix}}\leq kn\log(n/\varepsilon)$.
Note, the maximum degree $\Delta$ can be a function of $n$ in the above result of Jerrum [Jer95].
This was improved to $k>(11/6)\Delta$ by Vigoda [Vig00] using the following chain known as the flip dynamics. A transition $X_{t}\rightarrow X_{t+1}$ of the flip dynamics starts in the same manner as the Glauber dynamics. From $X_{t}\in\Omega$, we choose a vertex $v$ uniformly at random from $V$, and a color $c$ uniformly at random from all colors $[k]$. Let $S(v,c)$ denote the maximal $(c,X_{t}(v))$-colored component in $X_{t}$ which contains $v$, and let $\ell=|S(v,c)|$ denote its size. With probability $p_{\ell}$ we “flip” $S(v,c)$ to obtain $X_{t+1}$, where flipping $S(v,c)$ means interchanging colors $c$ and $X_{t}(c)$; and with probability $1-p_{\ell}$ we set $X_{t+1}=X_{t}$. Note, when $p_{1}=1$ and $p_{j}=0$ for all $j\geq 2$ then the flip dynamics is equivalent to the Glauber dynamics.
Vigoda [Vig00] introduced a set of flip probabilities with the following three properties: (i) $p_{1}=1$, (ii) $p_{i}\geq p_{i+1}$ for all $i\geq 1$, and (iii) $p_{i}=0$ for all $i\geq 7$. Later improvements [CDM+19, CV25] used slightly different flip probabilities but still maintained these three properties (i)-(iii). Under property (iii), the flip dynamics is local in that at most 6 vertices are recolored in every step, and hence the approach of Remark 9.31 applies. Therefore, $O(n\log(n/\varepsilon))$ mixing time of the flip dynamics implies $O(n)$ relaxation time of the Glauber dynamics for constant $\Delta$, and the generalization of Theorem 1.6 implies $O(n\log{n})$ mixing time of the Glauber dynamics.
The result of Vigoda [Vig00] was improved by Chen, Delcourt, Moitra, Perarnau, and Postle [CDM+19] to $k>(11/6-\epsilon)\Delta$ for $\epsilon\approx 10^{-5}$. This was further improved by Carlson and Vigoda [CV25] to $k\geq 1.809\Delta$ as stated in the following result.
Theorem 12.5 ([CV25]).
For $k\geq 1.809\Delta$ and $\Delta\geq 125$, the mixing time of the flip dynamics chain is $O(n\log{n})$ under a setting of the flip probabilities where $p_{i}=0$ for all $i\geq 7$. For constant $\Delta\geq 125$, using the approach of Theorem 1.6 and Remark 9.31, we obtain $O(n\log{n})$ mixing time of the Glauber dynamics.
In [CGŠV21, FGYZ21] it was shown that there is a constant $\eta>0$ where $\eta$-spectral independence holds when $k>1.763\dots\Delta$ for any triangle-free graph with maximum degree $\Delta$, where the constant $1.763$ derives from the root of $x=\exp(1/x)$.
Theorem 12.6 ([CGŠV21, FGYZ21]).
Let $\alpha\approx 1.763\dots$ denote the solution of $x=\exp(1/x)$. For all $\alpha>\alpha^{*}$ and $\Delta\geq 3$, there exists $C=C(\alpha,\Delta)>0$ such that for any $k\geq\alpha\Delta+1$ the Glauber dynamics on any triangle-free graph $G=(V,E)$ of maximum degree $\Delta$ has mixing time $\leq Cn\log{n}$.
Chen and Feng [CF25] extended Theorems 12.5 and 12.6 for optimal relaxation time for graphs with non-constant $\Delta$.
In contrast to Theorems 12.5 and 12.6, the following result of Chen, Liu, Matni, and Moitra [CLMM23] holds down to $k\geq\Delta+3$, which essentially matches the conjectured optimal bound of $k\geq\Delta+2$, but at the expense of a lower bound on the girth.
Theorem 12.7 ([CLMM23, Theorem 1.5]).
For all $\Delta\geq 3$, there exists $g=g(\Delta),C(\Delta)>0$ such that for any graph $G=(V,E)$ of maximum degree $\Delta$ and girth $\geq g$, then for any $k\geq\Delta+3$ the Glauber dynamics has mixing time $\leq Cn\log{n}$.
It is unclear how large the lower bound on the girth is in general, however in the case when $k>(1+\varepsilon)\Delta$ for $\varepsilon>0$ then the girth requirement is roughly $g=C^{\prime}(\varepsilon)\log^{2}{\Delta})$ for some constant $C^{\prime}(\varepsilon)>0$, see [CLMM23, Remark 5.16].
For edge colorings, which correspond to colorings of line graphs, recent work of Wang, Zhang, and Zhang [WZZ24] achieves close to the threshold of twice the maximum degree, since each edge shares an endpoint with $\leq 2(\Delta-1)$ edges so the degree of the line graph is $\leq 2(\Delta-1)$.
Theorem 12.8 ([WZZ24, Theorem 1]).
For all $\Delta\geq 3$, for any $k\geq(2+o_{\Delta}(1))\Delta$, for any $G=(V,E)$ with maximum degree $\Delta$ the Glauber dynamics for $k$-edge-colorings has mixing time $O(n\log{n})$.
In [WZZ24] the $o_{\Delta}(1)$-term in the bound on $k$ is more precisely quantified as $O(\Delta/\log{\Delta})$. They utilize a technique known as matrix trickle-down which was introduced in [ALO22].
By Vizing’s Theorem, we can construct a $k$-edge-coloring for all $k\geq\Delta+1$. It is a fascinating open problem to determine the computational complexity of approximating the number of $k$-edge colorings for $\Delta+1\leq k\leq 2\Delta$. More precisely, an interesting open problem is to improve Theorem 12.8 by establishing an $\mathsf{FPRAS}$ for $k>(2-\varepsilon_{0})\Delta$ for a fixed $\varepsilon_{0}>0$, or to establish computational hardness for $k>(1+\varepsilon_{1})\Delta$ for a fixed $\varepsilon_{1}>0$.
Acknowledgements
[…]